A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7400–7449

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Title word cross-reference

(log log n) [177]. 0 [385]. 1 [385]. 2 [546, 728, 442]. 2b [499]. 3 [566, 148, 475, 543, 546, 568, 284, 548, 728, 442, 539, 559]. 3/2 [166]. 3 × 3 [479]. 4 × 4 [479]. ^{+} [150, 639]. b [55]. d [594]. D^{2} [585]. F [252]. H [591]. K [835, 61, 585, 333, 219, 373]. ⌈n/2⌉ [719]. m [386]. N [342]. O(n log n) [34]. O_{1} [526]. q [747]. SOIQ [680]. T [245]. |m| [759].

- [747]. -3 [757]. -ATPase [150]. -Based [373]. -Gram [342]. -Hitting [594]. -Matching [55]. -Means [585, 219]. -minor-free [591]. -NN [61]. -Ring-Star [386]. -Rough [252]. -Sampling [585]. -tree [67].

/K [150].

1-Median [584]. 1.0 [671]. 10th [930, 911]. 11th [951, 949]. 12-14 [930]. 12th [945, 946, 932]. 13-15 [915, 919]. 13th [926, 935, 941]. 14-17 [939, 918].

14th [954, 948, 934]. 15-17 [916, 920, 917]. 15th [916, 935]. 16-18 [937, 938]. 16th [916, 918]. 17-20 [922, 921]. 17th [947, 943]. 18-20 [927, 926]. 18th [942, 940, 939]. 19-21 [929]. 19-23 [925].

2-Covered [710]. 2-to-1 [122]. 2.5 [631]. 20-22 [940]. 20-23 [935]. 2011 [411, 415, 400, 403, 398, 408, 278, 776, 409, 414, 407, 412, 413]. 2012 [935, 162]. 2012.Proceedings [921]. 23-25 [931]. 23rd [952, 953]. 24-26 [911]. 27-28 [943]. 27-29 [944]. 27-30 [932]. 27-31 [942]. 29-31 [933, 941].

3-5 [950]. 3-6 [947, 954, 948, 952, 953, 949]. 3-7 [951, 955]. 32nd [925].

4-7 [945, 946]. 4th [950].

‘5 [776]. 5-7 [928]. 5-9 [912]. 5th [919, 913].

6th [944, 931, 912, 920].

7th [922, 914, 917].

8-10 [914]. 802.15.4 [425, 347]. 8th [937, 938, 921].

9-10 [913]. 9-12 [934]. 9th [955].

Aarhus [939]. ABC [622]. Abdominal [529]. Abductive [260]. Abella [104]. Abnormal [614]. Absolute [446]. Abstract [648, 369, 109]. Abstraction [629, 751, 108]. Abstractions [515]. Accelerating [684]. Acceptance [200]. Access [891, 425, 894]. Accurate [229]. Achieving [308]. ACO [615]. ACO-Based [615]. Acquisition [669]. across [733]. Act [263]. Action [566, 758, 563, 450, 192]. Active [561, 199, 517, 863]. Activities [184]. Activity [878, 728, 251]. Actuators [282]. Ad [82, 432, 80, 410, 426]. Ad-Hoc [82, 426]. Adams [383]. Adaptable [364]. Adaptation [443, 350, 715, 467, 567]. Adaptive [537, 898, 445, 523, 498, 717, 355, 500, 501, 620, 83, 932, 28]. Adaptive-Width [83]. Adaptively [304]. Add [645]. Additive [124]. Adjectives [259]. Adjust [772]. Adjustment [541]. Adoption [194, 769]. Advances [897, 807, 937, 938, 935, 925, 917]. AES [477]. Affective [185]. Affine [210]. Affirmative [717]. Affordance [447]. after [529]. against [658]. Age [610, 927, 926]. Agencies [785]. Agent [522, 677, 523, 525, 521, 519]. Agent-Based [525, 521]. Agent-Driven [522]. Agents [675, 504]. Aggregate [876, 353]. Aggregating [14]. Aggregation [396]. Aggregations [875]. Agile [870]. Agreement [213]. AI [29]. Aided [102, 323, 228]. Aires [947]. Airports [525]. Algebra [636, 107, 948, 339]. Algebraic [748, 390, 106, 144, 105, 673, 483]. Algorithm [605, 24, 593, 584, 750, 54, 622, 623, 393, 756, 739, 336, 23, 112, 419, 686, 498, 25, 68, 763, 691, 883, 596, 884, 325, 35, 27, 687, 345, 434, 695, 88, 590, 606, 347, 250, 93]. Algorithmic [699, 488, 62]. Algorithms [856, 126, 380, 720, 30, 394, 834, 876, 145, 643, 916, 138, 615, 597, 52, 147, 526, 689, 852, 620, 914, 7, 945, 946]. Algorithms-Based [876]. Aligning [866]. Alignment [261]. Allocation [697, 880, 419, 444, 83, 94, 93]. Allocations [592]. Almost [179, 117]. Almost-Equivalence [179]. Almost-Full [117]. Alphabets [174]. Alternatives [266]. Alzheimer [533]. Ambiguity [471]. Amherst [403]. Ammonium [251]. Amortized [207]. Amounts [337]. AMPLE [83]. Amsterdam [410]. Analog [820]. Analogs [180]. Analyses [654]. Analysis [947, 838, 854, 653, 367, 433, 188, 866, 901, 774, 488, 834, 657, 898, 733, 234, 567, 858, 800, 861, 388, 22, 607, 316, 444, 442, 489, 477, 20, 6, 659, 787, 521, 230, 59, 620, 555, 156, 347, 356, 197]. Analytical [705, 352]. Analytics [848]. Analytics-Driven [848]. Analyzing [631, 365, 836]. Anatomical [530]. Ancient [568]. Angers [911]. Angles [47]. Animals [932]. Animat [459]. Animats [932]. ANN [617]. Annotated [109]. Annotation [850, 263]. Annotations [649]. Annual [940, 935, 925]. ANOVA [619]. Answer [264, 463]. Answer-Set [264]. Answering [847, 416, 664, 414]. Ant [17, 606]. Ant-Miner [606]. ante [901]. Antennae [40]. Anti [343, 726]. Anti-Bayesian [726]. Anti-theft [343]. Antibandwidth [27]. Antimicrobial [251]. AnySURF [290]. AODV [426]. Appearance [530]. Application [749, 315, 665, 134, 485, 362, 908, 498, 698, 711, 251, 706, 345, 716, 201, 223, 250]. Application-Dependent [711]. Applications [648, 947, 944, 891, 715, 758, 212, 761, 952, 953, 912, 686, 913, 692, 509, 729, 914]. Applied [159, 850, 617]. Applying [429, 112, 461, 616]. Approach [877, 569, 633, 411, 681, 390, 150, 247, 224, 837, 622, 337, 869, 107, 134, 576, 249, 334, 218, 353, 670, 13, 289, 850, 86, 608, 103, 6, 612, 611, 801, 230, 364, 83, 828, 220, 783]. Approaches [79, 69]. Approaching [448]. APPROX [916]. Approximability [55]. Approximate [70, 526, 68]. Approximating [540, 131, 135, 139, 600]. Approximation [124, 126, 54, 384, 876, 145, 138, 140, 245, 250, 916]. Approximations [221, 698, 374, 220]. April [929]. Apriori [883]. Architecture [468, 522, 449, 781, 455]. Architectures [363, 945, 946]. Area [42]. Argentina [947]. Argument [793]. Argumentation [847]. Arguments [794, 312, 311]. Arithmetic [481]. Arm [505, 497, 498]. ARQ [422]. Arrangement [42]. Arrays [699]. Arteries [529]. Articles [330]. Artifacts [820]. Artificial [605, 24, 440, 716, 911]. Artificielle [911]. ASPACE [177]. Aspects [233]. Assembly [651, 571, 575, 580, 810]. Assess [785]. Assessing [330]. Assessment [189, 610, 656, 202, 795]. Assignment [380, 394, 355, 696]. Assist [730]. Assistance [201]. Assisted [427, 478]. Associating [154]. Association [202]. Assumptions [638, 208]. Assurance [329]. Asymmetric [387, 494, 724]. Asymptotic [749]. Asynchronous [187]. Athens [929]. Atoms [167]. ATPase [150]. Attack [482, 739]. Attacks [485, 685, 483, 484]. Attenuation [433]. Attractiveness [732].

Attribute [221, 308, 309, 254, 257]. Attribute-Based [308, 309]. Attributes [905]. Attributive [259]. Auction [85]. Auditing [905]. August [927, 915, 944, 919, 926, 942, 940, 916, 935, 933, 931, 922, 912, 913, 925, 920, 939, 943, 914, 921, 918, 941, 932, 917, 776]. Austin [276]. Australia [940]. Austria [954, 955, 952, 953]. Authentication [890, 552, 213]. Authoring [461]. Authorities [776]. Auto [545]. Auto-Calibration [545]. Autocompletion [331]. Automata [629, 634, 163, 816, 179, 165, 180, 166, 181, 809, 518, 520, 5, 824]. Automated [648, 838, 652, 368, 610, 657, 555, 941]. Automatic [874, 415, 108, 556]. Automatically [871]. Automaton [651]. Automobile [627]. Autonomous [503, 633, 504, 935, 894, 502]. Avionics [652]. Aware [12, 432, 880, 855, 424, 861, 894, 840, 833]. Axiallities [887]. Axiom [62].

B [278]. B-Human [278]. Back [15, 36, 77, 160, 215, 274, 375, 473, 527, 582, 661, 803, 737, 282]. Back-Drivable [282]. Backbone [38]. Backdoor [476]. Backgrounds [565]. Backtracking [26]. Bacteria [823]. Bad [264]. Bag [111]. Balance [702]. Balanced [38]. Ball [285]. Band [351, 535]. Bandpass [54]. Bandpass-2 [54]. Bandwidth [421]. Banff [912]. Bank [558]. Barbara [925]. Bargaining [28]. Barrier [34]. Base [427, 94]. Based [877, 632, 838, 856, 532, 360, 681, 900, 315, 674, 890, 897, 547, 189, 837, 188, 623, 675, 30, 107, 252, 18, 656, 745, 868, 348, 876, 814, 739, 740, 319, 467, 344, 494, 297, 550, 235, 371, 585, 234, 791, 615, 466, 468, 577, 552, 522, 506, 562, 663, 677, 417, 895, 437, 308, 334, 190, 218, 455, 193, 862, 261, 246, 227, 492, 670, 295, 13, 738, 17, 316, 321, 289, 79, 69, 664, 882, 690, 513, 158, 883, 711, 831, 608, 810, 672, 741, 242, 243, 489, 373, 560, 309, 551, 426, 667, 611, 525]. Based [521, 322, 255, 679, 149, 95, 340, 81, 156, 331, 342, 799, 88, 232, 240, 245, 257, 565, 435, 723, 347, 324, 250, 241, 830, 579, 490]. Bases [805]. Bayesian [86, 726, 579]. BDD [643]. Be [901, 737, 889, 480]. Beam [687]. Beamforming [93]. Beauty [459]. Bee [605, 24, 880]. Behavior [623, 859, 445, 849, 87, 932]. Behavioral [457]. Behaviors [455, 822]. Behaviour [532, 285, 504, 4, 286]. Behavioural [648]. behind [277]. Belgium [934]. Belief [254]. bench [864]. Benchmark [864]. Benefit [362]. Benefits [312]. Better [600]. between [841, 170, 19]. Beyond [179, 298, 483, 205]. Biases [300]. BIDE [723]. BIDE-Based [723]. Bifurcations [754]. Bin [138]. Binaries [653]. Binding [152]. Bio [503, 451, 455, 448]. Bio-inspired [503, 451, 455, 448]. Bioconductor [149]. Bioinformatics [159, 917]. Biological [30, 524, 149]. Biology [917]. Biometric [899]. Biomimetics [504]. Biomolecular [819]. Bipartite [596, 46]. Bipartite-Like [596]. Biped [506, 293]. Birthday [298]. Birthday-Bound [298]. Bis [251]. Bis-quaternary [251]. Bit [494, 94]. Bitstream [694]. Black [34]. Black-Box [34]. Blackboard [189]. Blackboard-Based [189]. Blended [200, 192]. Block [65, 738]. Blockciphers [298]. Blogosphere [858]. Blue [598]. Board [23]. Body [753, 286]. Boltzmann [727]. Bonded [580]. Bone [610]. Book [399, 400, 401, 402, 404, 406]. Books [398, 405, 409]. Boolean [481]. Border [780]. Both [667]. Bound [134, 298]. Bounded [366, 307, 135, 48, 499]. Bounds [125, 574, 814, 600, 659, 89]. Box [34]. Boxes [479, 712]. BPMN [868]. BPMN-Based [868]. Brain [531]. Braitenberg [442, 499]. Branch [378, 388]. Branch-and-Cut [388]. Branch-and-Price [378]. Branched [845]. Brazil [941, 917]. Brazilian [917, 778]. Breaches [901]. Breaking [34, 299]. Breaks [401]. Breakthrough [476]. Breast [228]. Bridging [108]. Bringing [668]. Bristol [935]. BRMS [669]. Broadband [701]. Broadcast [91]. Browsers [889]. BSB [917]. BUAP [411]. Budget [697]. Budgeting [797]. Buenos [947]. Buffer [880]. Build [779]. Building [771, 505, 190, 193, 65, 328, 886]. Buildings [832]. Bundle [541]. Bursts [430]. Bus [705]. Business [665, 668, 774, 955, 669, 666, 670, 721, 770]. Business-Oriented [721]. Busted [491]. Bytecode [368].

C [649, 365, 654, 108]. CA [925]. Cache [831]. Caen [404]. Calculation [752]. Calculations [759]. Calculators [796]. Calculus [674]. Calibration [546, 545]. Calls [659]. Cambridge [916]. Camera [505]. Cameras [545, 728]. Campo [917]. Can [730]. Canada [912, 913, 920]. Canaries [475]. Cancelation [431]. Cancer [153, 228]. Cantor [105]. Capable [577]. Capacity [333]. Carbon [796]. Card [297]. Cardinality [391]. CASC [948]. Case [24, 872, 189, 792, 190, 364, 197, 779]. Cast [491]. Categorization [287, 850]. Categorizing [335]. Causal [672, 14]. Cause [854]. CBTC [435]. CC [619]. Cell [561, 701]. Cells [453]. Cellular [166, 809, 86, 824]. Census [554]. Center [757, 284]. Centered [42]. Centers [620]. Central [50]. Centric [411, 212, 409, 410, 407, 412]. Certain [205, 57]. Certification [118]. Certified [637]. Certifying [649]. CEUS [532]. Chain [652]. Chains [765]. Challenges [897, 784]. Champions [276]. Chance [384]. Chance-Constrained [384]. Channel [314, 425, 486, 483, 85, 83, 770, 82]. Characterisation [532]. Characteristic [224]. Characterization [748, 238]. Characterizing [84]. Characters [805]. Checking [366, 150, 655, 8, 5, 660, 844]. Chemical [573, 574, 754]. Chemicals [823]. Chengdu [922, 921]. CHES [934]. China [927, 919, 926, 922, 914, 921]. Chinese [200]. Chip [713, 476]. Chlorides [251]. Choice [504, 253, 770]. Chomsky [173]. Choosing [791]. Chordal [595]. Chosen [233]. Chromatic [597]. Chromosomes [144]. CIARP [947]. Ciphers [482, 712]. Ciphertext [309]. Circular [753]. CiteSpace [188]. Citizen [792, 793]. City [771, 798]. Class [741, 169, 606, 609]. Classes [261]. Classic [164]. Classical [163]. Classification [566, 222, 252, 609, 619, 734, 726]. Classifier [744, 151, 716]. Classifiers [287]. Classifying [246]. Cleaners [507]. Climate [796]. Clique [48]. Clique-Width [48]. Closed [433, 723]. Cloud [12, 697, 892, 541, 158, 721, 192, 866]. Cluster [64, 419]. Clustering [857, 125, 745, 744, 624, 740, 235, 585, 153, 253, 741, 743, 620, 219, 232]. Clusters [612]. Cluttered [543]. CNL [933, 461]. Coarse [580].

Coarse-Grained [580]. COCOA [912]. COCOON [940]. Code [494, 303, 700]. Code-Based [494]. Coded [177]. Codes [422, 213]. Codeword [576]. Coding [569, 423, 80, 435]. Coefficients [153]. Coercion [644]. Coercive [769]. Cognition [459]. Cognitive [97, 95, 94, 93]. Cographs [595]. Coherence [906]. Collaborative [189, 638, 190, 192]. Collatz [166]. Collecting [130]. Collection [44, 89]. Collective [936, 524]. Colleges [200]. Collision [485, 448]. Collisions [419]. Collusion [305]. Collusion-Preserving [305]. Collusions [307]. Colony [605, 24, 17]. Color [861, 288]. Color-Sentiment [861]. Colored [201]. Coloured [10, 7]. Colouring [599]. Columns [392]. Combination [744, 287, 852]. Combinatorial [916, 379, 912, 929]. Combinatorics [940]. Combine [3]. Combining [735, 654]. Commands [109]. Commitments [311]. Common [437]. Communication [189, 389, 792, 431, 90, 658, 693, 696, 96, 435]. Communications [99, 425]. Communities [800]. Commutative [231]. Compact [119, 246, 65, 829, 490]. Comparative [555, 156]. Comparators [550]. Compare [440]. Comparing [786, 559, 269]. Comparison [899, 619, 146]. Compass [506]. Compatible [422]. Compensate [433]. Competence [268]. Competition [361, 404]. Competitions [508]. Competitive [35]. Compiler [478]. Complete [177]. Completeness [178, 19]. Completion [565]. Complex [674, 675, 575, 526, 110, 560, 226, 232]. Complexities [167]. Complexity [806, 44, 207, 34, 814, 758, 818, 764, 821]. Compliance [666, 11]. Compliant [449, 13, 282]. Component [733, 522, 59]. Component-Based [522]. Components [641]. Composable [208, 291]. Composite [904]. Compositional [632]. Compositions [827]. Comprehensive [755, 193]. Compression [848, 420, 321, 340]. Compressive [564]. Computation [305, 750, 842, 751, 206, 573, 951, 53, 304, 307, 211, 74, 923, 881, 157, 290, 580, 581]. Computational [107, 936, 381, 524, 917]. Computations [205, 519, 761, 860]. Compute [303]. Computer [947, 838, 102, 228, 948]. Computer-Aided [102]. Computer-Based [838]. Computers [817, 820]. Computing [937, 938, 584, 603, 754, 202, 948, 816, 170, 685, 707, 763, 821, 811, 939, 823, 192, 257, 921, 940]. Concept [201]. Concepts [838, 866]. Conceptual [868, 896, 844]. Concern [795]. Concurrency [910]. Concurrent [703]. Conditional [630]. Conditioning [853]. Conditions [448]. Cone [384, 392, 385]. Conference [928, 915, 919, 954, 951, 955, 926, 940, 911, 935, 931, 922, 952, 953, 912, 913, 925, 949, 920, 939, 950, 914, 945, 946, 921, 918, 941, 932]. Conferencing [13]. Confidence [839]. Confidentiality [371]. Configuration [891]. Configurations [821]. Confusion [488]. Congress [947, 935]. Congruency [534]. Conjecture [166]. Connected [131, 90]. Connectivity [748, 132]. Consecutive [598]. Consensus [526, 520]. Consequent [613]. Considering [859, 696]. Consistency [709, 457]. Consistent [847, 391]. Constrained [697, 384, 41, 132, 391, 395, 541]. Constraint [637, 458, 333, 56]. Constraints [857, 882]. Constructing [538]. Construction [124, 106, 643, 319, 38, 707, 441, 689, 721, 49]. Consultation [794]. Consultations [790]. Consumed [507]. Consumption [907]. Content [399, 534, 930, 404, 560]. Contention [95]. Contention-Based [95]. Contest [8]. Context [175, 170, 171, 266, 172, 855, 176, 199, 894, 769, 174, 676, 269, 770]. Context-Aware [855, 894]. Context-Free [175, 170, 171, 172, 176, 174]. Contextual [886]. Contiguous [588]. Contingences [789]. Contingencies [445, 450]. Continual [260]. Continuity [859]. Continuous [329]. Contour [609]. Contours [561]. Contraction [340]. Control [10, 891, 640, 2, 451, 430, 506, 449, 895, 431, 498, 894, 458, 501, 282, 893, 454, 435]. Controlled [812, 462, 470, 466, 468, 471, 465, 463, 459, 933]. Convergence [931]. Conversational [523]. Convex [46]. Cooperation [94]. Cooperative [97, 83]. Coordinates [754]. Coordination [633, 4, 690]. Coping [471]. Coq [107, 113, 116, 106]. Core [257, 363, 695]. Corner [229]. Correct [835, 115]. Correction [687, 490]. Correctness [205, 643]. Correlated [843]. Correlation [73, 153]. Corruption [211]. Cosine [860]. Cost [649, 131, 41, 194, 904, 589, 429, 56, 361]. Cost-Constrained [41]. Council [798]. Count [181]. Counter [789]. Counter-Democracy [789]. Counterexample [629]. Countries [768]. Course [189, 193, 197]. Cover [122, 392, 598, 342]. Coverage [256, 869, 428]. Covered [710, 19]. Covering [138]. Covers [499]. CPGs [455]. Craniofacial [559]. Credentials [309]. Crete [937, 938]. Criteria [438, 726]. Critical [821, 268, 943]. Critique [799]. Cross [780]. Cross-Border [780]. Crossing [47]. Crowd [525]. Crowdsourcing [802]. CRYPTO [925, 493, 302]. Cryptographic [102, 492, 11, 934]. Cryptography [315, 487, 494]. Cryptology [925]. CSMA [347]. CSMA/CA [347]. CSPs [135]. CT [531]. Cube [750, 881]. Cubes [874]. Cubic [757]. CUDA [157]. Cup [924]. Curious [496]. Current [790, 784, 768, 921]. Curriculum [512]. Cut [388, 51, 393, 65]. Cut-Region [65]. Cuts [588, 741]. Cutting [393, 507]. Cutting-Plane [393]. Cycle [145, 494]. Cycled [96]. Cycles [45, 264, 551].

D [680, 546, 566, 148, 475, 543, 546, 568, 284, 548, 728, 764, 539, 559]. D-Finite [764]. D-LIPM [284]. D/ [546]. Daejeon [931]. Damage [901]. Data [320, 857, 874, 838, 411, 872, 222, 67, 231, 221, 866, 869, 901, 543, 624, 318, 348, 870, 878, 879, 319, 666, 864, 848, 800, 112, 437, 881, 461, 354, 541, 246, 91, 110, 420, 321, 328, 86, 721, 409, 743, 906, 410, 795, 785, 89, 853, 239, 340, 407, 412, 331, 232, 702, 941, 849, 326, 332, 358, 775, 779, 954]. Data-Centric [411, 409, 407, 412]. Data-Driven [239]. Database [854, 311, 349, 952, 953]. Databases [847, 826, 875, 845, 843, 329, 862, 734, 863, 236, 905, 156]. Datacentric [408]. DaWaK [954]. Days [776]. DCT [611]. DCT-Genetic-Fuzzy [611]. Deal [606]. Decade [281]. Decay [759]. December [930]. Decentralised [635]. Decidability [119, 168]. Deciding [578]. Decision [256, 247, 680, 252, 248, 233, 230, 885, 250, 490]. Decisions [217]. Decoder [718]. Decoding [718]. Decreasing [20]. Deductive [366]. Defeasible [677]. Definability [267]. Define [120]. Defined [372]. Definition [620]. Degree [719, 598, 590, 271]. Degree-4 [590]. Degrees [244]. Delay [80]. Delays [278, 693]. Delegation [206, 309]. Deleted [830]. Deletion [813]. Demand [773]. Democracy [789]. Denmark [939, 932]. Denotation [272]. Dense [306, 178, 596]. Densely [499]. Density [232]. Density-Based [232]. Denumerability [105]. Dependent [711]. Deployment [372]. Depth [505, 543, 544]. Derivation [428]. Derivations [172]. Descent [689]. Described [660]. Descriptor [548]. Descriptors [71, 559, 156]. Design [126, 130, 386, 576, 870, 190, 155, 711, 659]. Designing [508, 184]. Designs [652, 906]. Detecting [327]. Detection [688, 368, 897, 735, 547, 554, 561, 533, 235, 614, 558, 567, 738, 288, 289, 717, 616, 149, 250]. Detectors [735]. Determinants [186]. Determiner [270]. Determining [712]. Determinism [165]. Deterministic [573, 176, 697]. Developing [720, 441, 768]. Development [827, 872, 656, 773, 446, 522, 786, 192]. Developmental [822]. Developments [918]. Device [579]. Devices [191]. DEXA [952, 953]. DHE [313]. DIADEM [826]. Diagnosis [228]. Diagnostics [542]. Diagrams [836]. Dialogue [467, 260, 263]. Diameter [143, 41]. Dichotomies [745]. Dictionary [70, 614]. Did [569]. Difference [747]. Different [470, 448, 223]. Differential [747, 482, 31, 103]. Differentiated [875]. Diffusion [344, 858]. Digestion [152]. Digital [955, 568, 558]. Dimension [210]. Dimensional [762, 68, 765, 357, 874, 821]. Dimensionality [732]. Directed [377]. Directional [40]. Directions [923]. Dirty [332]. Discernibility [254]. Discontinuity [436, 557]. Discontinuity-Preserving [557]. Discovering [849]. Discovers [476]. Discovery [954, 843, 227]. Discrete [2]. Discrete-Event [2]. Discrimination [445]. Discussion [187]. Disease [636, 533]. Dishonest [304]. Disjoint [386]. Dispatch [35]. Dispersive [537]. Displacement [574, 581, 579]. Dissimilarity [738]. Distance [72, 191, 73, 497, 446, 170, 597, 408, 157, 882, 729]. Distance-Hereditary [597]. Distortion [545, 433]. Distributed [367, 699, 705, 2, 715, 265, 816, 645, 685, 23, 707, 881, 354, 68, 113, 658, 364]. Distribution [300]. Distributive [270]. Disturbing [187]. DLT [918]. D’MART [328]. DNA [939, 575, 576, 577, 939, 580, 581, 149, 579]. DNAQL [572]. DNase [152]. Do [181]. Document [841, 399, 234, 192]. Document-Sharing [192]. Documents [851]. Does [650]. DOF [500]. Doing [628]. Domain [550, 373, 6]. Domain-Specific [373]. Domains [637, 826]. Domestic [280, 507]. Dominance [218]. Dominance-Based [218]. Dominating [90]. Down [881, 507]. dP [163]. DPA [488, 484]. Drafting [470]. Dragon [162]. Drivable [282]. Driven [872, 665, 773, 848, 908, 522, 539, 239]. Drives [349]. Driving [356]. Droplet [817]. Drug [560]. DS [431]. DS/CDMA [431]. DSACL [67]. Dual [126, 538, 391, 509]. Dual-Rate [509]. during [776]. Dutch [779]. Duty [96]. Duty-Cycled [96]. Dynamic [709, 532, 61, 72, 842, 67, 567, 862, 17, 66, 684, 829, 35, 863, 591, 309, 580, 565, 830]. Dynamics [148, 636, 447, 266, 608].

E-Adoption [194]. E-Beam [687]. E-Consultations [790]. E-Government [769, 787, 784]. e-Learning [199]. e-Petitioning [792]. e-Service [773]. E-Training [194]. EA [911]. Easier [587]. Ecology [608]. Ecology-Based [608]. Economic [35]. ed [405]. Edge [42]. Edge-Centered [42]. Edges [21]. Edit [170, 408, 882]. Edit-Distance [170]. Editing [464]. Education [508, 838, 856, 196, 198]. Educational [820]. Effect [687]. Effective [841, 443, 194, 430, 59]. Effects [150]. Efficiency [778, 312, 786]. Efficient [688, 64, 634, 900, 510, 231, 481, 899, 294, 212, 724, 424, 843, 881, 353, 486, 831, 62, 581, 833, 696, 85]. Effort [653]. EGOV [949]. eGovernment [768]. Electric [452]. Electronic [320, 99, 949, 950]. Element [452]. Elevated [623]. Eliminating [278]. Embedded [512, 658, 934]. Embodied [262]. Embryomorphic [807]. EMD [607]. Emergence [496, 524]. Emergent [822]. Emission [477]. Empire [44]. Employing [718]. Enabled [540]. Enables [159]. Enciphering [297]. Enclave [99]. Encrypted [899, 898]. Encryption [208, 307, 494, 308, 309, 213, 310]. End [678, 302]. Energy [12, 576, 432, 424, 322, 581, 507]. Energy-Aware [12]. Energy-Based [322]. Energy-Efficient [424]. Enforcing [895]. Engaging [511]. Engine [75, 701]. Engineer [440]. Engineering [665, 807, 438, 941]. English [464, 189]. Enhanced [31, 352]. Enhancing [869, 707]. Enough [901]. Enrichment [840]. Enrollment [490]. Ensemble [744]. Enterprise [781]. Entity [236]. Entropy [250]. Environment [350, 618, 707, 198, 539, 356]. Environmental [17]. ePart [950]. eParticipation [791, 793, 801]. Epidemics [636]. Epistemic [266, 672]. Equation [701]. Equational [103]. Equations [747, 752, 344]. Equilibrium [753]. Equivalence [119, 111, 179, 245]. Era [159]. erasing [173]. Erratum [79]. Error [708, 642, 490]. Errors [179, 154, 621]. Escape [503, 21]. ESSLLI [923]. Estimation [543, 542, 332]. ETL [871, 868, 873]. Euler [752]. Eulerian [593]. Evacuation [832]. Evaluating [153]. Evaluation [210, 930, 491, 13, 764, 793, 381, 743, 787]. Evaluations [426]. Evaluative [271]. Evenly [699]. Event [634, 2, 614, 4]. Events [674]. Evidential [253]. Evolution [33, 911, 31, 845, 32, 158, 459]. Evolutionary [29, 30, 23, 25]. Evolving [817, 458]. Ex [901]. Ex-ante [901]. Exact [590]. Examples [300, 823]. Executable [844]. Execution [816, 660]. Executions [667]. Executive [196]. Existential [664]. Existing [328]. Expectations [791]. Experiences [512]. Experiments [408]. Expert [952, 953]. Explanations [682]. Explicit [638, 234]. Exploiting [688, 568, 801]. Exploration [636, 62, 860]. Explorations [605]. Explorative [513]. Exploring [713, 158, 26]. Exponential [486]. Expression [119, 153, 555]. Expressions [226]. Expressive [634, 869, 594]. Extended [390, 389]. Extending [877, 144, 3]. Extension [210]. Extensions [465, 606]. External [402, 769]. Extracting [665]. Extraction [648, 838, 837, 610, 404, 405, 851, 342]. Extreme [392, 616]. Eye [569, 554, 229]. f [303]. Face [736, 196, 732, 735, 622, 737, 733, 249, 738]. Face-to-Face [196]. Facebook [186]. Faces [379]. Faceted [399, 410]. Facial [735, 555]. Facility [134]. Fact [491]. Factor [134, 152, 40]. Factorizations [389]. Factorized [74]. Factors [200]. Fagin [180]. Failure [854, 717]. Fair [592, 434]. Fairness [322]. Family [482]. Far [480]. Fast [394, 74, 764, 698, 884, 96]. Fault [482, 852]. Feature [566, 735, 610, 71]. Features [288, 290]. Featuring [286]. Federations [12]. Feedback [433, 595, 506, 46]. Few [758, 141]. Fi [436]. FIDs [716]. Fields [759, 32]. FIFO [703]. Files [337, 79, 69]. Filter [558]. Filtering [294]. Filters [556]. Financial [521]. Finding [805, 835, 596]. Finger [552]. Fingerprint [436, 324]. Fingerprint-Based [324]. Finite [452, 637, 180, 181, 764, 464]. Finite-State [464]. Finnish [781]. FIRA [935]. First [2]. Fish [452]. Fiteng1000 [722]. Fixed [133, 613]. Flash [831, 349]. Flash-Based [831]. Flexibility [694]. Flexible [657, 902, 290, 280]. FlexRay [655]. Flow [129, 530, 374, 396, 702]. Fluoroscopy [556]. FM [942]. FMICS [943]. Focal [532]. Focused [415, 930]. Folding [148, 147]. Foliage [433]. Footstep [292]. Force [497]. Force-Distance [497]. Forcing [43]. Forecast [822]. Forest [52, 151]. Form [756, 331]. Form-Based [331]. Formal [626, 827, 633, 652, 891, 9, 627, 942, 644, 362, 372, 373, 659, 928, 943]. Formalism [144]. Formalization [116]. Formalize [680]. Formalizing [667]. Formed [244]. Forms [752]. Formulation [802]. Formulations [390, 389]. Forum [461, 327]. Forums [187]. Forwarding [903]. Foundries [852]. Four [753, 53]. Four-Body [753]. Four-Russians [53]. Fourier [127, 698]. FoVeOOS [928]. Fraction [19]. Frame [708, 543]. FrameNet [469]. Framework [367, 29, 665, 575, 644, 3, 184, 663, 851, 13, 896, 690, 672, 373, 853, 565, 357]. France [944, 951, 942, 911, 943]. Fredhopper [364]. Free [175, 170, 171, 172, 176, 592, 174, 591]. Freeness [644]. Frege [268]. French [259]. Frequency [127, 435]. Frequent [884, 325, 340, 723]. Front [16, 37, 528, 549, 570, 583, 604, 625, 646, 662, 683, 704, 725, 746, 767, 788, 804, 825, 846, 867, 888, 909, 58, 78, 100, 121, 142, 161, 182, 203, 216, 237, 258, 275, 296, 317, 338, 359, 376, 397, 418, 439, 460, 474, 495, 516]. Fronts [188]. Fukuoka [945, 946]. Full [404, 308, 25, 117]. Fully [492]. Fulton [763]. Function [605, 573, 210, 764]. Functional [307, 737, 441, 492, 310]. Functions [503, 316, 301, 535, 491]. Fundamental [521]. Fundamentalist [521]. Fusing [544]. Fusion [569]. Future [101, 784, 60]. Fuzzy [247, 246, 611, 225, 240, 245, 223, 220]. Fuzzy-Rough [246].

G [147]. G-Quadruplexes [147]. Gabor [558]. Gait [284, 506, 562, 551, 293]. Game [278, 93, 28]. Games [273, 525]. Gaming [706]. Gap [393, 108]. Gate [577]. Gathered [795]. Gathering [348]. Gaze [444]. GCM [299]. GDMM [927]. Gender [734]. Gene [153, 151, 438, 459]. General [812, 468, 68, 828, 214]. Generalization [267]. Generalized [138, 172, 51, 5]. Generated [399, 578, 712]. Generating [110]. Generation [670, 292]. Generator [492, 293]. Generic [762, 316, 357]. Genericity [760]. Genetic [805, 623, 876, 739, 6, 611, 695]. Genome [144, 822]. Genomic [159]. Genotype [154]. Genre [792]. Geo [779]. Geo-Data [779]. Geodesic [602]. Geographic [323, 851]. Geometric [47, 515, 576, 57]. Geometricity [168]. Geospatial [337]. Germany [930, 797]. Getting [798]. GF [472, 469, 467]. GF-Based [467]. Gibbs [576]. Giving [511]. Glamorgan [183]. Global [4, 80, 71]. Globally [561]. Going [737]. Good [221, 264, 737]. Goodstein [818]. Governance [782]. Government [769, 777, 949, 785, 787, 783, 770, 784]. Government-to-Business [770]. Governmental [775]. GPGPU [25]. GPGPUs [687]. GPU [684, 692]. GPUs [540, 691]. Gradient [289]. Grail [805]. Grain [482, 688]. Grained [580]. Gram [342]. Grammar [469]. Grammars [172]. Grande [917]. granular [231]. Granulation [240, 218]. Graph [42, 336, 417, 340, 342]. Graph-Based [417]. Graph-Structured [340]. Graphs [47, 306, 601, 126, 588, 45, 599, 130, 595, 462, 145, 57, 597, 48, 591, 46, 590, 43]. Grasp [513]. Grayscale [534]. Greece [937, 938, 929]. Grid [127, 453]. Griding [289]. Ground [295]. Ground-Truth [295]. Group [719, 186, 327, 96]. Growth [24]. Guangzhou [919]. Guarantee [120]. Guessing [204]. Guidance [529, 348]. Guide [669]. Guided [629, 378, 444, 542]. Guidelines [470]. Guiding [775].

Halin [597]. Halting [809]. Hamiltonian [45]. Hamiltonicity [714]. Hamming [483]. Hand [543]. Haplotype [154]. Harbin [927, 926]. Hard [724]. Hardness [122, 128, 55, 136]. Hardware [475, 494, 934, 277]. HARM [677]. Harvesting [424]. Hash [231, 316, 882]. Hash-Join [882]. HATS [364]. HDFS [337]. head [181]. Heads [181]. Heaps [766]. Hearted [162]. Helping [98]. Here [628]. Hereditary [597]. HERO [81]. Heterogeneous [98, 696, 695]. Heterotic [823]. Heuristic [146, 254, 695]. Heuristic-Genetic [695]. Hidden [562]. Hierarchical [427, 65, 435]. Hierarchy [454]. High [534, 708, 544, 759, 722, 68, 852, 325, 560, 509, 357]. High- [759]. High-Content [534, 560]. High-Dimensional [357]. High-Performance [722, 325]. High-Precision [852]. High-Resolution [544]. Higher [750, 553, 198]. Highly [85]. HiNet [435]. Hippocampus [533]. Histogram [288]. Histone [152]. Historic [636]. History [563]. Hitting [594]. Hoc [82, 432, 80, 410, 426]. HOGs [547]. Holds [509]. Holy [805]. Home [81, 84]. Homomorphisms [176]. Honest [205]. Hong [198]. Hop [395, 82, 83]. Hopcroft [112]. Hopf [754]. Host [897]. Host-Based [897]. Hotspots [188]. HPA [433]. HSE [656]. Human [451, 563, 728, 151, 278]. Human-Machine [451]. Humanoid [456, 497, 284, 514, 279, 498, 292]. Hybrid [742, 196, 415, 624, 195, 733, 422, 677, 931, 437, 548, 80, 246, 850, 611, 230, 695, 219, 828, 49, 241, 358, 197, 919]. Hybrid-ARQ [422]. Hyperbolic [809]. Hypergraph [598]. Hypergraphs [377, 55]. Hyperpaths [377]. Hyperspectral [351]. Hypertext [10]. Hypervolume [19]. hypotheses [294].

IaaS [697]. Iberoamerican [947]. iBGP [132]. IC0701 [361]. ICA3PP [945, 946]. Ice [514]. ICHIT [931]. ICHL [919]. ICITS [920]. ICTs [782]. Idea [737]. IDEAL [941]. Identification [622]. Identify [568]. Identity [900, 268, 780]. Identity-Based [900]. IEEE [425, 347]. IfAny [816]. IFIP [949, 950]. II [938, 188, 953, 946, 621]. Illumination [249, 552]. Illustrated [643]. Image [452, 569, 947, 550, 334, 75, 71, 557, 621]. Image-Based [334]. Imagery [351]. Images [736, 530, 531]. Imbalance [606]. IMITATOR [631]. Immigrants [17]. Impact [782]. Imperialist [35]. Implementation [370, 155, 769]. Implementations [479]. Implicit [330]. Importance [399]. Imposing [681]. Impossibility [209]. Impression [861]. Impression-Aware [861]. Improve [694, 461]. Improved [605, 125, 593, 128, 54, 622, 432, 741, 242, 27, 590]. Improvement [735]. Improving [24, 774, 358]. IMS [13]. IMS-Compliant [13]. In-situ [848]. Inapproximability [123]. Incentive [83]. Incidents [854]. Include [144]. Including [545]. Inclusion [592]. Inclusion-Free [592]. Incomplete [222, 224, 233, 743]. Inconsistency [682, 269]. Incorrect [835]. Incremental [842, 883, 333, 255, 860, 885, 220]. Independent [758]. Index [394, 597, 721, 63, 357]. Indexing [234, 848, 354, 65, 79, 69, 66, 521]. Indicators [19]. Indices [70]. Indonesian [769]. Induced [587, 392]. Induction [222, 221, 246, 248, 255, 617]. Inductive [806, 818]. Industrial [943]. Industry [362]. Inequalities [391, 393]. Inequality [227]. INEX [930, 412, 411, 415, 400, 416, 403, 398, 408, 417, 406, 409, 414, 407, 413]. INEX-2011 [413]. Inference [233, 154, 611, 149]. Influence [769, 521]. Influences [446]. Influencing [186]. Influential [200]. Infobright [226]. Information [778, 927, 224, 82, 926, 319, 619, 873, 858, 855, 786, 408, 931, 17, 607, 822, 557, 374, 795, 893, 679, 213, 245, 352, 920]. Information-Enhanced [352]. Information-Preserving [607]. Information-Theoretic [213]. Infrastructure [894]. Inheritance [115]. Inhibitory [256]. Initial [10]. Initialisation [556]. Initiative [930, 792]. Initiatives [796, 771]. Inner [593]. Inscriptions [568]. Insertion [813]. Insight [750]. Insights [507]. Inspired [449, 503, 451, 455, 448]. Instance [614, 315]. Instant [331]. Institutional [183, 769]. Instruction [196]. Instructions [511]. Instructive [300]. Instructor [200]. Instrumentation [660]. Integer [750]. Integral [554]. Integrating [152, 873]. Integration [366, 32, 453, 522]. Integrity [231, 329]. Intelligence [936, 716]. Intelligent [18, 523, 345, 941]. Inter [419, 423]. Inter-cluster [419]. Inter-session [423]. Interaction [365, 447, 880, 185, 280, 291, 197, 770]. Interaction-Aware [880]. Interactions [678]. Interactive [312, 234, 915]. Interception [285]. Interconnect [708]. Intercultural [189]. Interestingness [613, 437]. Interests [840]. Interface [451, 39, 500]. Interfaces [466]. Interference [431, 502]. Internal [685]. International [927, 937, 938, 928, 915, 944, 919, 954, 951, 955, 926, 948, 930, 942, 940, 916, 911, 933, 931, 922, 952, 953, 912, 929, 913, 934, 949, 920, 939, 943, 950, 914, 945, 946, 921, 918, 941, 932]. Internet [907]. Interpolated [436]. Interpretation [618, 109].

Intersection [763]. Interval [223]. Intervals [843, 592]. Intra [423]. Intra-session [423]. Intrinsic [732]. Introducing [816]. Introduction [727, 810, 442, 517]. Intrusion [897, 250]. intrusive [87]. Intuitionistic [247, 245]. Intuitive [500, 280]. Invariant [249, 761]. Invariants [632, 642, 103]. Invasions [30]. Inversion [765]. Inverted [354, 79, 69]. Investigate [150]. Investigating [783]. Investigation [623, 186, 396]. Invisible [343]. Involutions [815]. Involutive [755]. IP [381, 426]. IP-Based [426]. IR [415]. Iris [739, 550]. Iris-Based [739]. Irish [777]. IRIT [416]. ISCO [929]. iSCSI [350, 702]. Island [30]. Isolation [762, 363]. Isomorphism [587]. Issues [775]. ISVC [937, 938]. Italy [928]. Item [843]. Items [887]. Itemset [883, 884, 325]. ITP [915]. Itself [779]. IWSN [927].

J2EE [891]. Jacobson [756]. Japan [945, 946]. Japanese [795]. Java [751, 368, 665, 370, 374, 659]. JMS [648]. Jobs [133]. Join [353, 882, 882]. Joins [131, 882]. Joint [384, 935, 93]. Jointly [435]. Judicial [775]. Judiciary [778]. Juggrnaut [369]. Julienne [630]. July [937, 938]. Justifications [265]. JVM [369].

Kalman [294]. Keeps [641]. Kenneth [127]. Kernel [740]. Kernel-Based [740]. Kernelization [594]. Key [492, 213]. Key-Agreement [213]. Keypoint [547]. Keys [834]. Kind [162]. Kinematic [505]. Kinetics [758]. Kit [441]. Knapsack [589]. Know [303]. Knowledge [854, 311, 188, 207, 954, 265, 212, 437, 190, 193, 922, 227, 233]. Kong [198]. Korea [931]. Kripkean [272]. Kristiansand [949, 950].

Label [122, 1]. Labeling [829, 830]. Labels [830]. Ladders [832]. Landmark [530, 882]. Landmark-Join [882]. Landmarks [559]. Landscapes [576]. Language [462, 865, 170, 195, 470, 466, 468, 578, 933, 923, 471, 465, 464, 463, 6, 918]. Languages [168, 167, 175, 177, 176, 174, 373, 169, 310]. Laparoscopic [529]. Large [832, 32, 354, 862, 734, 62]. Large-Scale [354, 862]. Laser [295]. Latencies [658]. Latency [718]. Lateness [586]. Later [281]. Lattice [148]. Lattices [137]. Lausanne [798]. Layers [255]. Layout [123, 837]. Layout-Based [837]. LBP [738]. LDA [235]. LDPC [422]. Leader [327]. LeadingOnes [34]. League [276, 284, 279]. Leakage [82, 487, 486, 483]. Leakage-Resilient [487, 486]. Lean [772]. Learn [29]. Learn-and-Optimize [29]. Learner [227]. Learning [629, 742, 33, 515, 189, 919, 191, 195, 184, 614, 183, 819, 198, 186, 199, 200, 288, 852, 518, 616, 230, 201, 192, 885, 941, 197]. Least [733]. Lecture [838]. Left [813]. Legal [780]. Legislative [470]. LEGO [512]. lemon [472]. Length [256]. Lesion [532]. Leuven [934]. Level [890, 713, 531, 535, 715, 91, 690]. Levels [520]. Leveraging [706]. Library [699, 469, 692]. Lifetime [92]. Ligand [156]. Light [448]. Lightweight [751]. Like [596, 518]. Limits [63]. Line [601, 23]. Linear [300, 747, 584, 70, 384, 210, 144, 391, 485, 139, 698, 509, 693, 56]. Linear-Rate [210]. Linearization [506]. Link [841, 326]. Linked [795, 326]. Linking [866]. LIPM [284]. List [720]. Literature [768]. Lithography [687]. Liver [532]. Livestock [159]. Load [38]. Load-Balanced [38]. Local [796, 24, 510, 4, 762, 52, 769, 776, 596, 290, 777, 426, 559, 21]. Local-Repair [426]. Locales [120]. Localization [294, 345, 324]. Location [134, 333, 679, 229]. Locks [650]. LocoKit [441]. Log [849]. Logging [349]. Logic [682, 657, 266, 371, 577, 663, 677, 923, 120]. Logic-Based [371, 663]. Logical [577]. Logs [195]. London [776]. Lonely [143]. Long [650, 750]. Loop [540, 632, 433]. Lossless [848]. Lossy [208]. Lotka [757]. Low [544, 718]. Low-Resolution [544]. Lower [814, 600, 89]. Lowerbounding [72]. LPs [134]. LTE [434]. LTL [635]. Lumbar [556]. LuSH [357].

MA [916]. MAC [424, 322]. Machine [451, 814, 819, 461, 691, 852, 616, 228, 230]. Machine-Learning [852]. Machines [133, 727, 816, 553, 820, 660]. Macroparasites [560]. Magnetic [759]. Maintaining [709, 706]. Maintenance [786]. Majority [205, 304]. Making [800]. Malware [344]. Mammograms [558, 619]. Mammography [616]. Management [927, 904, 926, 346, 332, 780, 783]. Managing [820]. Manipulating [643]. Manipulation [515, 444, 280]. Manipulator [510]. Manufacturing [657, 879]. Many [429]. Many-to-Many [429]. Map [436, 201, 49]. Mapping [537, 500]. MapReduce [881, 355, 257, 723]. Maps [544]. Maribor [948]. Marketability [887]. Markets [521]. Markov [562]. Masking [481, 478]. Masks [458]. Mass [284, 758, 558, 76]. Mass-Action [758]. Massachusetts [403]. Massive [337]. Massively [722]. Matching [55, 544, 815, 141, 39, 236]. material [538]. Mathematical [751]. Matrix [392, 756, 722, 883, 254]. Matter [15, 36, 77, 160, 215, 274, 375, 473, 527, 582, 661, 803, 16, 37, 528, 549, 570, 583, 604, 625, 646, 662, 683, 704, 725, 746, 767, 788, 804, 825, 846, 867, 888, 909, 58, 78, 100, 121, 142, 161, 182, 203, 216, 237, 258, 275, 296, 317, 338, 359, 376, 397, 418, 439, 460, 474, 495, 516]. Matting [564]. Max [136, 592, 905, 393]. Max-Cut [393]. Max-Min [592]. Max-Not-2 [136]. Max/Min [905]. Maximal [632, 224, 597]. Maximization [128]. Maximizing [92]. Maximum [586, 595, 141, 39]. May [901]. Maze [623]. McEliece [208]. MDSP [927]. Me [730]. Mean [621]. Meaning [515]. Means [585, 219]. Measles [636]. Measurement [700]. Measures [613, 437, 243]. Mechanism [350, 85]. Mechanisms [820, 6]. Media [802, 789, 197]. Median [584]. Medical [530]. Medium [816, 791]. Membership [111]. Memetic [27]. Memory [709, 67, 691, 711, 706, 490]. Memory-Based [490]. Meta [548]. Metabolic [459]. Metadata [837]. Method [452, 742, 749, 231, 421, 538, 430, 234, 718, 431, 246, 248, 232, 229, 53]. Methodologies [787]. Methodology [194]. Methods [540, 481, 627, 942, 362, 308, 146, 943]. MetiTarski [101]. Metric [64, 134, 657, 550, 65, 68, 66, 839, 63]. MEXIR [413]. MGS [811].

Microblog [235]. Microcontroller [651]. Mid [831]. Mid-Tier [831]. Middleware [865]. MIG [351]. Military [476]. MIMO [93]. Min [592]. MINDSTORMS [512]. Miner [606]. Minimal [653, 595, 575, 813, 459]. Minimizing [179]. Minimum [588, 131, 41, 603, 392, 429, 139, 90, 381, 598]. Minimum-Cost [131]. Mining [320, 878, 319, 437, 883, 884, 325, 723, 326, 886]. Minmax [584]. minor [591]. miRNA [151]. Missing [221]. Mitigate [890, 893]. Mixed [385]. MLO [619]. Mobile [99, 191, 92, 202, 878, 184, 432, 707, 198, 80, 113, 287]. Mobility [280]. Modal [640, 267, 447, 334, 557]. Modalities [555]. Model [10, 366, 148, 872, 150, 505, 665, 623, 655, 497, 30, 656, 488, 898, 344, 235, 313, 372, 8, 677, 218, 548, 328, 438, 483, 273, 5, 797, 448, 785, 225, 580, 440, 799, 240, 223, 241]. Model-Based [656]. Model-Driven [872]. Modeling [365, 367, 12, 705, 869, 868, 774, 855, 199, 13, 518, 268, 521]. Modelling [827, 9, 447, 644, 4, 455, 373, 364]. Models [827, 648, 871, 380, 910, 562, 670, 896, 691, 381, 396]. Moderated [802]. Modification [152, 271]. Modified [554, 434]. Modular [360, 472, 657]. Modulation [435]. Modulo [765]. Molecular [148, 577, 158, 939, 156]. Monadic [112, 113]. Monitoring [635, 690]. Monitors [634, 370]. Monomials [758]. Montpellier [944]. Montreal [920]. Morphologies [441]. Morphology [534]. Most [333]. Motion [510, 563, 820, 542, 291]. Motions [536]. Motivation [186]. Motor [443, 449, 454, 617]. Motor-Primitive [449]. Mountains [914]. MRFM [353]. MSMapper [355]. Much [889]. Multi [688, 874, 749, 315, 588, 312, 231, 82, 705, 554, 715, 447, 538, 609, 304, 307, 31, 211, 519, 371, 614, 713, 294, 39, 181, 425, 334, 218, 541, 91, 738, 363, 690, 557, 741, 500, 692, 291, 695, 85, 269, 83, 240, 885, 356]. Multi-agent [519]. Multi-Block [738]. Multi-channel [82]. Multi-Class [741, 609]. Multi-Context [269]. Multi-core [363, 695]. Multi-dimensional [874]. Multi-DOF [500]. Multi-GPU [692]. Multi-grain [688]. Multi-granular [231]. Multi-Granulation [240, 218]. Multi-head [181]. Multi-hop [82, 83]. Multi-hypotheses [294]. Multi-Instance [614, 315]. Multi-Interface [39]. Multi-level [715, 91, 690]. Multi-material [538]. Multi-modal [447, 334, 557]. Multi-objective [31, 885]. Multi-Parameter [749]. Multi-Party [304, 211, 307]. Multi-processors [713]. Multi-prover [312]. Multi-radio [85]. Multi-robot [291]. Multi-scale [554]. Multi-sensor [356]. Multi-threaded [371]. Multi-transaction [705]. Multi-user [425]. Multi-view [541]. Multicast [98, 421]. Multiclass [621]. Multicore [724, 696]. Multicut [48]. Multidimensional [871, 875, 263, 352]. Multigrid [535]. Multihop [96]. Multilayer [621]. Multilevel [26]. Multilingual [472, 468]. Multimedia [863, 434]. Multimodal [453, 280]. Multiobjective [18, 620]. Multiparty [205]. Multiplayer [706]. Multiple [387, 395, 875, 542]. Multiplicative [207]. Multiplicities [763]. Multiply [722]. Multiprocessor [693]. Multisensory [32]. Multithreaded [722]. Multivariate [73, 484]. Multivoxel [607]. Multiway [51]. Municipal [791].

Must [889, 303]. Mutual [557]. MWSNs [344]. Myth [491].

Named [236]. Narrow [535]. NAS [350]. Natal [941]. Natural [951, 462, 466, 468, 933, 301, 465, 463, 440]. Nature [737]. Navigation [503, 453]. Near [124]. Near-Perfect [124]. Nearest [341, 68, 612]. Nearest-Neighbor [341]. Need [800, 772]. Negative [330]. Neighbor [341, 68, 612]. Neighborhoods [224]. NEO [9]. NEON [493]. Nerve [343]. Nested [745]. NestedCube [352]. Net [10]. Nets [3, 910, 4, 8, 7, 14]. Network [126, 130, 866, 92, 386, 889, 519, 432, 423, 80, 686, 341, 346, 711, 343, 428, 345, 620, 396, 459, 94, 84, 435, 347, 356, 352, 197]. Network-Flow [396]. Networks [97, 593, 98, 573, 710, 82, 574, 754, 816, 319, 38, 429, 685, 430, 424, 39, 40, 427, 425, 423, 90, 323, 80, 301, 420, 321, 86, 438, 426, 89, 322, 824, 21, 95, 81, 85, 83, 88, 96, 324, 87]. Neural [877, 447, 32, 620, 454]. Neural-Based [877]. Neuro [458]. Neuro-Control [458]. Neurocomputing [249]. Neuroevolution [457]. NFS [702]. NJ [915]. NN [61]. NoC [615]. Node [126, 130, 92, 345, 824]. Node-Localization [345]. Node-Weighted [126, 130]. Nodes [395]. Non [368, 697, 802, 657, 386, 485, 548, 173, 903, 509, 83, 87]. Non-cooperative [83]. Non-deterministic [697]. Non-Disjoint [386]. Non-erasing [173]. Non-intrusive [87]. Non-Linear [509]. Non-profiled [485]. Non-Repudiation [903]. Non-rigid [548]. Non-Standard [657]. Non-termination [368]. Nondeterminism [165]. Nondeterministic [180]. Nonlinear [433, 686, 608]. Nonnegative [389]. Nonparametric [86]. Nontermination [118]. Norm [125, 240]. Normal [752]. Normalization [552]. Normalized [741, 551]. Norway [949, 950]. Note [720, 764]. Notes [838]. Novel [488, 562, 498, 83]. NP [122, 806, 259, 177, 136]. NP-Complete [177]. NP-Hardness [122, 136]. Nuclei [561]. NullPointerExceptions [368]. Number [595, 603, 43]. Numbers [601, 106, 105, 26]. Numerical [605, 759].

Object [360, 928, 367, 713, 289, 287, 659, 280, 364, 220]. Object-Based [360]. Object-Level [713]. Object-Oriented [928, 367]. objective [31, 885]. Objects [448]. Oblivious [208, 210]. OBS [430]. Observer [617]. Observing [284]. Obstacle [288]. Occluded [736]. Occurrence [135]. OCL [871]. OCR [405]. OCR-ed [405]. October [928, 911]. Odd [264]. Odense [932]. Off [148, 127]. Off-Lattice [148]. Offline [292]. Offshore [656]. OLAP [877]. On-Board [23]. On-Line [23]. One [123, 749, 494]. One-Shot [123]. Ones [598]. Online [796, 189, 129, 133, 589, 800, 334, 793, 291, 352, 87, 197]. onto [25]. Ontological [801]. Ontologies [472, 668]. Ontology [856, 664, 839, 464]. Ontology-Based [856, 664]. Open [866, 795, 785]. Opening [790, 775]. Operated [501]. Operating [10, 715, 730]. Operation [231]. Operations [656, 577]. Operator [18, 103]. Opinion [327]. Opportunistic [618, 86]. Optical [708]. Optics [823]. Optima [21]. Optimal [720, 561, 576, 336, 691, 684, 726, 693, 95, 594]. Optimistic [240]. Optimization [605, 256, 18, 624, 870, 916, 912, 686, 929, 17, 379, 35, 687, 88, 885]. Optimizations [226]. Optimize [383, 29]. Optimized [485, 558]. Optimizer [31]. Optimizers [836]. Optimizing [658]. Optimum [377, 214]. or/and [71]. Oracles [164]. Orbital [382]. Order [602, 384, 385, 645, 509, 726, 830]. Order-Based [830]. Ordering [135, 139]. Ordinal [222]. Ore [756]. Oriented [928, 367, 319, 858, 721, 364]. Orléan [951]. Other [910, 585]. OUC [406]. Outbound [908]. Outcomes [783]. Outerplanar [597]. Outflow [893]. Outline [217]. Outsourced [329]. Overhead [129, 696]. Overview [398, 414, 407]. OWL [682, 681, 668, 464, 866].

P [806, 163, 812, 813]. PA [774]. Packaging [337]. Packet [615, 434, 88]. Packing [593]. Page [401]. Palytoxin [150]. Pan [545]. Pan-Tilt [545]. Papers [937, 938, 928, 930, 911, 923, 929]. Paradoxes [272]. Parallel [720, 252, 881, 718, 686, 346, 339, 79, 69, 691, 158, 700, 7, 945, 946, 723]. Parallelism [688, 713]. Parallelized [257]. Parameter [749, 29]. Parameterized [49]. Parameters [822]. Parametric [726]. Parasites [534]. Parent [202]. Paris [942, 943]. Parsing [146]. Part [937, 938, 952, 953, 945, 946]. Partial [747, 694, 733, 506, 548]. Partially [736]. Participate [798]. Participation [796, 789, 403, 406, 950]. Participatory [797]. Particle [624, 31, 556, 621]. Partitioned [349]. Partitioning [862, 863, 428]. Partitions [603]. Party [304, 211, 307]. Pass [235, 730]. Passes [141]. Passive [211]. Password [315]. Password-Based [315]. Past [101]. Path [710, 50]. Paths [593, 45, 283]. Patricia* [884]. Pattern [947, 632, 607, 174, 731, 726, 293]. Pattern-Based [632]. Patterned [575]. Patterns [647, 575, 587, 843, 326]. PCA [619, 351]. PCA-MIG [351]. PCM [833]. PCM-Aware [833]. Pebbling [123]. [546]. and [71]. CA [347]. CDMA [431]. K [150]. Min [905]. SONET [386]. PEPA [636]. Perception [446]. Perceptions [781]. Perceptron [621]. Perfect [124, 601]. Performance [24, 735, 834, 898, 864, 707, 703, 722, 13, 325, 426, 706, 700]. Permutation [301, 79, 69, 43]. Permutation-Based [79, 69]. Permutations [143, 316]. Persistence [567]. Person [567]. Personal [855]. Personalized [856]. Personalizing [679]. Perspective [626, 891, 792, 319, 80, 200, 769]. Perspectives [198, 524]. Petitioning [792]. Petri [910, 10, 3, 4, 8, 7]. Phase [534, 655, 718, 442, 617]. Phenomena [524]. Phenotypic [805, 560]. Phonetic [261]. Photonic [477]. Phrases [270]. Phylogenetic [157]. Phylogeny [124]. Physically [491]. Piecewise [698]. Pivot [61]. Place [453]. Placement [824]. Plan [836]. Planar [126, 588, 45, 820]. Plane [393, 442]. Planes [627]. Planning [832, 29, 515, 510, 870, 450, 513, 283, 292, 291]. Plans [656]. Plasma [700]. Plastic [456]. Platform [284, 615, 793]. Plattform [512]. Play [28]. Player [489]. Plus [623]. Plus-Maze [623]. Pneumoperitoneum [529]. pnml [3]. Pocket [81]. POI [336]. Point [122, 541]. Pointer [643]. Points [902]. Poisson [752]. Polaris [324]. Polarities [887]. Policies [895, 787, 679]. Policy [794, 802, 901, 183, 800, 799]. Policy-Making [800]. PolicyCommons [794]. Pollution [28]. Polygon [684]. Polygons [603]. Polyhedral [386, 388]. Polynomial [747, 756, 762, 766, 396]. Polynomial-Time [756]. Polynomials [756]. Polyphasic [63]. Polytopes [391]. Populating [328]. Population [24, 22, 839, 608]. Port [529]. Portable [295]. Portrait [442]. Ports [25]. Portuguese [471]. Pose [543, 733]. Position [259, 762]. Positioning [529]. Positions [753]. Positive [327, 330]. Potential [503]. Power [869, 519, 431, 90, 35, 524, 93]. Powered [459]. Powering [766]. Powers [750, 166]. Practical [189, 487, 362, 489, 63]. Practice [790, 318, 193, 480]. Pragmatic [676]. Pre [433]. Pre-distortion [433]. Precision [852]. Predication [270]. Predicting [259]. Prediction [877, 152, 450, 151, 852, 251, 840, 849]. Predicts [448]. Preemption [129]. Preference [886]. Preferences [842, 675, 195]. Prefix [691]. Prefix-Sums [691]. Preliminary [187]. Presence [129]. Preserving [305, 436, 609, 176, 341, 607, 557]. Pressures [769]. PRFs [486]. Price [378]. Pricing [128]. Primal [126]. Primal-Dual [126]. Primary [193, 94]. Primitive [449]. Princeton [915]. Principal [733, 59]. Principles [775]. Prior [561]. Priority [370, 347, 115]. Priority-Based [347]. Privacy [205, 955, 902, 341, 894, 11]. Private [786, 186]. Prize [130]. Prize-Collecting [130]. Probabilistic [367, 150, 221, 225, 579]. Problem [749, 387, 753, 380, 806, 1, 54, 710, 640, 41, 55, 386, 393, 757, 395, 589, 137, 171, 880, 388, 51, 68, 809, 146, 17, 381, 27, 598, 606, 49, 50, 44]. Problems [631, 123, 128, 394, 134, 139, 585, 140, 686, 526, 620]. Procedure [893]. Procedures [110]. Proceedings [935, 952, 953, 945, 946, 947, 927, 915, 944, 919, 954, 951, 955, 926, 948, 942, 940, 916, 933, 931, 922, 912, 913, 934, 925, 949, 920, 939, 943, 950, 914, 918, 941, 932, 917]. Process [790, 148, 636, 666, 670, 797, 893, 192]. Processes [871, 868, 774, 873, 233]. Processing [674, 447, 341, 339, 463, 828, 945, 946, 352]. Processor [722]. Processors [724, 713]. Producing [871]. Production [672, 667]. Prof. [162]. profiled [485]. Profiles [732, 153, 886]. Profiling [87]. Profit [128]. Profitability [904]. Program [654, 385, 201]. Programming [682, 384, 684, 591, 692, 463, 939, 693]. Programs [649, 360, 384, 264, 371, 112, 113, 690, 374, 667, 7]. Progress [947]. Promoting [99]. Proof [119, 111, 308, 85]. Proof-Relevant [111]. Proofs [102, 312, 639, 299, 301, 118]. Propagation [344]. Properties [681, 641, 57, 154, 374, 660]. Property [827, 598]. Proportional [434]. Proposal [790]. Proposals [799]. Proposed [893]. Protecting [685, 905]. Protein [148]. ProteinWorldDB [155]. Protocol [350, 10, 9, 432, 424, 427, 323, 426, 322, 115, 87]. Protocols [206, 207, 389, 212, 489, 11]. Prototype [895]. Provably [114, 481]. Prove [403]. Proved [115]. prover [312]. Providers [194]. Proving [915].

Provision [907]. Proximity [687, 63]. Pseudo [712, 396]. Pseudo-polynomial [396]. Pseudoknotted [53]. Pseudorandom [301]. PTA [202]. PTAS [585]. PTL [349]. Public [802, 773, 782, 786, 781, 772, 785, 156, 783]. Publication [236]. PUF [492, 489]. PUF-Based [492, 489]. PUFKY [492]. PUFs [491, 491, 490]. Puppy [445]. Pursuers [503]. Putting [676]. Puzzle [268]. Pyramid [714].

QA [415, 417, 414]. QAEE [424]. QC [920]. QoI [348]. QoI-Based [348]. QoS [424]. Quad [428]. Quad-Tree [428]. Quadruped [445]. Quadruplexes [147]. Quadtree [689]. Quality [318, 550, 91, 712, 743, 330]. Quality-Metric [550]. Quantification [530]. Quantified [634]. Quantifiers [262]. Quantitative [857]. Quantum [164]. Quasi [311, 760, 20]. Quasi-Random [20]. Quasi-stability [760]. quaternary [251]. Queen [880]. Queen-Bee [880]. Queries [842, 835, 329, 341]. Query [841, 847, 336, 835, 880, 664, 840, 905, 331, 828, 332]. Querying [682, 462, 845, 848, 855]. Question [416, 414]. Queues [703]. Quick [242]. Quorum [719]. Quotient [167].

Radial [545]. Radiation [49]. Radii [710]. Radio [436, 698, 95, 94, 435, 85]. Rainbow [748, 599, 600]. Ramsey [601]. RANDOM [916, 151, 712, 20]. Randomization [916]. Randomized [736, 389]. Range [653, 687, 332]. Rank [553]. Ranked [263]. Ranking [399, 857]. RANSAC [546]. Rapid [848]. Rate [708, 210, 422, 509]. Rates [759]. Ratings [330]. Ratio [392, 145]. Rats [623]. Ray [538]. Ray-Reps [538]. Rays [392]. RBF [620]. RDBMS [226]. RDF [462, 354]. Reachability [574]. Reachable [456, 291]. Reaching [443, 496, 63]. Reaction [573, 574, 754, 344, 671, 671]. Reaction-Diffusion [344]. Reactive [663, 673, 658]. Reactivity [675]. Real [647, 537, 536, 106, 105, 430, 453, 864, 724, 686, 692, 292, 293]. Real-Time [647, 537, 536, 430, 864, 724, 686, 692, 292]. Realistic [286]. Reals [105]. Rearrangements [144]. Reason [834]. Reasoning [682, 649, 260, 808, 672, 579]. Receipt [644]. Receiver [433]. Recognition [947, 736, 878, 737, 550, 733, 466, 563, 249, 562, 728, 731, 551, 611, 555]. Recognizing [821]. Recommendation [856, 402]. Reconfigurable [494]. Reconfiguration [694]. Reconfigurations [641]. Reconstruction [158, 539, 729, 535, 356]. Recovering [164]. Recruitment [893]. Rectilinear [603]. Recurrence [53]. Recursive [411, 714]. Red [598]. Red-Blue [598]. Reducing [33]. Reduct [242]. Reduction [64, 82, 245]. Reducts [252]. Redundant [449]. Reengineering [774]. Reference [797]. Refined [219]. Refinement [629, 107, 641, 112]. Refinement-Based [107]. Reflexes [497]. Refraction [537, 667]. Refraction-Based [667]. Refreshing [864]. REGEX [169]. Region [149, 65]. Regional [795]. Registration [546, 557]. Regression [542, 620]. Regret [584]. Regular [119, 168, 167, 719, 170, 578, 765, 174, 310]. Regulation [901, 902]. Regulatory [669, 6, 459]. Related [123, 133]. Relation [111, 239, 241].

Relational [847, 853, 331, 849]. Relations [207, 851]. Relationship [19]. Relative [256]. Relaxation [383]. Relaxations [385]. Release [213]. Relevance [841]. Relevant [111, 266]. Reliability [519]. Reliable [97, 430, 824]. Reliably [448]. Relocation [694]. Rely [120]. Rely-Guarantee [120]. Remote [137]. Remotely [501]. Removal [589]. Repair [426]. Repairing [299]. Repeats [146]. Replication [140]. Report [8]. Reports [328]. Representation [456]. Representations [566, 568, 817, 738]. Reps [538]. Repudiation [903]. Reputation [677]. Requests [273]. Requirements [873]. Rescue [283]. Research [944, 189, 188, 193, 768, 192, 775]. Researching [520]. Reservation [421]. Reshaping [612]. Residual [322]. Resiliency [685]. Resilient [487, 486]. Resistance [644]. Resizing [608]. Resolution [544]. Resolutions [269]. Resonant [757]. Resource [697, 469, 372, 140, 419, 346, 56]. Resource-Restricted [372]. Resources [402, 372]. Restarts [20]. Restricted [753, 727, 372, 592]. Restrictions [681]. Results [128, 44, 386, 335, 140]. Resurgence [404]. Rethymnon [937, 938]. Retrieval [930, 861, 408, 548, 75]. Reuse [830]. Revealing [134]. Reverse [665, 341, 438]. Review [790, 198, 268, 768]. Revised [937, 938, 928, 930, 911, 929]. Revision [342]. Revisited [209]. Revisiting [72]. Rewrite [630]. RF [435]. RF-CBTC [435]. RFMirTarget [151]. Right [813, 791]. Rigid [286, 548]. Rigorous [20]. Ring [386]. Riots [776]. Risk [656]. Risks [893]. RL [682]. RNA [53, 147]. RNN [438]. Road [341]. RoboCup [276, 284, 289, 924, 279]. RoboCupJunior [281]. Robot [456, 445, 514, 453, 506, 441, 498, 924, 501, 292, 277, 730, 293, 291, 505]. Robotic [510, 285, 507]. Robotics [508, 512, 935, 23]. Robots [449, 287, 283, 500, 280, 511, 507, 730]. RoboWorld [935]. Robust [114, 546, 294, 287, 542, 280]. Robustness [631, 205, 740, 658]. ROLAP [881]. Room [730]. Root [395, 762]. Rotation [92]. Rough [247, 252, 253, 249, 218, 246, 227, 251, 243, 226, 225, 219, 238, 240, 223, 257, 220, 241, 922, 921]. Route [336]. Routing [82, 348, 429, 615, 432, 427, 323, 81]. RSCTC [921]. RSKT [922]. RSL [244]. RSLIS [400]. RTDW [864]. RTDW-bench [864]. RUC [412]. Rule [681, 674, 221, 675, 677, 246, 227, 670, 248, 672, 667, 255, 679]. Rule-Based [681, 674, 675, 677, 672, 679]. RuleML [944, 671]. Rules [256, 222, 665, 668, 669, 613, 470, 678, 673, 664, 671, 863, 255, 676, 849, 886, 944]. Runs [410, 14]. Runtime [634, 20]. Russians [53]. Rydberg [759].

S [479, 712]. S-Boxes [479, 712]. Saarbrücken [930]. SAB [932]. Safe [656]. Salesman [387, 17]. sameAs [866]. Sample [242]. Samples [484]. Sampling [127, 510, 585, 542]. SAN [13]. SAN-Based [13]. Sand [821]. Sandboxing [908]. Sandpile [821]. Santa [925]. Satisfaction [200]. SCA [261]. Scaffolding [190]. Scalability [703]. Scalable [720, 68, 829, 242]. Scalar [700]. Scalar-Type [700]. Scale [354, 862, 62, 554]. Scales [454]. Scaling [586]. Scanning [476]. SCAR [567]. Scattered [175]. Scatterplots [22]. Scenarios [372, 289]. Scheduler [371]. Scheduler-Specific [371]. Scheduling [631, 129, 133, 880, 724, 322, 434, 693, 695, 96, 56]. Schema [845, 844]. Scheme [900, 890, 744, 31, 297, 425, 354, 17, 829, 355, 428, 94]. Schemes [306, 830]. School [193]. Schur [26]. Schützenberger [173]. Sciences [159]. Scientific [948, 158]. Scissors [730]. SCOOTER [829]. Scope [3]. Screening [534, 560]. SDH [386]. SDH/SONET [386]. SDWM [869]. Search [399, 64, 24, 400, 67, 402, 859, 335, 378, 398, 334, 71, 68, 913, 283, 76, 409, 26, 410, 612, 62, 559, 860, 459, 331, 358]. Searches [61]. Searching [70, 63, 60]. Second [384, 385, 929]. Second-Order [384, 385]. Secondary [67, 53]. Secrecy [204]. Secret [306, 214]. Sector [782, 786, 198, 781, 772]. Secure [99, 311, 481, 210, 304, 907]. Securely [303]. Security [626, 315, 314, 955, 889, 685, 299, 313, 491, 298, 308, 486, 316, 489, 906, 903, 920, 213]. See [569]. Segmentation [534, 533, 616, 531, 621]. Selected [937, 938, 928, 930, 911, 923, 929]. Selecting [484]. Selection [619, 351, 438, 333, 254]. Selective [688, 308]. Self [575, 810, 501, 580]. Self-Assembly [575, 580, 810]. Self-triggered [501]. Semantic [838, 314, 680, 234, 468, 199, 850, 671, 268]. Semantics [680, 703, 673, 271]. Semi [742, 838, 141, 556]. Semi-automatic [556]. Semi-streaming [141]. Semi-structured [838]. Semi-Supervised [742]. Semidefinite [385]. Semigroups [171]. Sense [437]. Sensing [97]. Sensitive [905]. Sensor [710, 618, 865, 38, 429, 685, 424, 40, 427, 423, 323, 295, 420, 321, 513, 343, 428, 89, 345, 322, 824, 88, 96, 347, 87, 356]. Sensor-Based [513]. Sensorimotor [445, 450]. Sentiment [861]. Separability [238]. Separation [892, 394]. September [947, 954, 951, 955, 948, 952, 953, 934, 949, 950, 945, 946]. Sequence [520, 154]. Sequenced [336]. Sequences [712, 723]. Sequential [667]. Server [908, 702]. Server-Driven [908]. Service [421, 773, 319, 903, 280]. Service-Oriented [319]. Services [827, 904, 91, 777, 192, 907]. Servo [282]. Servoing [509]. Session [890, 423]. Session-Based [890]. Sessions [923]. Set [247, 264, 137, 218, 90, 251, 463, 225, 598, 240, 223, 257, 535, 220, 241, 594]. Sets [224, 595, 575, 252, 922, 246, 227, 243, 517, 291, 46, 531, 921, 220]. Setting [485]. Shannon [116, 209]. Shape [561, 543, 548, 289]. Shape-Based [289]. Shared [711, 777, 524]. Sharing [306, 901, 95, 192, 214]. Sheng [162]. Sherali [383]. SHIELD [85]. Ship [536]. Shopping [334]. Short [900, 687]. Short-Range [687]. Shot [123]. Should [480]. Shrinking [382]. Shuffle [297]. Side [809, 486, 483]. Side-Channel [486, 483]. SIFT [552]. Signature [900]. Signatures [71]. Signed [729]. Significance [196]. Silence [187]. Silicon [491, 476]. Similarity [64, 67, 74, 548, 52, 530, 882, 76, 243, 62, 60, 913]. Simple [742, 585, 477, 511]. Simplified [455, 464]. SimTandem [76]. Simulating [148]. Simulation [276, 529, 536, 701, 519, 525, 286, 932]. Simulink [652]. Single [588, 543, 235, 490]. Single-Pass [235]. Single-Source-Multi-Sink [588]. Sink [588]. SIR [75]. SISAP [913]. Sites [152]. situ [848]. Situation [840]. Situation-Aware [840]. Size [598, 88]. Sized [138]. Skating [514].

Skeletal [563]. Skeletons [699]. SKUBA [277]. Skyline [842, 329]. Slicer [630]. Slicing [660]. Slotted [347]. Slovenia [948]. Small [337, 180, 721, 174]. Smart [771, 618, 75]. Smooth [729]. Smoothing [436, 541]. Snake [61]. Soccer [285, 924, 292]. Social [400, 802, 82, 789, 402, 183, 398, 198, 894, 409, 524, 197]. Soft [243, 286, 490]. Software [626, 709, 928, 404, 739, 183, 522, 658, 277]. Solid [349]. Solids [538]. Solution [443, 350, 749, 272]. Solutions [747, 693]. Solver [637]. Solving [710, 758, 526, 146]. Some [750, 587, 408]. Sort [833]. Sorting [145]. Sound [261]. Source [588, 770]. Space [456, 764, 282, 581, 700]. Spaces [64, 73, 855, 68, 809, 62, 860]. Spanning [429]. Sparse [566, 766]. Spatial [753, 869, 344, 329, 353, 811]. Spatial-temporal [344]. Spatio [566, 859]. Spatio-temporal [566, 859]. Spatiotemporal [465]. Specific [371, 373, 6]. Specification [647]. Specifications [651]. Spectra [76]. Spectral [125, 537, 741]. Spectral-Norm [125]. Spectrometer [698]. Spectrum [95]. Speech [466, 611]. Speed [586, 708, 617]. Speedup [53]. Spinal [556]. Splicing [578]. Split [599, 355]. Splitting [1]. Spreadsheet [837]. Sprint [870]. SPSA [88]. SQUALL [462]. Squared [134]. Squares [733]. SRNs [12]. SSReflect [116]. Stabbing [603, 381]. Stability [753, 760]. Stable [283]. Stacks [387]. Stage [571]. Stakeholder [791]. Standalone [897]. Standard [657, 284, 313]. Standardized [671]. Star [386]. Stars [44]. Start [133]. Startup [655]. State [464, 660, 768, 349]. States [759, 181]. Static [615, 66, 711, 659]. Static-Based [711]. Static-to-Dynamic [66]. Station [427, 94]. Statistical [866, 488, 211, 248, 905]. Statistically [210]. Statistics [726]. Status [784]. Steiner [41, 395]. Step [571, 2]. Step-Wise [571]. Stepsize [20]. Stepwise [827]. Stereo [544]. Stop [117]. Storage [350, 898, 349]. Storing [848]. Strand [574, 581, 579]. Strategic [893]. Strategies [33]. Strategy [85]. Strategy-Proof [85]. Stream [482, 878, 861, 712, 690]. Stream-Based [690]. Streaming [421, 141]. Streams [61, 865]. Strengths [706]. Stretch [40]. String [815, 882]. Strings [365, 164]. Strong [759, 597]. Structural [641, 388]. Structure [377, 67, 53, 930, 404, 185, 405, 526]. Structured [340, 838]. Structures [66, 884]. Structuring [401, 795]. Stuck [892]. Student [202, 195, 923, 186, 200, 197]. Students [185]. Studied [580]. Study [872, 732, 189, 191, 190, 364, 702, 187, 197]. Studying [441]. Subdivision [540]. Subgraph [587, 340]. Subgraphs [596]. Sublinear [127]. Subset [598]. Substantially [587]. Substitution [301]. Substitution-Permutation [301]. Successive [431]. Succinct [312]. Suggestions [399]. Summarization [415, 417]. Sums [691]. Super [70, 486]. Super-Exponential [486]. Super-Linear [70]. Supercomputer [700]. Supervised [742]. Supply [128, 56]. Support [652, 247, 843, 800, 553, 461, 233, 228]. Supporting [838, 231, 656, 91]. SURF [622]. Surface [42, 729, 535]. Surgery [529]. Surrogate [18]. Surrogate-Based [18]. Surveillance [502]. Survey [798]. Survivable [130]. Swarm [624, 31, 621]. Sweep [688]. Switched [81]. Switching [481].

Switzerland [933]. Sydney [940]. Symbol [817]. Symbolic [515, 876, 759, 660]. Symbolic-Numerical [759]. Symmetric [390, 487]. Symmetry [383]. Symposium [937, 938, 944, 942, 929, 917]. Synchronized [172]. Synergies [443]. Synergy [660]. Syntax [271]. Synthesis [443, 651, 113]. Synthesising [449]. Synthesizing [575]. System [778, 415, 572, 719, 640, 451, 202, 757, 578, 249, 562, 861, 523, 417, 244, 346, 295, 518, 343, 706, 611, 228, 345, 434, 695, 459, 435, 356, 324, 332]. Systematic [134]. Systems [367, 897, 224, 705, 574, 812, 2, 715, 657, 813, 391, 755, 739, 758, 467, 663, 786, 431, 762, 952, 953, 91, 233, 658, 934, 35, 943, 581, 700, 440, 696, 914, 269, 364, 245, 93].

T [593, 131, 783]. T-Government [783]. T-Joins [131]. T-Paths [593]. Table [61, 680, 248]. Tables [874, 328]. Tabu [26]. Tag [860]. Taipei [918]. Taiwan [918]. Take [650]. Tandem [146, 76]. Tangent [820]. Tangent-Managing [820]. Tanzania [784]. Target [151]. TAROS [935]. Task [651, 416, 724, 444, 282, 696, 695]. Task-Space [282]. Tasks [443, 453, 724]. Teacher [202, 518]. Teaching [186, 197]. Team [277]. Technical [470]. Technique [467, 694, 420]. Techniques [72, 916, 308]. Technological [198]. Technologies [277]. Technology [778, 931, 922, 525]. Teenagers [798]. Teleoperation [500]. Telescope [698]. Temperature [571]. Template [565]. Template-Based [565]. Templates [899]. Temporal [681, 641, 657, 677, 861, 120, 849, 566, 859, 344]. Tensor [553]. Tensors [74]. termination [368]. Terrain [445]. Tessellation [540]. Test [547, 248]. Testing [638, 708, 645, 864, 5, 844]. Tests [898]. Tetris [33]. Text [334, 850, 358]. Text-Based [334]. Textbooks [320]. Texts [669, 839]. theft [343]. Their [312, 568, 485]. Theorem [915, 818, 180, 173]. Theorems [116, 377]. Theoretic [920, 213]. Theoretical [396]. Theories [630]. Theory [318, 761, 244, 480, 251, 257, 217, 918, 93]. Thermal [555]. Third [915, 933]. Though [808]. threaded [371]. Three [300, 394, 244, 146, 316, 217, 617]. Three-Index [394]. Three-Phase [617]. Three-Way [217]. Threshold [479, 599, 212, 219, 214]. Throughput [80]. Tier [831]. Tile [575, 810]. Tile-Based [810]. Tilt [545]. Time [647, 72, 584, 537, 33, 129, 536, 756, 645, 845, 430, 864, 724, 686, 484, 692, 292, 89, 454, 293]. Timed [651, 213]. Timed-Release [213]. Times [133]. TLA [639]. TLS [313]. TLS-DHE [313]. TOC [405]. Tolerance [239, 241]. Tolerant [502]. Tool [631, 652, 328, 659, 779]. Tools [647]. Top [835, 881, 333]. Top- [333, 835]. Top-Down [881]. Topic [235, 858]. Topic-Oriented [858]. Topological [851]. Topologies [812]. Topology [92, 707, 323, 321]. Topology-Aided [323]. Topology-Based [321]. Toronto [913]. Total [90]. Trace [630]. Track [411, 400, 403, 398, 408, 417, 406, 409, 414, 407, 412]. Tracking [532, 556, 560]. Tractional [820]. Trading [205]. Traffic [430, 434, 84]. Train [380, 435]. Training [194, 621]. Trains [627]. Trajectory [292]. Transaction [845, 705]. Transactional [709, 843, 706]. Transactions [910, 936]. Transcription [152]. Transfer [10, 208, 86].

Transform [554, 698]. Transformation [874, 66]. Transformers [110]. Transition [640]. Translation [461]. Transmission [10]. Transmit [431]. Transportation [731, 345]. Transposition [143]. Transpositions [145]. Traveling [387]. Travelling [17]. Tree [584, 41, 395, 429, 526, 428, 46, 885, 250, 67]. Trees [51]. Treewidth [123]. Trends [921, 60]. Triangular [762, 240]. Triangulation [684, 381]. triggered [501]. Trio [105]. Trip [859]. Trust [896, 273, 955]. TrustBus [955]. Trustworthy [904]. Truth [244, 295]. TSP [590]. Tumor [616]. Tumors [531]. Tuning [709, 29]. Turbo [718]. Turin [928]. Turing [571]. Tutorial [104, 518]. Tutorial- [518]. Tutoring [523]. Tweakable [298]. Tweets [859]. Twitter [776]. Two [206, 720, 165, 25, 821, 489, 225, 598, 223]. Two-dimensional [821]. Two-List [720]. Two-Player [489]. Two-Way [165]. Type [572, 700, 621]. Types [602, 3, 113, 363].

Ubiquitous [99, 878]. UCNC [951]. UK [935]. Ultrasound [701]. Unary [177, 181]. Uncapacitated [387]. Uncertain [853]. Uncertainty [530]. Unclonable [491]. Unconstrained [567]. Unconventional [163, 951]. Undeniable [900]. Undergraduate [512]. Understanding [771, 260, 782, 781, 524]. Unification [517]. Unified [485]. Unimodality [547]. Unit [380]. Universal [311]. Universality [571, 809]. Universally [208]. Universes [225, 223]. University [403, 404, 183, 410]. Unlimited [128]. Unobtrusive [99]. Unscented [294]. Unsupervised [851, 839]. Updates [829]. Updating [462, 220]. UPF [409]. Upper [600, 659, 245]. Urban [283]. USA [915, 916, 925]. Usage [895]. USDM [927]. Use [191, 795]. Used [776]. User [399, 890, 859, 835, 678, 372, 466, 212, 840, 539, 849, 886, 425]. User-Centric [212]. User-Defined [372]. User-Generated [399]. User-Level [890]. Users [668]. Using [796, 540, 534, 871, 854, 566, 148, 636, 832, 622, 610, 472, 561, 847, 401, 680, 543, 744, 92, 754, 624, 546, 145, 878, 701, 335, 645, 445, 519, 336, 558, 678, 615, 843, 881, 431, 548, 526, 120, 157, 248, 450, 607, 288, 113, 79, 69, 766, 556, 698, 712, 438, 233, 158, 518, 383, 608, 884, 442, 542, 458, 863, 5, 687, 428, 801, 228, 330, 525, 726, 824, 254, 291, 531, 7, 364, 435, 783, 490]. Usually [650]. UT [276].

Vacuum [507]. Vague [270]. Validity [844]. Value [783]. Valued [244, 239, 223]. Values [221, 330, 905]. VANETs [348]. Variable [710, 138]. Variable-Sized [138]. Variants [173, 458]. Variation [18, 220]. Variety [577]. Various [700]. Vector [73, 289, 228, 830]. Vectorized [689]. Vehicle [633, 499]. Vehicles [442]. Vein [552]. Verbal [230]. Verbalisation [472]. Verification [827, 360, 652, 928, 366, 361, 231, 9, 638, 654, 657, 371, 906, 374, 11, 87]. Verified [699, 643, 108, 370]. Verifying [110, 7]. Versioning [709]. versus [806, 163, 760]. Vertebrae [556]. Vertex [595, 46]. Vertical [863]. Very [306]. VI [910]. via [24, 111, 307, 614, 186, 463, 535, 93, 87]. Video [861, 556]. Video-Fluoroscopy [556]. Videos [567, 565]. Vienna [954, 955, 952, 953]. View [542, 84, 541]. Views [195, 619]. VIII [936]. Villa [276]. Virtual [618, 38]. Virtualization [346]. Visible [889]. Vision [947, 560]. Vision-Based [560]. Visited [336]. Visual [937, 938, 453, 71, 22, 288, 509, 555]. Visualisation [793]. Visualization [188, 861, 232]. Visualizing [794, 185]. Visually [444]. Vlasov [700]. Voice [425]. Volterra [757]. Volumes [563]. vs [165]. Vulnerabilities [890]. Vulnerability [739].

WAIM [927, 926]. Waiting [650]. Walk [551, 286]. Walking [446, 445, 455, 293]. Wall [529]. Wandering [499]. Warehouse [872, 870, 864, 328, 906]. Warehousing [954, 869, 879]. Warping [72]. WASA [914]. Watermarking [114]. Wavelength [814]. Wavelength-Based [814]. Waves [536]. Way [165, 217]. Weak [26]. Weakly [452]. Wearable [451]. Web [944, 676, 827, 874, 927, 890, 891, 926, 319, 889, 908, 894, 839, 795, 849]. Web-Age [927, 926]. Web-Application [908]. Web-Based [319]. WebSand [908]. Weight [483]. Weighted [126, 588, 130, 839]. Westervelt [701]. WG [949, 950]. Where [362]. Whether [578]. While [706, 511]. Whispering [678]. Who [628]. Wi [436]. Wideband [698]. Width [48, 83]. Wiki [468, 190]. Wiki-Based [190]. Wikipedia [335, 330, 342]. Wild [627]. Winners [279]. Wireless [350, 98, 710, 38, 429, 685, 424, 40, 427, 423, 90, 323, 420, 321, 343, 345, 322, 824, 95, 85, 83, 88, 96, 324, 87, 914]. Wiretap [314]. Wise [571]. Withdrawal [497]. within [512, 680, 769]. without [145, 419, 374]. WK [714]. WK-Recursive [714]. Word [171]. WordNet [335]. Words [175]. Workflow [158, 149, 14]. Workflows [697]. Workload [862]. Workload-Based [862]. Workplace [199]. Workshop [948, 930, 916, 933, 934, 943]. Workshops [927]. Workspace [499]. World [99, 924]. Would [737]. WP [110]. WrightEagle [276]. Write [833]. Writer [568]. Writing [470]. WSN [419].

XACML [895]. XML [930, 834, 408, 829, 828, 830]. XMLDM [927]. XPath [835]. XQuery [836, 339]. XV [924].

Yao [57]. Yellow [914]. Yours [253]. Youth [798]. Yu [162].

Z [149]. Z-DNA [149]. Zero [207, 212, 762, 765, 43]. Zero-Dimensional [762, 765]. Zero-Knowledge [212]. Zoom [545]. Zurich [933].

References

Carmona:2012:LSP

[1] Josep Carmona. The label splitting problem. Lecture Notes in Computer Science, 7400:1–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35179-2_1/.

Darondeau:2012:DCD

[2] Philippe Darondeau and Laurie Ricker. Distributed control of discreteevent systems: a first step. Lecture Notes in Computer Science, 7400: 24–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35179-2_2/.

Hillah:2012:EPS

[3] Lom-Messan Hillah, Fabrice Kordon, Charles Lakos, and Laure Petrucci. Extending pnml scope: a framework to combine Petri nets types. Lecture Notes in Computer Science, 7400:46–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35179-2_3/.

Kindler:2012:MLG

[4] Ekkart Kindler. Modelling local and global behaviour: Petri nets and event coordination. Lecture Notes in Computer Science, 7400:71–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35179-2_4/.

Salem:2012:MCU

[5] Ala-Eddine Ben Salem, Alexandre Duret-Lutz, and Fabrice Kordon. Model checking using generalized testing automata. Lecture Notes in Computer Science, 7400:94–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35179-2_5/.

Sedlmajer:2012:DSL

[6] Nicolas Sedlmajer, Didier Buchs, Steve Hostettler, Alban Linard, and Edmundo López Bóbeda. A domain specific language approach for genetic regulatory mechanisms analysis. Lecture Notes in Computer Science, 7400: 123–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35179-2_6/.

Westergaard:2012:VPA

[7] Michael Westergaard. Verifying parallel algorithms and programs using coloured Petri nets. Lecture Notes in Computer Science, 7400: 146–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35179-2_7/.

Kordon:2012:RMC

[8] Fabrice Kordon, Alban Linard, Didier Buchs, Maximilien Colange, and Sami Evangelista. Report on the model checking contest at Petri nets 2011. Lecture Notes in Computer Science, 7400:169–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35179-2_8/.

Choppy:2012:MFV

[9] Christine Choppy, Anna Dedova, Sami Evangelista, Kaïs Klaï, and Laure Petrucci. Modelling and formal verification of the NEO protocol. Lecture Notes in Computer Science, 7400:197–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35179-2_9/.

Arnold:2012:ICP

[10] Sonya Arnold and Jonathan Billington. An initial coloured Petri net model of the hypertext transfer protocol operating over the transmission control protocol. Lecture Notes in Computer Science, 7400:226– 250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35179-2_10/.

Suriadi:2012:PCV

[11] Suriadi Suriadi, Chun Ouyang, and Ernest Foo. Privacy compliance verification in cryptographic protocols. Lecture Notes in Computer Science, 7400:251–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35179-2_11/.

Bruneo:2012:MEA

[12] Dario Bruneo, Francesco Longo, and Antonio Puliafito. Modeling energyaware cloud federations with SRNs. Lecture Notes in Computer Science, 7400:277–307, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35179-2_12/.

Marrone:2012:SBM

[13] Stefano Marrone, Nicola Mazzocca, Roberto Nardone, Roberta Presta, and Simon Pietro Romano. A SAN-based modeling approach to performance evaluation of an IMS-compliant conferencing framework. Lecture Notes in Computer Science, 7400:308–333, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35179-2_13/.

vanDongen:2012:ACR

[14] Boudewijn F. van Dongen, Jörg Desel, and Wil M. P. van der Aalst. Aggregating causal runs into Workflow Nets. Lecture Notes in Computer Science, 7400:334–363, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35179-2_14/.

Anonymous:2012:BMa

[15] Anonymous. Back matter. Lecture Notes in Computer Science, 7400: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35179-2/1.

Anonymous:2012:FMa

[16] Anonymous. Front matter. Lecture Notes in Computer Science, 7400: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35179-2/1.

Mavrovouniotis:2012:ISB

[17] Michalis Mavrovouniotis and Shengxiang Yang. An immigrants scheme based on environmental information for ant colony optimization for the dynamic travelling salesman problem. Lecture Notes in Computer Science, 7401:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35533-2_1/.

Diaz-Manriquez:2012:SBI

[18] Alan Díaz-Manríquez, Gregorio Toscano-Pulido, and Ricardo Landa-Becerra. A surrogate-based intelligent variation operator for multiobjective optimization. Lecture Notes in Computer Science, 7401:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35533-2_2/.

daFonseca:2012:RBC

[19] Viviane Grunert da Fonseca and Carlos M. Fonseca. The relationship between the covered fraction, completeness and hypervolume indicators. Lecture Notes in Computer Science, 7401:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35533-2_3/.

Schoenauer:2012:RRA

[20] Marc Schoenauer, Fabien Teytaud, and Olivier Teytaud. A rigorous runtime analysis for quasi-random restarts and decreasing stepsize. Lecture Notes in Computer Science, 7401:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35533-2_4/.

Verel:2012:LON

[21] Sébastien Vérel, Fabio Daolio, Gabriela Ochoa, and Marco Tomassini. Local optima networks with escape edges. Lecture Notes in Computer Science, 7401:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35533-2_5/.

Lutton:2012:VAP

[22] Evelyne Lutton, Julie Foucquier, Nathalie Perrot, Jean Louchet, and Jean-Daniel Fekete. Visual analysis of population scatterplots. Lecture Notes in Computer Science, 7401:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35533-2_6/.

Huijsman:2012:LBD

[23] Robert-Jan Huijsman, Evert Haasdijk, and A. E. Eiben. An on-line onboard distributed algorithm for evolutionary robotics. Lecture Notes in Computer Science, 7401:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35533-2_7/.

Aydin:2012:IPP

[24] Doğan Aydın, Tianjun Liao, Marco A. Montes de Oca, and Thomas Stützle. Improving performance via population growth and local search: The case of the artificial bee colony algorithm. Lecture Notes in Computer Science, 7401:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35533-2_8/.

Maitre:2012:TPF

[25] Ogier Maitre, Nicolas Lachiche, and Pierre Collet. Two ports of a full evolutionary algorithm onto GPGPU. Lecture Notes in Computer Science, 7401:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35533-2_9/.

Robilliard:2012:MTS

[26] Denis Robilliard, Cyril Fonlupt, Virginie Marion-Poty, and Amine Boumaza. A multilevel tabu search with backtracking for exploring weak Schur numbers. Lecture Notes in Computer Science, 7401:109– 119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35533-2_10/.

Rodriguez-Tello:2012:IMA

[27] Eduardo Rodriguez-Tello and Luis Carlos Betancourt. An improved memetic algorithm for the antibandwidth problem. Lecture Notes in Computer Science, 7401:121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35533-2_11/.

vanderGoes:2012:APP

[28] Vincent van der Goes. Adaptive play in a pollution bargaining game. Lecture Notes in Computer Science, 7401:133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35533-2_12/.

Brendel:2012:LOP

[29] Mátyás Brendel and Marc Schoenauer. Learn-and-Optimize: a parameter tuning framework for evolutionary AI planning. Lecture Notes in Computer Science, 7401:145–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35533-2_13/.

DeFalco:2012:MBB

[30] Ivanoe De Falco, Antonio Della Cioppa, Domenico Maisto, and Umberto Scafuri. A model based on biological invasions for Island evolutionary algorithms. Lecture Notes in Computer Science, 7401:157–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35533-2_14/.

Hernandez-Dominguez:2012:MOP

[31] Jorge Sebastian Hernández-Domínguez, Gregorio Toscano-Pulido, and Carlos A. Coello Coello. A multi-objective particle swarm optimizer enhanced with a differential evolution scheme. Lecture Notes in Computer Science, 7401:169–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35533-2_15/.

Inden:2012:EMI

[32] Benjamin Inden, Yaochu Jin, Robert Haschke, and Helge Ritter. Evolution of multisensory integration in large neural fields. Lecture Notes in Computer Science, 7401:181–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35533-2_16/.

Boumaza:2012:RLT

[33] Amine Boumaza. Reducing the learning time of Tetris in evolution strategies. Lecture Notes in Computer Science, 7401:193–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35533-2_17/.

Doerr:2012:BBC

[34] Benjamin Doerr and Carola Winzen. Black-box complexity: Breaking the O(n log n) barrier of LeadingOnes. Lecture Notes in Computer Science, 7401:205–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35533-2_18/.

Roche:2012:ICA

[35] Robin Roche, Lhassane Idoumghar, Benjamin Blunier, and Abdellatif Miraoui. Imperialist competitive algorithm for dynamic optimization of economic dispatch in power systems. Lecture Notes in Computer Science, 7401:217–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35533-2_19/.

Anonymous:2012:BMb

[36] Anonymous. Back matter. Lecture Notes in Computer Science, 7401: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35533-2/1.

Anonymous:2012:FMb

[37] Anonymous. Front matter. Lecture Notes in Computer Science, 7401: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35533-2/1.

He:2012:LBV

[38] Jing (Selena) He, Shouling Ji, and Yi Pan. Load-balanced virtual backbone construction for wireless sensor networks. Lecture Notes in Computer Science, 7402:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31770-5_1/.

Kosowski:2012:MMM

[39] Adrian Kosowski and Alfredo Navarra. Maximum matching in multiinterface networks. Lecture Notes in Computer Science, 7402:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31770-5_2/.

Kranakis:2012:SFW

[40] Evangelos Kranakis and Fraser MacQuarrie. Stretch factor in wireless sensor networks with directional antennae. Lecture Notes in Computer Science, 7402:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31770-5_3/.

Ding:2012:MDC

[41] Wei Ding and Guoliang Xue. On the minimum diameter cost-constrained Steiner tree problem. Lecture Notes in Computer Science, 7402:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31770-5_4/.

Cheng:2012:ECS

[42] Eddie Cheng, Ke Qiu, and Zhizhang Shen. The edge-centered surface area of the arrangement graph. Lecture Notes in Computer Science, 7402:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31770-5_5/.

Yi:2012:ZFN

[43] Eunjeong Yi. On zero forcing number of permutation graphs. Lecture Notes in Computer Science, 7402:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31770-5_6/.

Couetoux:2012:CRE

[44] Basile Couetoux and Jérome Monnot. Complexity results for the Empire Problem in collection of stars. Lecture Notes in Computer Science, 7402:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31770-5_7/.

Biswas:2012:HPC

[45] Sudip Biswas and Stephane Durocher. Hamiltonian paths and cycles in planar graphs. Lecture Notes in Computer Science, 7402:83– 94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31770-5_8/.

Wang:2012:FVS

[46] Chaoyi Wang, Tian Liu, Wei Jiang, and Ke Xu. Feedback vertex sets on tree convex bipartite graphs. Lecture Notes in Computer Science, 7402: 95–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31770-5_9/.

Arikushi:2012:CAG

[47] Karin Arikushi and Csaba D. Tóth. Crossing angles of geometric graphs. Lecture Notes in Computer Science, 7402:103–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31770-5_10/.

Lackner:2012:MGB

[48] Martin Lackner and Reinhard Pichler. Multicut on graphs of bounded clique-width. Lecture Notes in Computer Science, 7402:115–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31770-5_11/.

Zhang:2012:RHM

[49] Chihao Zhang, Haitao Jiang, and Binhai Zhu. Radiation hybrid map construction problem parameterized. Lecture Notes in Computer Science, 7402:127–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31770-5_12/.

Zhu:2012:CPP

[50] Yongding Zhu and Jinhui Xu. On the central path problem. Lecture Notes in Computer Science, 7402:138–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31770-5_13/.

Liu:2012:GMC

[51] Hong Liu and Peng Zhang. On the generalized multiway cut in trees problem. Lecture Notes in Computer Science, 7402:151–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31770-5_14/.

Liang:2012:AFL

[52] Zhewei Liang and Kaizhong Zhang. Algorithms for forest local similarity. Lecture Notes in Computer Science, 7402:163–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31770-5_15/.

Frid:2012:SRP

[53] Yelena Frid and Dan Gusfield. Speedup of RNA pseudoknotted secondary structure recurrence computation with the Four-Russians Method. Lecture Notes in Computer Science, 7402:176–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31770-5_16/.

Chen:2012:IAA

[54] Zhi-Zhong Chen and Lusheng Wang. An improved approximation algorithm for the bandpass-2 problem. Lecture Notes in Computer Science, 7402:188–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31770-5_17/.

ElOuali:2012:MPH

[55] Mourad El Ouali and Gerold Jäger. The b-matching problem in hypergraphs: Hardness and approximability. Lecture Notes in Computer Science, 7402:200–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31770-5_18/.

Zhang:2012:RSS

[56] Qiang Zhang, Weiwei Wu, and Minming Li. Resource scheduling with supply constraint and linear cost. Lecture Notes in Computer Science, 7402:212–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31770-5_19/.

Kanj:2012:CGP

[57] Iyad A. Kanj and Ge Xia. On certain geometric properties of the Yao–Yao graphs. Lecture Notes in Computer Science, 7402:223–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31770-5_20/.

Anonymous:2012:FMc

[58] Anonymous. Front matter. Lecture Notes in Computer Science, 7402: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31770-5/1.

Vempala:2012:EPC

[59] Santosh S. Vempala. Effective principal component analysis. Lecture Notes in Computer Science, 7404:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_1/.

Zezula:2012:FTS

[60] Pavel Zezula. Future trends in similarity searching. Lecture Notes in Computer Science, 7404:8–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_2/.

Barrios:2012:STD

[61] Juan Manuel Barrios, Benjamin Bustos, and Tomáš Skopal. Snake table: a dynamic pivot table for streams of k-NN searches. Lecture Notes in Computer Science, 7404:25–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_3/.

Skopal:2012:AEA

[62] Tomáš Skopal and Tomáš Bartoš. Algorithmic exploration of axiom spaces for efficient similarity search at large scale. Lecture Notes in Computer Science, 7404:40–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32153-5_4/.

Tellez:2012:PMI

[63] Eric Sadit Tellez, Edgar Chavez, and Karina Figueroa. Polyphasic metric index: Reaching the practical limits of proximity searching. Lecture Notes in Computer Science, 7404:54–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_5/.

Ares:2012:ESS

[64] Luis G. Ares, Nieves R. Brisaboa, Alberto Ordóñez Pereira, and Oscar Pedreira. Efficient similarity search in metric spaces with cluster reduction. Lecture Notes in Computer Science, 7404:70–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32153-5_6/.

Lokoc:2012:CRC

[65] Jakub Lokoč, Přemysl Čech, Jiří Novák, and Tomáš Skopal. Cut-Region: a compact building block for hierarchical metric indexing. Lecture Notes in Computer Science, 7404:85–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_7/.

Naidan:2012:SDT

[66] Bilegsaikhan Naidan and Magnus Lie Hetland. Static-to-dynamic transformation for metric indexing structures. Lecture Notes in Computer Science, 7404:101–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32153-5_8/.

Britos:2012:DTD

[67] Luis Britos, A. Marcela Printista, and Nora Reyes. DSACL+-tree: a dynamic data structure for similarity search in secondary memory. Lecture Notes in Computer Science, 7404:116–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32153-5_9/.

Malkov:2012:SDA

[68] Yury Malkov, Alexander Ponomarenko, Andrey Logvinov, and Vladimir Krylov. Scalable distributed algorithm for approximate nearest neighbor search problem in high dimensional general metric spaces. Lecture Notes in Computer Science, 7404:132–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_10/.

Mohamed:2012:PAP

[69] Hisham Mohamed and Stéphane Marchand-Maillet. Parallel approaches to permutation-based indexing using inverted files. Lecture Notes in Computer Science, 7404:148–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_11/. See erratum [79].

Boytsov:2012:SLI

[70] Leonid Boytsov. Super-linear indices for approximate dictionary searching. Lecture Notes in Computer Science, 7404:162–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32153-5_12/.

Lokoc:2012:VIS

[71] Jakub Lokoč, David Novák, Michal Batko, and Tomáš Skopal. Visual image search: Feature signatures or/and global descriptors. Lecture Notes in Computer Science, 7404:177–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32153-5_13/.

Bartos:2012:RTL

[72] Tomáš Bartoš and Tomáš Skopal. Revisiting techniques for lowerbounding the dynamic time warping distance. Lecture Notes in Computer Science, 7404:192–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32153-5_14/.

Connor:2012:MCD

[73] Richard Connor and Robert Moss. A multivariate correlation distance for vector spaces. Lecture Notes in Computer Science, 7404:209–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32153-5_15/.

Houle:2012:FSC

[74] Michael E. Houle, Hisashi Kashima, and Michael Nett. Fast similarity computation in factorized tensors. Lecture Notes in Computer Science, 7404:226–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32153-5_16/.

Lokoc:2012:SSI

[75] Jakub Lokoč, Tomáš Grošup, and Tomáš Skopal. SIR: The smart image retrieval engine. Lecture Notes in Computer Science, 7404:240–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32153-5_17.

Novak:2012:SSS

[76] Jiří Novák, Jakub Galgonek, David Hoksza, and Tomáš Skopal. Sim-Tandem: Similarity search in tandem mass spectra. Lecture Notes in Computer Science, 7404:242–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32153-5_18.

Anonymous:2012:BMc

[77] Anonymous. Back matter. Lecture Notes in Computer Science, 7404: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32153-5/1.

Anonymous:2012:FMd

[78] Anonymous. Front matter. Lecture Notes in Computer Science, 7404: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32153-5/1.

Mohamed:2012:EPA

[79] Hisham Mohamed and Stéphane Marchand-Maillet. Erratum: Parallel approaches to permutation-based indexing using inverted files. Lecture Notes in Computer Science, 7404:E1–E2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-32153-5_19. See [69].

Li:2012:TDN

[80] Jian Li, Luoyi Fu, and Xinbing Wang. Throughput and delay with network coding in hybrid mobile ad hoc networks: a global perspective. Lecture Notes in Computer Science, 7405:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31869-6_1/.

Wang:2012:HHB

[81] Shengling Wang, Min Liu, and Xiuzhen Cheng. HERO — a home based routing in pocket switched networks. Lecture Notes in Computer Science, 7405:20–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31869-6_2/.

Cheng:2012:RIL

[82] Wei Cheng, Dengyuan Wu, and Xiuzhen Cheng. Routing for information leakage reduction in multi-channel multi-hop ad-hoc social networks. Lecture Notes in Computer Science, 7405:31–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31869-6_3/.

Wu:2012:ANI

[83] Chunyang Wu, Fan Wu, and Guihai Chen. AMPLE: a novel incentive approach to adaptive-width channel allocation in multi-hop, noncooperative wireless networks. Lecture Notes in Computer Science, 7405: 43–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31869-6_4/.

Xu:2012:CHN

[84] Kuai Xu, Feng Wang, Lin Gu, and Jianhua Gao. Characterizing home network traffic: An inside view. Lecture Notes in Computer Science, 7405:60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31869-6_5/.

Wei:2012:SSP

[85] Zuying Wei, Tianrong Zhang, and Fan Wu. SHIELD: a strategy-proof and highly efficient channel auction mechanism for multi-radio wireless networks. Lecture Notes in Computer Science, 7405:72–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31869-6_6/.

Nguyen:2012:NBA

[86] Nam Tuan Nguyen, Yichuan Wang, and Xin Liu. A nonparametric Bayesian approach for opportunistic data transfer in cellular networks. Lecture Notes in Computer Science, 7405:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31869-6_7/.

Zhou:2012:OPV

[87] Yangfan Zhou, Xinyu Chen, and Michael R. Lyu. Online protocol verification in wireless sensor networks via non-intrusive behavior profiling. Lecture Notes in Computer Science, 7405:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31869-6_8/.

Xia:2012:SBP

[88] Na Xia, Ruji Feng, and Lina Xu. SPSA based packet size optimization algorithm in wireless sensor networks. Lecture Notes in Computer Science, 7405:112–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31869-6_9/.

Sun:2012:LBD

[89] Xianwei Sun, Scott C.-H. Huang, and Minming Li. Lower bounds on data collection time in sensor networks. Lecture Notes in Computer Science, 7405:120–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31869-6_10/.

Li:2012:MTC

[90] Deying Li, Donghyun Kim, and Qinghua Zhu. Minimum total communication power connected dominating set in wireless networks. Lecture Notes in Computer Science, 7405:132–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31869-6_11/.

Lv:2012:SML

[91] Jingsong Lv, Victor C. S. Lee, and Minming Li. Supporting multi-level quality of services in data broadcast systems. Lecture Notes in Computer Science, 7405:142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31869-6_12/.

El-Moukaddem:2012:MNT

[92] Fatme El-Moukaddem and Eric Torng. Maximizing network topology lifetime using mobile node rotation. Lecture Notes in Computer Science, 7405:154–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31869-6_13/.

Zhao:2012:JBP

[93] Feng Zhao, Bin Li, and Hongbin Chen. Joint beamforming and power allocation algorithm for cognitive MIMO systems via game theory. Lecture Notes in Computer Science, 7405:166–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31869-6_14/.

Xu:2012:BAS

[94] Xiaorong Xu, Aiping Huang, and Jianwu Zhang. Bit allocation scheme with primary base station cooperation in cognitive radio network. Lecture Notes in Computer Science, 7405:178–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31869-6_15/.

Wadhwa:2012:OSS

[95] Manish Wadhwa, Chunsheng Xin, and Min Song. Optimal spectrum sharing for contention-based cognitive radio wireless networks. Lecture Notes in Computer Science, 7405:187–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31869-6_16/.

Xu:2012:FGC

[96] Xiaohua Xu, Jiannong Cao, and Peng-Jun Wan. Fast group communication scheduling in duty-cycled multihop wireless sensor networks. Lecture Notes in Computer Science, 7405:197–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31869-6_17/.

Abdelhakim:2012:RCS

[97] Mai Abdelhakim, Jian Ren, and Tongtong Li. Reliable cooperative sensing in cognitive networks. Lecture Notes in Computer Science, 7405: 206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31869-6_18/.

Cao:2012:HMN

[98] Xuanyu Cao, Jinbei Zhang, and Guanglin Zhang. Heterogeneous multicast networks with wireless helping networks. Lecture Notes in Computer Science, 7405:218–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31869-6_19/.

Champion:2012:EPU

[99] Adam C. Champion, Xinfeng Li, and Qiang Zhai. Enclave: Promoting unobtrusive and secure mobile communications with a ubiquitous electronic world. Lecture Notes in Computer Science, 7405:235–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31869-6_20/.

Anonymous:2012:FMe

[100] Anonymous. Front matter. Lecture Notes in Computer Science, 7405: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31869-6/1.

Paulson:2012:MPF

[101] Lawrence C. Paulson. MetiTarski: Past and future. Lecture Notes in Computer Science, 7406:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32347-8_1/.

Barthe:2012:CAC

[102] Gilles Barthe, Juan Manuel Crespo, Benjamin Grégoire, and César Kunz. Computer-aided cryptographic proofs. Lecture Notes in Computer Science, 7406:11–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32347-8_2/.

Platzer:2012:DOA

[103] André Platzer. A differential operator approach to equational differential invariants. Lecture Notes in Computer Science, 7406:28–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32347-8_3/.

Gacek:2012:AT

[104] Andrew Gacek. Abella: a tutorial. Lecture Notes in Computer Science, 7406:49–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-32347-8_4.

Gamboa:2012:CTD

[105] Ruben Gamboa and John Cowles. A Cantor trio: Denumerability, the reals, and the real algebraic numbers. Lecture Notes in Computer Science, 7406:51–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32347-8_5/.

Cohen:2012:CRA

[106] Cyril Cohen. Construction of real algebraic numbers in Coq. Lecture Notes in Computer Science, 7406:67–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32347-8_6/.

Denes:2012:RBA

[107] Maxime Dénès, Anders Mörtberg, and Vincent Siles. A refinement-based approach to computational algebra in coq. Lecture Notes in Computer Science, 7406:83–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32347-8_7/.

Greenaway:2012:BGA

[108] David Greenaway, June Andronick, and Gerwin Klein. Bridging the gap: Automatic verified abstraction of C. Lecture Notes in Computer Science, 7406:99–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32347-8_8/.

Nipkow:2012:AIA

[109] Tobias Nipkow. Abstract interpretation of annotated commands. Lecture Notes in Computer Science, 7406:116–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32347-8_9/.

Michel:2012:VGW

[110] Patrick Michel and Arnd Poetzsch-Heffter. Verifying and generating WP transformers for procedures on complex data. Lecture Notes in Computer Science, 7406:133–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32347-8_10/.

Danielsson:2012:BEP

[111] Nils Anders Danielsson. Bag equivalence via a proof-relevant membership relation. Lecture Notes in Computer Science, 7406:149–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32347-8_11/.

Lammich:2012:ADR

[112] Peter Lammich and Thomas Tuerk. Applying data refinement for monadic programs to Hopcroft’s algorithm. Lecture Notes in Computer Science, 7406:166–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32347-8_12/.

Miculan:2012:SDM

[113] Marino Miculan and Marco Paviotti. Synthesis of distributed mobile programs using monadic types in coq. Lecture Notes in Computer Science, 7406:183–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32347-8_13/.

Baelde:2012:TPR

[114] David Baelde, Pierre Courtieu, David Gross-Amblard, and Christine Paulin-Mohring. Towards provably robust watermarking. Lecture Notes in Computer Science, 7406:201–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32347-8_14/.

Zhang:2012:PIP

[115] Xingyuan Zhang, Christian Urban, and Chunhan Wu. Priority inheritance protocol proved correct. Lecture Notes in Computer Science, 7406: 217–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32347-8_15/.

Affeldt:2012:FST

[116] Reynald Affeldt and Manabu Hagiwara. Formalization of Shannon’s theorems in SSReflect–Coq. Lecture Notes in Computer Science, 7406: 233–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32347-8_16/.

Vytiniotis:2012:SWY

[117] Dimitrios Vytiniotis, Thierry Coquand, and David Wahlstedt. Stop when you are almost-full. Lecture Notes in Computer Science, 7406: 250–265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32347-8_17/.

Sternagel:2012:CNP

[118] Christian Sternagel and René Thiemann. Certification of nontermination proofs. Lecture Notes in Computer Science, 7406:266–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32347-8_18/.

Asperti:2012:CPD

[119] Andrea Asperti. A compact proof of decidability for regular expression equivalence. Lecture Notes in Computer Science, 7406:283–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32347-8_19/.

Mansky:2012:ULD

[120] William Mansky and Elsa L. Gunter. Using locales to define a relyguarantee temporal logic. Lecture Notes in Computer Science, 7406: 299–314, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32347-8_20/.

Anonymous:2012:FMf

[121] Anonymous. Front matter. Lecture Notes in Computer Science, 7406: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32347-8/1.

Austrin:2012:NPN

[122] Per Austrin, Ryan O’Donnell, and John Wright. A new point of NPhardness for 2-to-1 label cover. Lecture Notes in Computer Science, 7408:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32512-0_1/.

Austrin:2012:ITO

[123] Per Austrin, Toniann Pitassi, and Yu Wu. Inapproximability of treewidth, one-shot pebbling, and related layout problems. Lecture Notes in Computer Science, 7408:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32512-0_2/.

Awasthi:2012:AAN

[124] Pranjal Awasthi, Avrim Blum, Jamie Morgenstern, and Or Sheffet. Additive approximation for near-perfect phylogeny construction. Lecture Notes in Computer Science, 7408:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32512-0_3/.

Awasthi:2012:ISN

[125] Pranjal Awasthi and Or Sheffet. Improved spectral-norm bounds for clustering. Lecture Notes in Computer Science, 7408:37–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32512-0_4/.

Berman:2012:PDA

[126] Piotr Berman and Grigory Yaroslavtsev. Primal-dual approximation algorithms for node-weighted network design in planar graphs. Lecture Notes in Computer Science, 7408:50–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32512-0_5/.

Boufounos:2012:WFK

[127] Petros Boufounos, Volkan Cevher, Anna C. Gilbert, Yi Li, and Martin J. Strauss. What’s the frequency, Kenneth?: Sublinear Fourier sampling off the grid. Lecture Notes in Computer Science, 7408:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32512-0_6/.

Chalermsook:2012:IHR

[128] Parinya Chalermsook, Julia Chuzhoy, Sampath Kannan, and Sanjeev Khanna. Improved hardness results for profit maximization pricing problems with unlimited supply. Lecture Notes in Computer Science, 7408: 73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32512-0_7/.

Chan:2012:OFT

[129] Ho-Leung Chan, Tak-Wah Lam, and Rongbin Li. Online flow time scheduling in the presence of preemption overhead. Lecture Notes in Computer Science, 7408:85–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32512-0_8/.

Chekuri:2012:PCS

[130] Chandra Chekuri, Alina Ene, and Ali Vakilian. Prize-collecting survivable network design in node-weighted graphs. Lecture Notes in Computer Science, 7408:98–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32512-0_9/.

Cheriyan:2012:AMC

[131] Joseph Cheriyan, Zachary Friggstad, and Zhihan Gao. Approximating minimum-cost connected T-joins. Lecture Notes in Computer Science, 7408:110–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32512-0_10/.

Dinitz:2012:ICC

[132] Michael Dinitz and Gordon Wilfong. iBGP and constrained connectivity. Lecture Notes in Computer Science, 7408:122–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32512-0_11/.

Epstein:2012:OSJ

[133] Leah Epstein, Łukasz Jeż, Jiří Sgall, and Rob van Stee. Online scheduling of jobs with fixed start times on related machines. Lecture Notes in Computer Science, 7408:134–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32512-0_12/.

Fernandes:2012:SAB

[134] Cristina G. Fernandes, Luís A. A. Meira, Flávio K. Miyazawa, and Lehilton L. C. Pedrosa. A systematic approach to bound factor revealing LPs and its application to the metric and squared metric facility location problems. Lecture Notes in Computer Science, 7408: 146–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32512-0_13/.

Guruswami:2012:ABO

[135] Venkatesan Guruswami and Yuan Zhou. Approximating bounded occurrence ordering CSPs. Lecture Notes in Computer Science, 7408: 158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32512-0_14/.

Haastad:2012:NHM

[136] Johan Håstad. On the NP-hardness of max-not-2. Lecture Notes in Computer Science, 7408:170–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32512-0_15/.

Haviv:2012:RSP

[137] Ishay Haviv. The remote set problem on lattices. Lecture Notes in Computer Science, 7408:182–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32512-0_16/.

Hellwig:2012:AAG

[138] Matthias Hellwig and Alexander Souza. Approximation algorithms for generalized and variable-sized bin covering. Lecture Notes in Computer Science, 7408:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32512-0_17/.

Iwata:2012:AML

[139] Satoru Iwata, Prasad Tetali, and Pushkar Tripathi. Approximating minimum linear ordering problems. Lecture Notes in Computer Science, 7408: 206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32512-0_18/.

Khuller:2012:NAR

[140] Samir Khuller, Barna Saha, and Kanthi K. Sarpatwar. New approximation results for resource replication problems. Lecture Notes in Computer Science, 7408:218–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32512-0_19/.

Konrad:2012:MMS

[141] Christian Konrad, Frédéric Magniez, and Claire Mathieu. Maximum matching in semi-streaming with few passes. Lecture Notes in Computer Science, 7408:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32512-0_20/.

Anonymous:2012:FMg

[142] Anonymous. Front matter. Lecture Notes in Computer Science, 7408: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32512-0/1.

Cunha:2012:TDL

[143] Luís Felipe I. Cunha and Luis Antonio B. Kowada. Transposition diameter and lonely permutations. Lecture Notes in Computer Science, 7409:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31927-3_1/.

Feijao:2012:EAF

[144] Pedro Feijao and Joao Meidanis. Extending the algebraic formalism for genome rearrangements to include linear chromosomes. Lecture Notes in Computer Science, 7409:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31927-3_2/.

Galvao:2012:ARA

[145] Gustavo Rodrigues Galvão and Zanoni Dias. On the approximation ratio of algorithms for sorting by transpositions without using cycle graphs. Lecture Notes in Computer Science, 7409:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31927-3_3/.

Matroud:2012:CTH

[146] A. A. Matroud, C. P. Tuffley, and D. Bryant. A comparison of three heuristic methods for solving the parsing problem for tandem repeats. Lecture Notes in Computer Science, 7409:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31927-3_4/.

Lorenz:2012:RFA

[147] Ronny Lorenz and Stephan H. Bernhart. RNA folding algorithms with G-quadruplexes. Lecture Notes in Computer Science, 7409:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31927-3_5/.

Benitez:2012:MDS

[148] César Manuel Vargas Benítez. Molecular dynamics for simulating the protein folding process using the 3D AB off-lattice model. Lecture Notes in Computer Science, 7409:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31927-3_6/.

Vilela:2012:BBW

[149] Halian Vilela and Tainá Raiol. A bioconductor based workflow for Z-DNA region detection and biological inference. Lecture Notes in Computer Science, 7409:73–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31927-3_7/.

Braz:2012:PMC

[150] Fernando A. F. Braz and Jader S. Cruz. A probabilistic model checking approach to investigate the palytoxin effects on the Na^{+} /K^{+}-ATPase. Lecture Notes in Computer Science, 7409:84–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31927-3_8/.

Mendoza:2012:RRF

[151] Mariana R. Mendoza and Guilherme C. da Fonseca. RFMirTarget: a random forest classifier for human miRNA target gene prediction. Lecture Notes in Computer Science, 7409:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31927-3_9/.

Gusmao:2012:PTF

[152] Eduardo G. Gusmão and Christoph Dieterich. Prediction of transcription factor binding sites by integrating DNase digestion and histone modification. Lecture Notes in Computer Science, 7409:109–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31927-3_10/.

Jaskowiak:2012:ECC

[153] Pablo A. Jaskowiak. Evaluating correlation coefficients for clustering gene expression profiles of cancer. Lecture Notes in Computer Science, 7409: 120–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31927-3_11/.

Rosa:2012:AGS

[154] Rogério S. Rosa and Rafael H. S. Santos. Associating genotype sequence properties to haplotype inference errors. Lecture Notes in Computer Science, 7409:132–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31927-3_12/.

Lifschitz:2012:DIP

[155] Sérgio Lifschitz and Carlos Juliano M. Viana. Design and implementation of ProteinWorldDB. Lecture Notes in Computer Science, 7409: 144–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31927-3_13/.

Winck:2012:CAP

[156] Ana T. Winck and Christian V. Quevedo. A comparative analysis of public ligand databases based on molecular descriptors. Lecture Notes in Computer Science, 7409:156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31927-3_14/.

Martins:2012:PDC

[157] Wellington S. Martins and Thiago F. Rangel. Phylogenetic distance computation using CUDA. Lecture Notes in Computer Science, 7409: 168–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31927-3_15/.

Ocana:2012:EME

[158] Kary A. C. S. Ocaña and Daniel de Oliveira. Exploring molecular evolution reconstruction using a parallel cloud based scientific workflow. Lecture Notes in Computer Science, 7409:179–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31927-3_16/.

Garcia:2012:HBE

[159] José Fernando Garcia. How bioinformatics enables livestock applied sciences in the genomic era. Lecture Notes in Computer Science, 7409: 192–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31927-3_17/.

Anonymous:2012:BMd

[160] Anonymous. Back matter. Lecture Notes in Computer Science, 7409: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31927-3/1.

Anonymous:2012:FMh

[161] Anonymous. Front matter. Lecture Notes in Computer Science, 7409: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31927-3/1.

Szilard:2012:KHD

[162] Andrew L. Szilard. The kind hearted dragon Prof. Sheng Yu, 1950– 2012. Lecture Notes in Computer Science, 7410:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31653-1_1/.

Csuhaj-Varju:2012:PDA

[163] Erzsébet Csuhaj-Varjú. P and dP automata: Unconventional versus classical automata. Lecture Notes in Computer Science, 7410:7–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31653-1_2/.

Iwama:2012:RSO

[164] Kazuo Iwama. Recovering strings in oracles: Quantum and classic. Lecture Notes in Computer Science, 7410:23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31653-1_3.

Hromkovic:2012:DVN

[165] Juraj Hromkovič, Rastislav Královič, Richard Královič, and Richard Štefanec. Determinism vs. nondeterminism for two-way automata. Lecture Notes in Computer Science, 7410:24–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31653-1_4/.

Kari:2012:CAC

[166] Jarkko Kari. Cellular automata, the Collatz Conjecture and powers of 3/2. Lecture Notes in Computer Science, 7410:40–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31653-1_5/.

Brzozowski:2012:QCA

[167] Janusz Brzozowski and Hellis Tamm. Quotient complexities of atoms of regular languages. Lecture Notes in Computer Science, 7410:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31653-1_6/.

Beal:2012:DGR

[168] Marie-Pierre Béal, Jean-Marc Champarnaud, Jean-Philippe Dubernard, and Hadrien Jeanne. Decidability of geometricity of regular languages.

Lecture Notes in Computer Science, 7410:62–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31653-1_7/.

Schmid:2012:ICR

[169] Markus L. Schmid. Inside the class of REGEX languages. Lecture Notes in Computer Science, 7410:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31653-1_8/.

Han:2012:CED

[170] Yo-Sub Han, Sang-Ki Ko, and Kai Salomaa. Computing the edit-distance between a regular language and a context-free language. Lecture Notes in Computer Science, 7410:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31653-1_9/.

Hoffmann:2012:SCF

[171] Michael Hoffmann, Derek F. Holt, Matthew D. Owens, and Richard M. Thomas. Semigroups with a context-free word problem. Lecture Notes in Computer Science, 7410:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31653-1_10/.

Holzer:2012:GDS

[172] Markus Holzer, Sebastian Jakobi, and Ian McQuillan. Generalized derivations with synchronized context-free grammars. Lecture Notes in Computer Science, 7410:109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31653-1_11/.

Okhotin:2012:NEV

[173] Alexander Okhotin. Non-erasing variants of the Chomsky–Schützenberger theorem. Lecture Notes in Computer Science, 7410:121–129, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-31653-1_12/

.

Reidenbach:2012:RCF

[174] Daniel Reidenbach and Markus L. Schmid. Regular and context-free pattern languages over small alphabets. Lecture Notes in Computer Science, 7410:130–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31653-1_13/.

Esik:2012:CFL

[175] Zoltan Ésik and Satoshi Okawa. On context-free languages of scattered words. Lecture Notes in Computer Science, 7410:142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31653-1_14/.

Lehtinen:2012:HPD

[176] Tommi Lehtinen and Alexander Okhotin. Homomorphisms preserving deterministic context-free languages. Lecture Notes in Computer Science, 7410:154–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31653-1_15/.

Geffert:2012:UCN

[177] Viliam Geffert and Dana Pardubská. Unary coded NP-complete languages in ASPACE(log log n). Lecture Notes in Computer Science, 7410: 166–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31653-1_16/.

Krebs:2012:DC

[178] Andreas Krebs and Klaus-Jörn Lange. Dense completeness. Lecture Notes in Computer Science, 7410:178–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31653-1_17/.

Holzer:2012:EAE

[179] Markus Holzer and Sebastian Jakobi. From equivalence to almostequivalence, and beyond— minimizing automata with errors. Lecture Notes in Computer Science, 7410:190–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31653-1_18/.

Kapoutsis:2012:AFT

[180] Christos A. Kapoutsis and Nans Lefebvre. Analogs of Fagin’s theorem for small nondeterministic finite automata. Lecture Notes in Computer Science, 7410:202–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31653-1_19/.

Kutrib:2012:SHD

[181] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. States and heads do count for unary multi-head finite automata. Lecture Notes in Computer Science, 7410:214–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31653-1_20/.

Anonymous:2012:FMi

[182] Anonymous. Front matter. Lecture Notes in Computer Science, 7410: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31653-1/1.

Jones:2012:SSL

[183] Norah Jones and Esyin Chew. Social software for learning — the institutional policy of the University of Glamorgan. Lecture Notes in Computer Science, 7411:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32018-7_1/.

Huang:2012:FDL

[184] Ronghuai Huang, Haisen Zhang, Yushun Li, and Junfeng Yang. A framework of designing learning activities for mobile learning. Lecture Notes in Computer Science, 7411:9–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32018-7_2/.

Krcadinac:2012:VAS

[185] Uroš Krčadinac, Jelena Jovanović, and Vladan Devedžić. Visualizing the affective structure of students interaction. Lecture Notes in Computer Science, 7411:23–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32018-7_3/.

Lau:2012:IDI

[186] Nga-Sin Lau and Lui Lam. An investigation of the determinants influencing student learning motivation via Facebook private group in teaching and learning. Lecture Notes in Computer Science, 7411:35– 44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32018-7_4/.

Yen:2012:PSD

[187] Cheng-Huang Yen. A preliminary study of disturbing silence in asynchronous discussion forums. Lecture Notes in Computer Science, 7411: 45–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32018-7_5/.

Chen:2012:ARH

[188] Huiling Chen, Guoqing Zhao, and Ningyi Xu. The analysis of research hotspots and fronts of knowledge visualization based on CiteSpace II. Lecture Notes in Computer Science, 7411:57–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32018-7_6/.

Chai:2012:PRA

[189] Shaoming Chai and Meirong Ding. Practical research on the assessment of online collaborative English learning — a case study of blackboardbased course “intercultural communication”. Lecture Notes in Computer Science, 7411:69–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32018-7_7/.

Li:2012:CSS

[190] Shuang Li, Pengfeng Shi, and Qi Tang. A case study on scaffolding design for Wiki-based collaborative knowledge building. Lecture Notes in Computer Science, 7411:77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32018-7_8/.

Cheung:2012:SUM

[191] Simon K. S. Cheung. A study on the use of mobile devices for distance learning. Lecture Notes in Computer Science, 7411:89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32018-7_9/.

Yang:2012:DCA

[192] Harrison Hao Yang. The development of collaborative action research through cloud computing document-sharing services and blended learning process. Lecture Notes in Computer Science, 7411:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32018-7_10/.

Li:2012:RCP

[193] Wenchong Li and Xiuping Tu. Research on the comprehensive practice course of primary school based on knowledge building. Lecture Notes in Computer Science, 7411:109–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32018-7_11/.

Diu:2012:CEM

[194] Chin Kee Diu and Wai Yu Ng. Cost effective methodology in E-adoption for E-training providers. Lecture Notes in Computer Science, 7411: 120–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32018-7_12/.

Hirata:2012:LLH

[195] Yoko Hirata and Yoshihiro Hirata. Learning logs in hybrid language learning: Students’ views and preferences. Lecture Notes in Computer Science, 7411:132–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32018-7_13/.

Bedi:2012:SFF

[196] Kanishka Bedi. The significance of face-to-face instruction in hybrid executive education. Lecture Notes in Computer Science, 7411:144–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32018-7_14/.

Zuo:2012:SNA

[197] Pingping Zuo, Su Mu, and Xiaomeng Han. A social network analysis of students’ online interaction in hybrid learning — a case study of “media and teaching” course. Lecture Notes in Computer Science, 7411: 155–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32018-7_15/.

Lam:2012:RML

[198] J. Lam and C. G. Duan. A review of mobile learning environment in higher education sector of hong kong: Technological and social perspectives. Lecture Notes in Computer Science, 7411:165–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32018-7_16/.

Li:2012:MSC

[199] Yanyan Li and Li Wang. Modeling semantic context for active elearning in the workplace. Lecture Notes in Computer Science, 7411: 174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32018-7_17/.

Lu:2012:IFB

[200] Xing Lu, Guodong Zhao, and Jianfeng Jiang. Influential factors of blended learning in Chinese colleges: From the perspective of instructor’s acceptance and students’ satisfaction. Lecture Notes in Computer Science, 7411: 186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32018-7_18/.

Xu:2012:CCM

[201] Ningyi Xu, Guoqing Zhao, Huiling Chen, and Leisi Pei. The colored concept map and its application in learning assistance program. Lecture Notes in Computer Science, 7411:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32018-7_19/.

Fong:2012:PSM

[202] Joseph Fong, Kenneth Ting Yan Wong, and Tracy Wu. PTA system: Mobile computing student assessment by parent and teacher association. Lecture Notes in Computer Science, 7411:210–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32018-7_20/.

Anonymous:2012:FMj

[203] Anonymous. Front matter. Lecture Notes in Computer Science, 7411: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32018-7/1.

Alimomeni:2012:GS

[204] Mohsen Alimomeni and Reihaneh Safavi-Naini. Guessing secrecy. Lecture Notes in Computer Science, 7412:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32284-6_1/.

Broadbent:2012:TRC

[205] Anne Broadbent, Stacey Jeffery, and Samuel Ranellucci. Trading robustness for correctness and privacy in certain multiparty computations, beyond an honest majority. Lecture Notes in Computer Science, 7412: 14–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32284-6_2/.

Canetti:2012:TPD

[206] Ran Canetti, Ben Riva, and Guy N. Rothblum. Two protocols for delegation of computation. Lecture Notes in Computer Science, 7412: 37–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32284-6_3/.

Cramer:2012:ACZ

[207] Ronald Cramer, Ivan Damgård, and Valerio Pastro. On the amortized complexity of zero knowledge protocols for multiplicative relations. Lecture Notes in Computer Science, 7412:62–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32284-6_4/.

David:2012:UCO

[208] Bernardo Machado David and Anderson C. A. Nascimento. Universally composable oblivious transfer from lossy encryption and the McEliece assumptions. Lecture Notes in Computer Science, 7412:80– 99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32284-6_5/.

Dodis:2012:SIR

[209] Yevgeniy Dodis. Shannon impossibility, revisited. Lecture Notes in Computer Science, 7412:100–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32284-6_6/.

Dottling:2012:SSL

[210] Nico Döttling, Daniel Kraschewski, and Jörn Müller-Quade. Statistically secure linear-rate dimension extension for oblivious affine function evaluation. Lecture Notes in Computer Science, 7412:111–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32284-6_7/.

Hirt:2012:PCS

[211] Martin Hirt, Christoph Lucas, Ueli Maurer, and Dominik Raub. Passive corruption in statistical multi-party computation. Lecture Notes in Computer Science, 7412:129–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32284-6_8/.

Keller:2012:ETZ

[212] Marcel Keller, Gert Læssøe Mikkelsen, and Andy Rupp. Efficient threshold zero-knowledge with applications to user-centric protocols. Lecture Notes in Computer Science, 7412:147–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32284-6_9/.

Watanabe:2012:ITT

[213] Yohei Watanabe, Takenobu Seito, and Junji Shikata. Informationtheoretic timed-release security: Key-agreement, encryption, and authentication codes. Lecture Notes in Computer Science, 7412:167–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32284-6_10/.

Yoshida:2012:OGT

[214] Maki Yoshida, Toru Fujiwara, and Marc Fossorier. Optimum general threshold secret sharing. Lecture Notes in Computer Science, 7412: 187–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32284-6_11/.

Anonymous:2012:BMe

[215] Anonymous. Back matter. Lecture Notes in Computer Science, 7412: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32284-6/1.

Anonymous:2012:FMk

[216] Anonymous. Front matter. Lecture Notes in Computer Science, 7412: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32284-6/1.

Yao:2012:OTT

[217] Yiyu Yao. An outline of a theory of three-way decisions. Lecture Notes in Computer Science, 7413:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32115-3_1/.

Li:2012:MGM

[218] Shaoyong Li, Tianrui Li, and Chuan Luo. A multi-granulation model under dominance-based rough set approach. Lecture Notes in Computer Science, 7413:18–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32115-3_2/.

Wang:2012:RRM

[219] Hailiang Wang and Mingtian Zhou. A refined rough k-means clustering with hybrid threshold. Lecture Notes in Computer Science, 7413: 26–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32115-3_3/.

Zeng:2012:IAU

[220] Anping Zeng, Tianrui Li, Junbo Zhang, and Dun Liu. An incremental approach for updating approximations of rough fuzzy sets under the variation of the object set. Lecture Notes in Computer Science, 7413: 36–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32115-3_4/.

Clark:2012:HGP

[221] Patrick G. Clark, Jerzy W. Grzymala-Busse, and Zdzislaw S. Hippe. How good are probabilistic approximations for rule induction from data with missing attribute values? Lecture Notes in Computer Science, 7413: 46–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32115-3_5/.

Blaszczynski:2012:IOC

[222] Jerzy Błaszczyński, Roman Słowiński, and Marcin Szelag. Induction of ordinal classification rules from incomplete data. Lecture Notes in Computer Science, 7413:56–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32115-3_6/.

Yang:2012:IVF

[223] Hai-Long Yang. Interval valued fuzzy rough set model on two different universes and its application. Lecture Notes in Computer Science, 7413:66–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32115-3_7/.

Chan:2012:MCS

[224] Chien-Chung Chan. Maximal characteristic sets and neighborhoods approach to incomplete information systems. Lecture Notes in Computer Science, 7413:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32115-3_8/.

Sun:2012:PFR

[225] Bingzhen Sun, Weimin Ma, Haiyan Zhao, and Xinxin Wang. Probabilistic fuzzy rough set model over two universes. Lecture Notes in Computer Science, 7413:83–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32115-3_9/.

Slezak:2012:ROC

[226] Dominik Ślezak, Piotr Synak, Jakub Wróblewski, Janusz Borkowski, and Graham Toppin. Rough optimizations of complex expressions in Infobright’s RDBMS. Lecture Notes in Computer Science, 7413:94–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32115-3_10/.

Liu:2012:RSB

[227] Yang Liu, Guohua Bai, Qinglei Zhou, and Elisabeth Rakus-Andersson. Rough sets based inequality rule learner for knowledge discovery. Lecture Notes in Computer Science, 7413:100–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32115-3_11/.

Soliman:2012:CAD

[228] Omar S. Soliman and Aboul Ella Hassanien. A computer aided diagnosis system for breast cancer using support vector machine. Lecture Notes in Computer Science, 7413:106–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32115-3_12/.

Yang:2012:NMA

[229] Yong Yang and Yunxia Lu. A new method of the accurate eye corner location. Lecture Notes in Computer Science, 7413:116–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32115-3_13/.

Tamanini:2012:HAV

[230] Isabelle Tamanini, Plácido Rogério Pinheiro, and Cícero Nogueira dos Santos. A hybrid approach of verbal decision analysis and machine learning. Lecture Notes in Computer Science, 7413:126–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32115-3_14/.

Chen:2012:EDI

[231] Long Chen, Banglan Liu, and Wei Song. An efficient data integrity verification method supporting multi-granular operation with commutative hash. Lecture Notes in Computer Science, 7413:132–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32115-3_15/.

Xieski:2012:DBM

[232] Tomasz Xieski, Agnieszka Nowak-Brzezińska, and Alicja Wakulicz-Deja. Density-based method for clustering and visualization of complex data. Lecture Notes in Computer Science, 7413:142–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-32115-3_16/.

Nowak-Brzezinska:2012:IPU

[233] Agnieszka Nowak-Brzezińska, Tomasz Jach, and Alicja Wakulicz-Deja. Inference processes using incomplete knowledge in decision support systems — chosen aspects. Lecture Notes in Computer Science, 7413: 150–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32115-3_17/.

Janusz:2012:IDI

[234] Andrzej Janusz, Wojciech Świeboda, Adam Krasuski, and Hung Son Nguyen. Interactive document indexing method based on explicit semantic analysis. Lecture Notes in Computer Science, 7413:156–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32115-3_18/.

Huang:2012:MTD

[235] Bo Huang, Yan Yang, Amjad Mahmood, and Hongjun Wang. Microblog topic detection based on LDA model and single-pass clustering. Lecture Notes in Computer Science, 7413:166–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32115-3_19/.

Szczuka:2012:NEM

[236] Marcin Szczuka, Paweł Betliński, and Kamil Herba. Named entity matching in publication databases. Lecture Notes in Computer Science, 7413: 172–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32115-3_20/.

Anonymous:2012:FMl

[237] Anonymous. Front matter. Lecture Notes in Computer Science, 7413: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32115-3/1.

Wasilewski:2012:CRS

[238] Piotr Wasilewski and Andrzej Skowron. A characterization of rough separability. Lecture Notes in Computer Science, 7414:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31900-6_1/.

Wang:2012:DDV

[239] Guoyin Wang and Lihe Guan. Data-driven valued tolerance relation. Lecture Notes in Computer Science, 7414:11–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31900-6_2/.

Xu:2012:OMG

[240] Weihua Xu, Wenxin Sun, and Yufeng Liu. Optimistic multi-granulation fuzzy rough set model based on triangular norm. Lecture Notes in Computer Science, 7414:20–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31900-6_3/.

Zhou:2012:RSM

[241] Junyi Zhou and Xibei Yang. Rough set model based on hybrid tolerance relation. Lecture Notes in Computer Science, 7414:28–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31900-6_4/.

Prasad:2012:SIQ

[242] P. S. V. S. Sai Prasad and C. Raghavendra Rao. Scalable improved quick reduct: Sample based. Lecture Notes in Computer Science, 7414: 34–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31900-6_5/.

Qin:2012:SRS

[243] Keyun Qin, Zhenming Song, and Yang Xu. Soft rough sets based on similarity measures. Lecture Notes in Computer Science, 7414:40–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31900-6_6/.

Liu:2012:TTD

[244] Huan Liu, Yingcang Ma, and Xuezhen Dai. Theory of truth degrees in three valued formed system RSL. Lecture Notes in Computer Science, 7414:49–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31900-6_7/.

Xu:2012:UAR

[245] Weihua Xu, Yufeng Liu, and Wenxin Sun. Upper approximation reduction based on intuitionistic fuzzy T equivalence information systems. Lecture Notes in Computer Science, 7414:55–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31900-6_8/.

Liu:2012:FRS

[246] Yang Liu, Qinglei Zhou, Elisabeth Rakus-Andersson, and Guohua Bai. A fuzzy-rough sets based compact rule induction method for classifying hybrid data. Lecture Notes in Computer Science, 7414:63–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31900-6_9/.

Chai:2012:NIF

[247] Junyi Chai, James N. K. Liu, and Anming Li. A new intuitionistic fuzzy rough set approach for decision support. Lecture Notes in Computer Science, 7414:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-31900-6_10/.

Matsubayashi:2012:NRI

[248] Tsukasa Matsubayashi, Yuichi Kato, and Tetsuro Saeki. A new rule induction method from a decision table using a statistical test. Lecture Notes in Computer Science, 7414:81–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31900-6_11/.

Kavita:2012:RNA

[249] Singh Kavita, Zaveri Mukesh, and Raghuwanshi Mukesh. A rough neurocomputing approach for illumination invariant face recognition system. Lecture Notes in Computer Science, 7414:91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31900-6_12/.

Zhao:2012:ADE

[250] Hongbo Zhao, Feng Jiang, and Chunping Wang. An approximation decision entropy based decision tree algorithm and its application in intrusion detection. Lecture Notes in Computer Science, 7414:101–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31900-6_13/.

Palkowski:2012:ARS

[251] Łukasz Pałkowski, Jerzy Błaszczyński, Jerzy Krysiński, and Roman Słowiński. Application of rough set theory to prediction of antimicrobial activity of bis-quaternary ammonium chlorides. Lecture Notes in Computer Science, 7414:107–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31900-6_14/.

Deng:2012:CDB

[252] Dayong Deng, Lin Chen, and Dianxun Yan. Classification and decision based on parallel reducts and F -rough sets. Lecture Notes in Computer Science, 7414:117–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31900-6_15/.

Joshi:2012:ECR

[253] Manish Joshi and Pawan Lingras. Evidential clustering or rough clustering: The choice is yours. Lecture Notes in Computer Science, 7414: 123–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31900-6_16/.

Trabelsi:2012:HAS

[254] Salsabil Trabelsi, Zied Elouedi, and Pawan Lingras. Heuristic for attribute selection using belief discernibility matrix. Lecture Notes in Computer Science, 7414:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31900-6_17/.

Tsumoto:2012:IRI

[255] Shusaku Tsumoto and Shoji Hirano. Incremental rules induction based on rule layers. Lecture Notes in Computer Science, 7414:139–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31900-6_18/.

Alsolami:2012:OID

[256] Fawaz Alsolami, Igor Chikalov, Mikhail Moshkov, and Beata Zielosko. Optimization of inhibitory decision rules relative to length and coverage. Lecture Notes in Computer Science, 7414:149–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31900-6_19/.

Yang:2012:PCA

[257] Yong Yang and Zhengrong Chen. Parallelized computing of attribute core based on rough set theory and MapReduce. Lecture Notes in Computer Science, 7414:155–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31900-6_20/.

Anonymous:2012:FMm

[258] Anonymous. Front matter. Lecture Notes in Computer Science, 7414: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31900-6/1.

Fox:2012:PPA

[259] Gwendoline Fox and Juliette Thuilier. Predicting the position of attributive adjectives in the French NP. Lecture Notes in Computer Science, 7415:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31467-4_1/.

Janicek:2012:ARC

[260] Miroslav Janíček. Abductive reasoning for continual dialogue understanding. Lecture Notes in Computer Science, 7415:16–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-31467-4_2/.

List:2012:SPA

[261] Johann-Mattis List. SCA: Phonetic alignment based on sound classes. Lecture Notes in Computer Science, 7415:32–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31467-4_3/.

Pauw:2012:EQ

[262] Simon Pauw and Michael Spranger. Embodied quantifiers. Lecture Notes in Computer Science, 7415:52–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31467-4_4/.

Wlodarczak:2012:RMD

[263] Marcin Włodarczak. Ranked multidimensional dialogue act annotation. Lecture Notes in Computer Science, 7415:67–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31467-4_5/.

Fichte:2012:GBO

[264] Johannes Klaus Fichte. The good, the bad, and the odd: Cycles in answer-set programs. Lecture Notes in Computer Science, 7415:78–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31467-4_6/.

Ghari:2012:DKJ

[265] Meghdad Ghari. Distributed knowledge with justifications. Lecture Notes in Computer Science, 7415:91–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31467-4_7/.

Holliday:2012:ELR

[266] Wesley H. Holliday. Epistemic logic, relevant alternatives, and the dynamics of context. Lecture Notes in Computer Science, 7415:109–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31467-4_8/.

Perkov:2012:TGM

[267] Tin Perkov. Towards a generalization of modal definability. Lecture Notes in Computer Science, 7415:130–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31467-4_9/.

Rendsvig:2012:MSC

[268] Rasmus K. Rendsvig. Modeling semantic competence: a critical review of Frege’s puzzle about identity. Lecture Notes in Computer Science, 7415: 140–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31467-4_10/.

Weinzierl:2012:CIR

[269] Antonius Weinzierl. Comparing inconsistency resolutions in multicontext systems. Lecture Notes in Computer Science, 7415:158–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31467-4_11/.

Burnett:2012:VDP

[270] Heather Burnett. Vague determiner phrases and distributive predication. Lecture Notes in Computer Science, 7415:175–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31467-4_12/.

deVries:2012:SSE

[271] Hanna de Vries. The syntax and semantics of evaluative degree modification. Lecture Notes in Computer Science, 7415:195–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31467-4_13/.

Hansen:2012:KSP

[272] Casper Storm Hansen. A Kripkean solution to paradoxes of denotation. Lecture Notes in Computer Science, 7415:212–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31467-4_14/.

Quinley:2012:TGM

[273] Jason Quinley. Trust games as a model for requests. Lecture Notes in Computer Science, 7415:221–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31467-4_15/.

Anonymous:2012:BMf

[274] Anonymous. Back matter. Lecture Notes in Computer Science, 7415: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31467-4/1.

Anonymous:2012:FMn

[275] Anonymous. Front matter. Lecture Notes in Computer Science, 7415: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31467-4/1.

Bai:2012:WUA

[276] Aijun Bai, Xiaoping Chen, Patrick MacAlpine, Daniel Urieli, Samuel Barrett, and Peter Stone. WrightEagle and UT Austin Villa: RoboCup 2011 Simulation League Champions. Lecture Notes in Computer Science, 7416:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32060-6_1/.

Sukvichai:2012:RHS

[277] Kanjanapan Sukvichai, Teeratath Ariyachartphadungkit, and Krit Chaiso. Robot hardware, software, and technologies behind the SKUBA robot team. Lecture Notes in Computer Science, 7416:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32060-6_2/.

Laue:2012:BHE

[278] Tim Laue, Thomas Röfer, Katharina Gillmann, Felix Wenk, Colin Graf, and Tobias Kastner. B-Human 2011 — eliminating game delays. Lecture Notes in Computer Science, 7416:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32060-6_3/.

Lee:2012:RHL

[279] Daniel D. Lee, Seung-Joon Yi, Stephen McGill, Yida Zhang, Sven Behnke, and Marcell Missura. Robocup 2011 Humanoid League winners. Lecture Notes in Computer Science, 7416:37–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32060-6_4/.

Stuckler:2012:TRM

[280] Jörg Stückler, David Droeschel, Kathrin Gräve, Dirk Holz, and Jochen Kläß. Towards robust mobility, flexible object manipulation, and intuitive multimodal interaction for domestic service robots. Lecture Notes in Computer Science, 7416:51–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32060-6_5/.

Eguchi:2012:RDL

[281] Amy Eguchi, Nicky Hughes, Matthias Stocker, Jiayao Shen, and Naomi Chikuma. RoboCupJunior — a decade later. Lecture Notes in Computer Science, 7416:63–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32060-6_6/.

Stuckler:2012:CTS

[282] Jörg Stückler and Sven Behnke. Compliant task-space control with backdrivable servo actuators. Lecture Notes in Computer Science, 7416: 78–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32060-6_7/.

Norouzi:2012:PSP

[283] Mohammad Norouzi, Freek De Bruijn, and Jaime Valls Miró. Planning stable paths for urban search and rescue robots. Lecture Notes in Computer Science, 7416:90–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32060-6_8/.

Graf:2012:CMO

[284] Colin Graf and Thomas Röfer. A center of mass observing 3D-LIPM gait for the RoboCup standard platform league humanoid. Lecture Notes in Computer Science, 7416:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32060-6_9/.

Cunha:2012:BIB

[285] João Cunha, Nuno Lau, and João Rodrigues. Ball interception behaviour in robotic soccer. Lecture Notes in Computer Science, 7416: 114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32060-6_10/.

Urbann:2012:RSB

[286] Oliver Urbann, Sören Kerner, and Stefan Tasse. Rigid and soft body simulation featuring realistic walk behaviour. Lecture Notes in Computer Science, 7416:126–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32060-6_11/.

Mueller:2012:TRO

[287] Christian A. Mueller, Nico Hochgeschwender, and Paul G. Ploeger. Towards robust object categorization for mobile robots with combination of classifiers. Lecture Notes in Computer Science, 7416:137–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32060-6_12/.

Metzler:2012:LVO

[288] Saskia Metzler, Matthias Nieuwenhuisen, and Sven Behnke. Learning visual obstacle detection using color histogram features. Lecture Notes in Computer Science, 7416:149–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32060-6_13/.

Moballegh:2012:GVG

[289] Hamid Moballegh, Naja von Schmude, and Raúl Rojas. Gradient vector griding: An approach to shape-based object detection in RoboCup scenarios. Lecture Notes in Computer Science, 7416:162–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32060-6_14/.

Sadeh-Or:2012:AFL

[290] Eran Sadeh-Or and Gal A. Kaminka. AnySURF: Flexible local features computation. Lecture Notes in Computer Science, 7416:174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32060-6_15/.

Valtazanos:2012:OMP

[291] Aris Valtazanos and Subramanian Ramamoorthy. Online motion planning for multi-robot interaction using composable reachable sets. Lecture Notes in Computer Science, 7416:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32060-6_16/.

Schmitz:2012:RTT

[292] Andreas Schmitz, Marcell Missura, and Sven Behnke. Real-time trajectory generation by offline footstep planning for a humanoid soccer robot. Lecture Notes in Computer Science, 7416:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32060-6_17/.

Xue:2012:RTB

[293] Feng Xue, Xiaoping Chen, Jinsu Liu, and Daniele Nardi. Real time biped walking gait pattern generator for a real robot. Lecture Notes in Computer Science, 7416:210–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32060-6_18/.

Jochmann:2012:EMH

[294] Gregor Jochmann, Sören Kerner, Stefan Tasse, and Oliver Urbann. Efficient multi-hypotheses unscented Kalman filtering for robust localization. Lecture Notes in Computer Science, 7416:222–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32060-6_19/.

Marchant:2012:PGT

[295] Román Marchant, Pablo Guerrero, and Javier Ruiz del Solar. A portable ground-truth system based on a laser sensor. Lecture Notes in Computer Science, 7416:234–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32060-6_20/.

Anonymous:2012:FMo

[296] Anonymous. Front matter. Lecture Notes in Computer Science, 7416: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32060-6/1.

Hoang:2012:ESB

[297] Viet Tung Hoang, Ben Morris, and Phillip Rogaway. An enciphering scheme based on a card shuffle. Lecture Notes in Computer Science, 7417:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32009-5_1/.

Landecker:2012:TBB

[298] Will Landecker, Thomas Shrimpton, and R. Seth Terashima. Tweakable blockciphers with beyond birthday-bound security. Lecture Notes in Computer Science, 7417:14–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32009-5_2/.

Iwata:2012:BRG

[299] Tetsu Iwata, Keisuke Ohashi, and Kazuhiko Minematsu. Breaking and repairing GCM security proofs. Lecture Notes in Computer Science, 7417: 31–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32009-5_3/.

Abdelraheem:2012:DLB

[300] Mohamed Ahmed Abdelraheem and Martin Ågren. On the distribution of linear biases: Three instructive examples. Lecture Notes in Computer Science, 7417:50–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32009-5_4/.

Miles:2012:SPN

[301] Eric Miles and Emanuele Viola. Substitution-permutation networks, pseudorandom functions, and natural proofs. Lecture Notes in Computer Science, 7417:68–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32009-5_5/.

Zittrain:2012:EC

[302] Jonathan Zittrain. The end of crypto. Lecture Notes in Computer Science, 7417:86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-32009-5_6.

Rosulek:2012:MYK

[303] Mike Rosulek. Must you know the code of f to securely compute f? Lecture Notes in Computer Science, 7417:87–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32009-5_7/.

Garg:2012:ASM

[304] Sanjam Garg and Amit Sahai. Adaptively secure multi-party computation with dishonest majority. Lecture Notes in Computer Science, 7417:

105–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32009-5_8/.

Alwen:2012:CPC

[305] Joël Alwen, Jonathan Katz, and Ueli Maurer. Collusion-preserving computation. Lecture Notes in Computer Science, 7417:124–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32009-5_9/.

Beimel:2012:SSS

[306] Amos Beimel, Oriol Farràs, and Yuval Mintz. Secret sharing schemes for very dense graphs. Lecture Notes in Computer Science, 7417:144– 161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32009-5_10/.

Gorbunov:2012:FEB

[307] Sergey Gorbunov, Vinod Vaikuntanathan, and Hoeteck Wee. Functional encryption with bounded collusions via multi-party computation. Lecture Notes in Computer Science, 7417:162–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32009-5_11/.

Lewko:2012:NPM

[308] Allison Lewko and Brent Waters. New proof methods for attribute-based encryption: Achieving full security through selective techniques. Lecture Notes in Computer Science, 7417:180–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32009-5_12/.

Sahai:2012:DCC

[309] Amit Sahai, Hakan Seyalioglu, and Brent Waters. Dynamic credentials and ciphertext delegation for attribute-based encryption. Lecture Notes in Computer Science, 7417:199–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32009-5_13/.

Waters:2012:FER

[310] Brent Waters. Functional encryption for regular languages. Lecture Notes in Computer Science, 7417:218–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32009-5_14/.

Chase:2012:SDC

[311] Melissa Chase and Ivan Visconti. Secure database commitments and universal arguments of quasi knowledge. Lecture Notes in Computer Science, 7417:236–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32009-5_15/.

Bitansky:2012:SAM

[312] Nir Bitansky and Alessandro Chiesa. Succinct arguments from multiprover interactive proofs and their efficiency benefits. Lecture Notes in Computer Science, 7417:255–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32009-5_16/.

Jager:2012:STD

[313] Tibor Jager, Florian Kohlar, and Sven Schäge. On the security of TLS-DHE in the standard model. Lecture Notes in Computer Science, 7417: 273–293, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32009-5_17/.

Bellare:2012:SSW

[314] Mihir Bellare, Stefano Tessaro, and Alexander Vardy. Semantic security for the wiretap channel. Lecture Notes in Computer Science, 7417: 294–311, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32009-5_18/.

Bellare:2012:MIS

[315] Mihir Bellare, Thomas Ristenpart, and Stefano Tessaro. Multi-instance security and its application to password-based cryptography. Lecture Notes in Computer Science, 7417:312–329, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32009-5_19/.

Mennink:2012:HFB

[316] Bart Mennink and Bart Preneel. Hash functions based on three permutations: a generic security analysis. Lecture Notes in Computer Science, 7417:330–347, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32009-5_20/.

Anonymous:2012:FMp

[317] Anonymous. Front matter. Lecture Notes in Computer Science, 7417: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32009-5/1.

Fan:2012:DQT

[318] Wenfei Fan. Data quality: Theory and practice. Lecture Notes in Computer Science, 7418:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32281-5_1/.

Han:2012:CWB

[319] Jiawei Han. Construction of Web-based, service-oriented information networks: a data mining perspective. Lecture Notes in Computer Science, 7418:17–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_2/.

Agrawal:2012:ETD

[320] Rakesh Agrawal, Sreenivas Gollapudi, Anitha Kannan, and Krishnaram Kenthapadi. Electronic textbooks and data mining. Lecture Notes in Computer Science, 7418:20–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32281-5_3.

Mo:2012:TBD

[321] Shangfeng Mo, Hong Chen, and Yinglong Li. Topology-based data compression in wireless sensor networks. Lecture Notes in Computer Science, 7418:22–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_4/.

Tan:2012:REB

[322] Long Tan. A residual energy-based fairness scheduling MAC protocol for wireless sensor networks. Lecture Notes in Computer Science, 7418: 35–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_5/.

Li:2012:TAG

[323] Guilin Li, Longjiang Guo, Jian Zhang, and Minghong Liao. Topologyaided geographic routing protocol for wireless sensor networks. Lecture Notes in Computer Science, 7418:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32281-5_6/.

Zhang:2012:PFB

[324] Nan Zhang and Jianhua Feng. Polaris: a fingerprint-based localization system over wireless networks. Lecture Notes in Computer Science, 7418: 58–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_7/.

Qu:2012:HPA

[325] Jun-Feng Qu and Mengchi Liu. A high-performance algorithm for frequent itemset mining. Lecture Notes in Computer Science, 7418:71– 82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32281-5_8/.

Zhang:2012:MLP

[326] Xiang Zhang, Cuifang Zhao, Peng Wang, and Fengbo Zhou. Mining link patterns in linked data. Lecture Notes in Computer Science, 7418: 83–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_9/.

Song:2012:DPO

[327] Kaisong Song, Daling Wang, Shi Feng, Dong Wang, and Ge Yu. Detecting positive opinion leader group from forum. Lecture Notes in Computer Science, 7418:95–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32281-5_10/.

Negi:2012:DTB

[328] Sumit Negi, Manish A. Bhide, Vishal S. Batra, Mukesh K. Mohania, and Sunil Bajpai. D’MART: a tool for building and populating data warehouse model from existing reports and tables. Lecture Notes in Computer Science, 7418:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32281-5_11/.

Lin:2012:CSQ

[329] Xin Lin, Jianliang Xu, and Junzhong Gu. Continuous skyline queries with integrity assurance in outsourced spatial databases. Lecture Notes in Computer Science, 7418:114–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32281-5_12/.

Suzuki:2012:AQV

[330] Yu Suzuki. Assessing quality values of Wikipedia articles using implicit positive and negative ratings. Lecture Notes in Computer Science, 7418: 127–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_13/.

Wu:2012:FBI

[331] Hao Wu and Lizhu Zhou. Form-based instant search and query autocompletion on relational data. Lecture Notes in Computer Science, 7418: 139–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_14/.

Zhang:2012:RQE

[332] Yan Zhang, Long Yang, and Hongzhi Wang. Range query estimation for dirty data management system. Lecture Notes in Computer Science, 7418: 152–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32281-5_15/.

Sun:2012:TMI

[333] Yu Sun, Jin Huang, Yueguo Chen, Xiaoyong Du, and Rui Zhang. Topk most incremental location selection with capacity constraint. Lecture Notes in Computer Science, 7418:165–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32281-5_16/.

Li:2012:ATB

[334] Renfei Li, Daling Wang, Yifei Zhang, Shi Feng, and Ge Yu. An approach of text-based and image-based multi-modal search for online shopping. Lecture Notes in Computer Science, 7418:172–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32281-5_17/.

Hemayati:2012:CSR

[335] Reza Taghizadeh Hemayati, Weiyi Meng, and Clement Yu. Categorizing search results using WordNet and Wikipedia. Lecture Notes in Computer Science, 7418:185–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32281-5_18/.

Htoo:2012:OSR

[336] Htoo Htoo, Yutaka Ohsawa, Noboru Sonehara, and Masao Sakauchi. Optimal sequenced route query algorithm using visited POI graph. Lecture Notes in Computer Science, 7418:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32281-5_19/.

Cui:2012:PAM

[337] Jifeng Cui, Yong Zhang, Chao Li, and ChunXiao Xing. A packaging approach for massive amounts of small geospatial files with HDFS. Lecture Notes in Computer Science, 7418:210–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32281-5_20/.

Anonymous:2012:FMq

[338] Anonymous. Front matter. Lecture Notes in Computer Science, 7418: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32281-5/1.

Miao:2012:APX

[339] Haixu Miao, Tiezheng Nie, Dejun Yue, Tiancheng Zhang, and Jinshen Liu. Algebra for parallel XQuery processing. Lecture Notes in Computer Science, 7419:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33050-6_1/.

Wang:2012:GSD

[340] Cong Wang and Hongzhi Wang. Graph-structured data compression based on frequent subgraph contraction. Lecture Notes in Computer Science, 7419:11–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_2/.

Lin:2012:PPR

[341] Xin Lin, Lingchen Zhou, Peng Chen, and Junzhong Gu. Privacy preserving reverse nearest-neighbor queries processing on road network. Lecture Notes in Computer Science, 7419:19–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33050-6_3/.

Wu:2012:WRG

[342] Jianmin Wu and Mizuho Iwaihara. Wikipedia revision graph extraction based on N-gram cover. Lecture Notes in Computer Science, 7419: 29–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_4/.

Ren:2012:WNI

[343] Meirui Ren, Jinsheng Duan, Hao Qu, Xinjing Wang, and Lei Du. Wireless nerve: Invisible anti-theft system in wireless sensor network. Lecture Notes in Computer Science, 7419:39–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33050-6_5/.

He:2012:STM

[344] Zaobo He and Xiaoming Wang. A spatial-temporal model for the malware propagation in MWSNs based on the reaction-diffusion equations. Lecture Notes in Computer Science, 7419:45–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33050-6_6/.

Tan:2012:ANL

[345] Long Tan and Liyan Chen. The application of a node-localization algorithm of wireless sensor network in intelligent transportation system. Lecture Notes in Computer Science, 7419:57–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33050-6_7/.

Luo:2012:PNV

[346] Juan Luo, Lei Chen, Shan Fu, and Renfa Li. Parallel network virtualization resource management system. Lecture Notes in Computer Science, 7419:69–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_8/.

Zao:2012:APB

[347] Yang Zao, Liu Yan, and Li Renfa. The analysis of priority-based slotted CSMA/CA algorithm in IEEE 802.15.4 sensor network. Lecture Notes in Computer Science, 7419:78–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33050-6_9/.

Feng:2012:QBD

[348] Cheng Feng, Rui Zhang, Shouxu Jiang, and Zhijun Li. QoI-based data gathering and routing guidance in VANETs. Lecture Notes in Computer Science, 7419:87–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33050-6_10/.

Yang:2012:PPL

[349] Robin Jun Yang and Qiong Luo. PTL: Partitioned logging for database storage on flash solid state drives. Lecture Notes in Computer Science, 7419:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_11/.

Allayear:2012:AMI

[350] Shaikh Muhammad Allayear, Sung Soon Park, Shamim H. Ripon, and Gyeong Hun Kim. Adaptation mechanism of iSCSI protocol for NAS storage solution in wireless environment. Lecture Notes in Computer Science, 7419:109–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33050-6_12/.

Koonsanit:2012:BSH

[351] Kitti Koonsanit, Chuleerat Jaruskulchai, and Apisit Eiumnoh. Band selection for hyperspectral imagery with PCA-MIG. Lecture Notes in Computer Science, 7419:119–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33050-6_13/.

Zhang:2012:NTO

[352] Jing Zhang, Xiaoguang Hong, Zhaohui Peng, and Qingzhong Li. NestedCube: Towards online analytical processing on information-enhanced multidimensional network. Lecture Notes in Computer Science, 7419: 128–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_14/.

Liu:2012:MEA

[353] Yi Liu, Luo Chen, Ning Jing, and Wei Xiong. MRFM: An efficient approach to spatial join aggregate. Lecture Notes in Computer Science, 7419: 140–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_15/.

Li:2012:DII

[354] Xu Li, Xin Wang, Hong Shi, Zhaohua Sheng, and Zhiyong Feng. A distributed inverted indexing scheme for large-scale RDF data. Lecture Notes in Computer Science, 7419:151–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33050-6_16/.

Pan:2012:MAS

[355] Wei Pan, Zhanhuai Li, Qun Chen, Shanglian Peng, Bo Suo, and Jian Xu. MSMapper: An adaptive split assignment scheme for MapReduce. Lecture Notes in Computer Science, 7419:162–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33050-6_17/.

Zhang:2012:DER

[356] Chunyu Zhang, Yong Su, Jiyang Chen, and Wen Wang. Driving environment reconstruction and analysis system on multi-sensor network. Lecture Notes in Computer Science, 7419:173–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33050-6_18/.

Yu:2012:LGH

[357] Zhou Yu, Jian Shao, and Fei Wu. LuSH: a generic high-dimensional index framework. Lecture Notes in Computer Science, 7419:181–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33050-6_19/.

Zhu:2012:ITS

[358] Huaijie Zhu, Xiaochun Yang, Bin Wang, and Yue Wang. Improving text search on hybrid data. Lecture Notes in Computer Science, 7419: 192–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33050-6_20/.

Anonymous:2012:FMr

[359] Anonymous. Front matter. Lecture Notes in Computer Science, 7419: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33050-6/1.

Banerjee:2012:MVO

[360] Anindya Banerjee. Modular verification of object-based programs. Lecture Notes in Computer Science, 7421:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-31762-0_1.

Bormer:2012:CIV

[361] Thorsten Bormer, Marc Brockschmidt, Dino Distefano, and Gidon Ernst. The COST IC0701 verification competition 2011. Lecture Notes in Computer Science, 7421:3–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31762-0_2/.

Hunt:2012:PAF

[362] James J. Hunt. The practical application of formal methods: Where is the benefit for industry? Lecture Notes in Computer Science, 7421: 22–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31762-0_3/.

Mycroft:2012:ITM

[363] Alan Mycroft. Isolation types and multi-core architectures. Lecture Notes in Computer Science, 7421:33–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31762-0_4/.

Wong:2012:MAD

[364] Peter Y. H. Wong, Nikolay Diakov, and Ina Schaefer. Modelling adaptable distributed object oriented systems using the HATS approach: a Fredhopper case study. Lecture Notes in Computer Science, 7421:49– 66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31762-0_5/.

Balakrishnan:2012:MAI

[365] Gogul Balakrishnan, Naoto Maeda, Sriram Sankaranarayanan, Franjo Ivančić, and Aarti Gupta. Modeling and analyzing the interaction of C and C++ strings. Lecture Notes in Computer Science, 7421:67–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31762-0_6/.

Beckert:2012:IBM

[366] Bernhard Beckert, Thorsten Bormer, Florian Merz, and Carsten Sinz. Integration of bounded model checking and deductive verification. Lecture Notes in Computer Science, 7421:86–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31762-0_7/.

Bentea:2012:PFO

[367] Lucian Bentea and Olaf Owe. A probabilistic framework for objectoriented modeling and analysis of distributed systems. Lecture Notes in Computer Science, 7421:105–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31762-0_8/.

Brockschmidt:2012:ADN

[368] Marc Brockschmidt, Thomas Ströder, Carsten Otto, and Jürgen Giesl. Automated detection of non-termination and NullPointerExceptions for Java bytecode. Lecture Notes in Computer Science, 7421:123–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31762-0_9/.

Heinen:2012:JAJ

[369] Jonathan Heinen, Henrik Barthels, and Christina Jansen. Juggrnaut — an abstract JVM. Lecture Notes in Computer Science, 7421:142– 159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31762-0_10/.

Herranz:2012:VIP

[370] Ángel Herranz and Julio Mariño. A verified implementation of priority monitors in Java. Lecture Notes in Computer Science, 7421:160– 177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31762-0_11/.

Huisman:2012:SSC

[371] Marieke Huisman and Tri Minh Ngo. Scheduler-specific confidentiality for multi-threaded programs and its logic-based verification. Lecture Notes in Computer Science, 7421:178–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31762-0_12/.

Johnsen:2012:FMU

[372] Einar Broch Johnsen, Rudolf Schlatte, and S. Lizeth Tapia Tarifa. A formal model of user-defined resources in resource-restricted deployment scenarios. Lecture Notes in Computer Science, 7421:196–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-31762-0_13/.

Rusu:2012:BFF

[373] Vlad Rusu and Dorel Lucanu. A K-based formal framework for domainspecific modelling languages. Lecture Notes in Computer Science, 7421: 214–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-31762-0_14/.

Scheben:2012:VIF

[374] Christoph Scheben and Peter H. Schmitt. Verification of information flow properties of Java programs without approximations. Lecture Notes in Computer Science, 7421:232–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31762-0_15/.

Anonymous:2012:BMg

[375] Anonymous. Back matter. Lecture Notes in Computer Science, 7421: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31762-0/1.

Anonymous:2012:FMs

[376] Anonymous. Front matter. Lecture Notes in Computer Science, 7421: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31762-0/1.

Ausiello:2012:STO

[377] Giorgio Ausiello, Giuseppe F. Italiano, Luigi Laura, Umberto Nanni, and Fabiano Sarracco. Structure theorems for optimum hyperpaths in directed hypergraphs. Lecture Notes in Computer Science, 7422: 1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32147-4_1/.

Hewitt:2012:BPG

[378] Mike Hewitt, George L. Nemhauser, and Martin Savelsbergh. Branchand-price guided search. Lecture Notes in Computer Science, 7422: 15–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32147-4_2/.

Papadimitriou:2012:NFC

[379] Christos H. Papadimitriou. The new faces of combinatorial optimization. Lecture Notes in Computer Science, 7422:19–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32147-4_3/.

Cacchiani:2012:MAT

[380] Valentina Cacchiani, Alberto Caprara, and Paolo Toth. Models and algorithms for the train unit assignment problem. Lecture Notes in Computer Science, 7422:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32147-4_4/.

Piva:2012:MST

[381] Breno Piva and Cid C. de Souza. The minimum stabbing triangulation problem: IP models and computational evaluation. Lecture Notes in Computer Science, 7422:36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32147-4_5/.

Fischetti:2012:OS

[382] Matteo Fischetti and Leo Liberti. Orbital shrinking. Lecture Notes in Computer Science, 7422:48–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32147-4_6/.

Ostrowski:2012:USO

[383] James Ostrowski. Using symmetry to optimize over the Sherali–Adams relaxation. Lecture Notes in Computer Science, 7422:59–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32147-4_7/.

Cheng:2012:SOC

[384] Jianqiang Cheng, Céline Gicquel, and Abdel Lisser. A second-order cone programming approximation to joint chance-constrained linear programs. Lecture Notes in Computer Science, 7422:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32147-4_8/.

Gorge:2012:SRM

[385] Agnès Gorge, Abdel Lisser, and Riadh Zorgati. Semidefinite relaxations for mixed 0–1 second-order cone program. Lecture Notes in Computer Science, 7422:81–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32147-4_9/.

Fouilhoux:2012:NDR

[386] Pierre Fouilhoux and Aurélien Questel. The non-disjoint m-ring-star problem: Polyhedral results and SDH/SONET network design. Lecture Notes in Computer Science, 7422:93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32147-4_10/.

Borne:2012:UAT

[387] Sylvie Borne, Roland Grappe, and Mathieu Lacroix. The uncapacitated asymmetric traveling salesman problem with multiple stacks. Lecture Notes in Computer Science, 7422:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32147-4_11/.

Lacroix:2012:PAB

[388] Mathieu Lacroix, A. Ridha Mahjoub, and Sébastien Martin. Polyhedral analysis and branch-and-cut for the structural analysis problem. Lecture Notes in Computer Science, 7422:117–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32147-4_12/.

Faenza:2012:EFN

[389] Yuri Faenza, Samuel Fiorini, Roland Grappe, and Hans Raj Tiwary. Extended formulations, nonnegative factorizations, and randomized communication protocols. Lecture Notes in Computer Science, 7422:129– 140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32147-4_13/.

Braun:2012:AAS

[390] Gábor Braun and Sebastian Pokutta. An algebraic approach to symmetric extended formulations. Lecture Notes in Computer Science, 7422: 141–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32147-4_14/.

Fujishige:2012:DCS

[391] Satoru Fujishige and Jens Maßberg. Dual consistent systems of linear inequalities and cardinality constrained polytopes. Lecture Notes in Computer Science, 7422:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32147-4_15/.

Freire:2012:MRC

[392] A. S. Freire, V. Acuña, P. Crescenzi, C. E. Ferreira, V. Lacroix, and P. V. Milreu. Minimum ratio cover of matrix columns by extreme rays of its induced cone. Lecture Notes in Computer Science, 7422: 165–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32147-4_16/.

Galli:2012:GIM

[393] Laura Galli, Konstantinos Kaparis, and Adam N. Letchford. Gap inequalities for the Max-Cut problem: a cutting-plane algorithm. Lecture Notes in Computer Science, 7422:178–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32147-4_17/.

Dokka:2012:FSA

[394] Trivikram Dokka, Ioannis Mourtos, and Frits C. R. Spieksma. Fast separation algorithms for three-index assignment problems. Lecture Notes in Computer Science, 7422:189–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32147-4_18/.

Gouveia:2012:HCS

[395] Luis Gouveia, Markus Leitner, and Ivana Ljubić. On the hop constrained Steiner tree problem with multiple root nodes. Lecture Notes in Computer Science, 7422:201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32147-4_19/.

Voge:2012:TIA

[396] Marie-Emilie Voge and François Clautiaux. Theoretical investigation of aggregation in pseudo-polynomial network-flow models. Lecture Notes in Computer Science, 7422:213–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32147-4_20/.

Anonymous:2012:FMt

[397] Anonymous. Front matter. Lecture Notes in Computer Science, 7422: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32147-4/1.

Koolen:2012:OIB

[398] Marijn Koolen, Gabriella Kazai, Jaap Kamps, Antoine Doucet, and Monica Landoni. Overview of the INEX 2011 books and social search track. Lecture Notes in Computer Science, 7424:1–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35734-3_1/.

Adriaans:2012:IDR

[399] Frans Adriaans, Jaap Kamps, and Marijn Koolen. The importance of document ranking and user-generated content for faceted search and book suggestions. Lecture Notes in Computer Science, 7424:30–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35734-3_2/.

Bogers:2012:RIS

[400] Toine Bogers, Kirstine Wilfred Christensen, and Birger Larsen. RSLIS at INEX 2011: Social book search track. Lecture Notes in Computer Science, 7424:45–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35734-3_3/.

Dejean:2012:UPB

[401] Hervé Déjean. Using page breaks for book structuring. Lecture Notes in Computer Science, 7424:57–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35734-3_4/.

Deveaud:2012:SRE

[402] Romain Deveaud, Eric SanJuan, and Patrice Bellot. Social recommendation and external resources for book search. Lecture Notes in Computer Science, 7424:68–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35734-3_5/.

Feild:2012:UMA

[403] Henry A. Feild, Marc-Allen Cartright, and James Allan. The University of Massachusetts Amherst’s participation in the INEX 2011 prove it track. Lecture Notes in Computer Science, 7424:80–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35734-3_6/.

Giguet:2012:BSE

[404] Emmanuel Giguet and Nadine Lucas. The book structure extraction competition with the resurgence full content software at Caen University. Lecture Notes in Computer Science, 7424:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35734-3_7/.

Liu:2012:TSE

[405] Caihua Liu, Jiajun Chen, Xiaofeng Zhang, Jie Liu, and Yalou Huang. TOC structure extraction from OCR-ed books. Lecture Notes in Computer Science, 7424:98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35734-3_8/.

Preminger:2012:OPI

[406] Michael Preminger and Ragnar Nordlie. OUC’s participation in the 2011 INEX book track. Lecture Notes in Computer Science, 7424: 109–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35734-3_9/.

Wang:2012:OID

[407] Qiuyue Wang, Georgina Ramírez, Maarten Marx, Martin Theobald, and Jaap Kamps. Overview of the INEX 2011 data-centric track. Lecture Notes in Computer Science, 7424:118–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35734-3_10/.

Laitang:2012:EDX

[408] Cyril Laitang, Karen Pinel-Sauvagnat, and Mohand Boughanem. Edit distance for XML information retrieval: Some experiments on the datacentric track of INEX 2011. Lecture Notes in Computer Science, 7424: 138–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35734-3_11/.

Ramirez:2012:UIB

[409] Georgina Ramírez. UPF at INEX 2011: Books and social search track and data-centric track. Lecture Notes in Computer Science, 7424:146– 154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35734-3_12/.

Schuth:2012:UAD

[410] Anne Schuth and Maarten Marx. University of Amsterdam data centric ad hoc and faceted search runs. Lecture Notes in Computer Science, 7424: 155–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35734-3_13/.

Ayala:2012:BRA

[411] Darnes Vilariño Ayala, David Pinto, Saúl León Silverio, and Esteban Castillo. BUAP: a recursive approach to the data-centric track of INEX 2011. Lecture Notes in Computer Science, 7424:161–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35734-3_14/.

Wang:2012:RDC

[412] Qiuyue Wang, Yantao Gan, and Yu Sun. RUC@INEX 2011 data-centric track. Lecture Notes in Computer Science, 7424:167–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35734-3_15/.

Wichaiwong:2012:MI

[413] Tanakorn Wichaiwong and Chuleerat Jaruskulchai. MEXIR at INEX-2011. Lecture Notes in Computer Science, 7424:180–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35734-3_16/.

SanJuan:2012:OIQ

[414] Eric SanJuan, Véronique Moriceau, Xavier Tannier, Patrice Bellot, and Josiane Mothe. Overview of the INEX 2011 question answering track (QA@INEX). Lecture Notes in Computer Science, 7424:188–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35734-3_17/.

Bhaskar:2012:HQS

[415] Pinaki Bhaskar, Somnath Banerjee, Snehasis Neogi, and Sivaji Bandyopadhyay. A hybrid QA system with focused IR and automatic summarization for INEX 2011. Lecture Notes in Computer Science, 7424: 207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35734-3_18/.

Ermakova:2012:IIQ

[416] Liana Ermakova and Josiane Mothe. IRIT at INEX: Question answering task. Lecture Notes in Computer Science, 7424:219–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35734-3_19/.

Laureano-Cruces:2012:GBS

[417] Ana Lilia Laureano-Cruces and Javier Ramírez-Rodríguez. A graphbased summarization system at QA@INEX track 2011. Lecture Notes in Computer Science, 7424:227–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35734-3_20/.

Anonymous:2012:FMu

[418] Anonymous. Front matter. Lecture Notes in Computer Science, 7424: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35734-3/1.

Leem:2012:ARA

[419] Intaek Leem, Mary Wu, and Chonggun Kim. An algorithm for resource allocation without inter-cluster collisions in WSN. Lecture Notes in Computer Science, 7425:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32645-5_1/.

Min:2012:DCT

[420] Joonki Min, Jookyoung Kim, and Youngmi Kwon. Data compression technique for wireless sensor networks. Lecture Notes in Computer Science, 7425:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32645-5_2/.

Cho:2012:BRM

[421] Jeong hyun Cho, Yong do Choi, and Sung ho Kim. A bandwidth reservation method for multicast streaming service. Lecture Notes in Computer Science, 7425:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32645-5_3/.

Kim:2012:HAR

[422] Dongho Kim. Hybrid-ARQ with rate compatible LDPC codes. Lecture Notes in Computer Science, 7425:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32645-5_4/.

Li:2012:ISI

[423] Shu Li and Jeong Geun Kim. Inter-session and intra-session network coding for wireless sensor networks. Lecture Notes in Computer Science, 7425:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32645-5_5/.

Kim:2012:QAE

[424] Seong Cheol Kim, Jun Heon Jeon, and Hyun Joo Park. QoS aware energyefficient (QAEE) MAC protocol for energy harvesting wireless sensor networks. Lecture Notes in Computer Science, 7425:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32645-5_6/.

Lee:2012:NCA

[425] Kyungrak Lee, Juhyun Maeng, and Inwhee Joe. A new channel access scheme for multi-user voice communications in IEEE 802.15.4 networks. Lecture Notes in Computer Science, 7425:49–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32645-5_7/.

Shin:2012:PEL

[426] Hyun-Ho Shin and Byung-Seo Kim. Performance evaluations on localrepair of AODV protocol over IP-based ad-hoc networks. Lecture Notes in Computer Science, 7425:57–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32645-5_8/.

Kusdaryono:2012:BSA

[427] Aries Kusdaryono, Gan Bayar, and Lee Kyoung Oh. Base station assisted hierarchical routing protocol in wireless sensor networks. Lecture Notes in Computer Science, 7425:64–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32645-5_9/.

Shin:2012:DCS

[428] Hee Jae Shin, Do Hyun Bae, and Jin Suk Kim. Derivation of coverage in sensor network using a quad-tree partitioning scheme. Lecture Notes in Computer Science, 7425:72–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32645-5_10/.

Hosen:2012:AMR

[429] A. S. M. Sanwar Hosen and Gihwan Cho. Applying a minimum routing cost spanning tree to many-to-many in wireless sensor networks. Lecture Notes in Computer Science, 7425:79–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32645-5_11/.

Imanishi:2012:ETC

[430] Ryota Imanishi, Yuki Kondo, Young-Bok Choi, and Hiromi Okada. An effective traffic control method for real-time and reliable bursts in OBS networks. Lecture Notes in Computer Science, 7425:87–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32645-5_12/.

Lee:2012:TPC

[431] Ye Hoon Lee. A transmit power control method for DS/CDMA communication systems using successive interference cancelation. Lecture Notes in Computer Science, 7425:98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32645-5_13/.

Kang:2012:IEA

[432] SeokHoon Kang, Gwanggil Jeon, and Young-Sup Lee. Improved energy aware routing protocol in mobile ad hoc network. Lecture Notes in Computer Science, 7425:106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32645-5_14/.

Bok:2012:ARF

[433] Junyeong Bok and Heung-Gyoon Ryu. Analysis of receiver feedback closed loop pre-distortion to compensate for nonlinear HPA and foliage attenuation. Lecture Notes in Computer Science, 7425:114–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32645-5_15/.

Tuan:2012:MPF

[434] Le Thanh Tuan, Daeseung Yoo, Hyungjoo Kim, Gwangja Jin, Byungtae Jang, and Soong Hwan Ro. The modified proportional fair packet scheduling algorithm for multimedia traffic in LTE system. Lecture Notes in Computer Science, 7425:122–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32645-5_16/.

Yoo:2012:HRF

[435] Kwang Kiun Yoo and Ronny Yongho Kim. HiNet: Radio frequency communication based train control (RF-CBTC) system jointly using hierarchical modulation and network coding. Lecture Notes in Computer Science, 7425:130–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32645-5_17/.

Bong:2012:FWF

[436] Wonsun Bong and Yong Cheol Kim. Fingerprint Wi–Fi radio map interpolated by discontinuity preserving smoothing. Lecture Notes in Computer Science, 7425:138–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32645-5_18/.

Lee:2012:CSK

[437] Ingi Lee and Hwan-Seung Yong. Common sense knowledge based hybrid interestingness measures for data mining. Lecture Notes in Computer Science, 7425:146–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32645-5_19/.

Noman:2012:MSC

[438] Nasimul Noman, Leon Palafox, and Hitoshi Iba. On model selection criteria in reverse engineering gene networks using RNN model. Lecture Notes in Computer Science, 7425:155–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32645-5_20/.

Anonymous:2012:FMv

[439] Anonymous. Front matter. Lecture Notes in Computer Science, 7425: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32645-5/1.

Vassie:2012:NAS

[440] Kelly Vassie and Giuseppe Morlino. Natural and artificial systems: Compare, model or engineer? Lecture Notes in Computer Science, 7426: 1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33093-3_1/.

Larsen:2012:LRC

[441] Jørgen Christian Larsen, David Brandt, and Kasper Stoy. LocoKit: a robot construction kit for studying and developing functional morphologies. Lecture Notes in Computer Science, 7426:12–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33093-3_2/.

Rano:2012:IAB

[442] Iñaki Rañó. An introduction to the analysis of Braitenberg vehicles 2 and 3 using phase plane portrait. Lecture Notes in Computer Science, 7426:23–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33093-3_3/.

Alessandro:2012:SAE

[443] Cristiano Alessandro, Juan Pablo Carbajal, and Andrea d’Avella. Synthesis and adaptation of effective motor synergies for the solution of reaching tasks. Lecture Notes in Computer Science, 7426:33–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33093-3_4/.

Nunez-Varela:2012:GAA

[444] Jose Nunez-Varela, Balaraman Ravindran, and Jeremy L. Wyatt. Gaze allocation analysis for a visually guided manipulation task. Lecture Notes in Computer Science, 7426:44–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33093-3_5/.

Hoffmann:2012:USC

[445] Matej Hoffmann, Nico M. Schmidt, Rolf Pfeifer, Andreas K. Engel, and Alexander Maye. Using sensorimotor contingencies for terrain discrimination and adaptive walking behavior in the quadruped robot puppy. Lecture Notes in Computer Science, 7426:54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33093-3_6/.

Grzyb:2012:HWI

[446] Beata J. Grzyb, Angel P. del Pobil, and Linda B. Smith. How walking influences the development of absolute distance perception. Lecture Notes in Computer Science, 7426:65–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33093-3_7/.

Duran:2012:MIM

[447] Boris Durán and Serge Thill. Modelling interaction in multi-modal affordance processing with neural dynamics. Lecture Notes in Computer Science, 7426:75–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33093-3_8/.

Silva:2012:BIM

[448] Ana Carolina Silva and Cristina Peixoto dos Santos. A bio-inspired model reliably predicts the collision of approaching objects under different light conditions. Lecture Notes in Computer Science, 7426:85– 95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33093-3_9/.

Kuppuswamy:2012:SMP

[449] Naveen Kuppuswamy, Hugo Gravato Marques, and Helmut Hauser. Synthesising a motor-primitive inspired control architecture for redundant compliant robots. Lecture Notes in Computer Science, 7426:96–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33093-3_10/.

Maye:2012:USC

[450] Alexander Maye and Andreas K. Engel. Using sensorimotor contingencies for prediction and action planning. Lecture Notes in Computer Science, 7426:106–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33093-3_11/.

Folgheraiter:2012:BIC

[451] Michele Folgheraiter, Mathias Jordan, Jan Albiez, and Frank Kirchner. A bio-inspired control system for a wearable human-machine interface. Lecture Notes in Computer Science, 7426:117–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33093-3_12/.

Ahn:2012:FEM

[452] Sejoon Ahn and DaeEun Kim. A finite element method of electric image in weakly electric fish. Lecture Notes in Computer Science, 7426: 127–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33093-3_13/.

Jauffret:2012:MIV

[453] Adrien Jauffret, Nicolas Cuperlier, Philippe Gaussier, and Philippe Tarroux. Multimodal integration of visual place cells and grid cells for navigation tasks of a real robot. Lecture Notes in Computer Science, 7426: 136–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33093-3_14/.

Waegeman:2012:TNH

[454] Tim Waegeman, Francis Wyffels, and Benjamin Schrauwen. Towards a neural hierarchy of time scales for motor control. Lecture Notes in Computer Science, 7426:146–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33093-3_15/.

Li:2012:MWB

[455] Cai Li, Robert Lowe, and Tom Ziemke. Modelling walking behaviors based on CPGs: a simplified bio-inspired architecture. Lecture Notes in Computer Science, 7426:156–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33093-3_16/.

Antonelli:2012:PRR

[456] Marco Antonelli, Beata J. Grzyb, Vicente Castelló, and Angel P. del Pobil. Plastic representation of the reachable space for a humanoid robot. Lecture Notes in Computer Science, 7426:167–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33093-3_17/.

Ollion:2012:TBC

[457] Charles Ollion and Stéphane Doncieux. Towards behavioral consistency in neuroevolution. Lecture Notes in Computer Science, 7426: 177–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33093-3_18/.

Rempis:2012:EVN

[458] Christian Rempis and Frank Pasemann. Evolving variants of neurocontrol using constraint masks. Lecture Notes in Computer Science, 7426: 187–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33093-3_19/.

Wrobel:2012:SBE

[459] Borys Wróbel, Michał Joachimczak, Alberto Montebelli, and Robert Lowe. The search for beauty: Evolution of minimal cognition in an animat controlled by a gene regulatory network and powered by a metabolic system. Lecture Notes in Computer Science, 7426:198–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33093-3_20/.

Anonymous:2012:FMw

[460] Anonymous. Front matter. Lecture Notes in Computer Science, 7426: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33093-3/1.

Lehmann:2012:ACA

[461] Sabine Lehmann, Ben Gottesman, Robert Grabowski, and Mayo Kudo. Applying CNL authoring support to improve machine translation of forum data. Lecture Notes in Computer Science, 7427:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32612-7_1/.

Ferre:2012:SCN

[462] Sébastien Ferré. SQUALL: a controlled natural language for querying and updating RDF graphs. Lecture Notes in Computer Science, 7427: 11–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32612-7_2/.

Schwitter:2012:ASP

[463] Rolf Schwitter. Answer set programming via controlled natural language processing. Lecture Notes in Computer Science, 7427:26–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32612-7_3/.

Power:2012:OSE

[464] Richard Power. OWL simplified english: a finite-state language for ontology editing. Lecture Notes in Computer Science, 7427:44–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32612-7_4/.

Murray:2012:SEC

[465] William R. Murray and Tomas Singliar. Spatiotemporal extensions to a controlled natural language. Lecture Notes in Computer Science, 7427: 61–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32612-7_5/.

Kaljurand:2012:CNL

[466] Kaarel Kaljurand and Tanel Alumäe. Controlled natural language in speech recognition based user interfaces. Lecture Notes in Computer Science, 7427:79–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32612-7_6/.

Hasibi:2012:ATG

[467] Faegheh Hasibi. An adaptation technique for GF-based dialogue systems. Lecture Notes in Computer Science, 7427:95–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32612-7_7/.

Kaljurand:2012:GAC

[468] Kaarel Kaljurand. General architecture of a controlled natural language based multilingual semantic Wiki. Lecture Notes in Computer Science, 7427:110–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32612-7_8/.

Gruzitis:2012:FRG

[469] Normunds Gruzitis, Peteris Paikens, and Guntis Barzdins. FrameNet resource grammar library for GF. Lecture Notes in Computer Science, 7427: 121–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32612-7_9/.

Hofler:2012:LDG

[470] Stefan Höfler. Legislative drafting guidelines: How different are they from controlled language rules for technical writing? Lecture Notes in Computer Science, 7427:138–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32612-7_10/.

Marrafa:2012:PCL

[471] Palmira Marrafa, Raquel Amaro, Nuno Freire, and Sara Mendes. Portuguese controlled language: Coping with ambiguity. Lecture Notes in Computer Science, 7427:152–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32612-7_11/.

Davis:2012:MVM

[472] Brian Davis, Ramona Enache, and Jeroen van Grondelle. Multilingual verbalisation of modular ontologies using GF and lemon. Lecture Notes in Computer Science, 7427:167–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32612-7_12/.

Anonymous:2012:BMh

[473] Anonymous. Back matter. Lecture Notes in Computer Science, 7427: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32612-7/1.

Anonymous:2012:FMx

[474] Anonymous. Front matter. Lecture Notes in Computer Science, 7427: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32612-7/1.

Briais:2012:HC

[475] Sébastien Briais, Stéphane Caron, Jean-Michel Cioranesco, and Jean-Luc Danger. 3D hardware canaries. Lecture Notes in Computer Science, 7428:1–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33027-8_1/.

Skorobogatov:2012:BSS

[476] Sergei Skorobogatov and Christopher Woods. Breakthrough silicon scanning discovers backdoor in military chip. Lecture Notes in Computer Science, 7428:23–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33027-8_2/.

Schlosser:2012:SPE

[477] Alexander Schlösser, Dmitry Nedospasov, Juliane Krämer, and Susanna Orlic. Simple photonic emission analysis of AES. Lecture Notes in Computer Science, 7428:41–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33027-8_3/.

Moss:2012:CAM

[478] Andrew Moss, Elisabeth Oswald, Dan Page, and Michael Tunstall. Compiler assisted masking. Lecture Notes in Computer Science, 7428:58– 75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33027-8_4/.

Bilgin:2012:TIA

[479] Begül Bilgin, Svetla Nikova, Ventzislav Nikov, Vincent Rijmen, and Georg Stütz. Threshold implementations of all 3 × 3 and 4 × 4 S-boxes. Lecture Notes in Computer Science, 7428:76–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33027-8_5/.

Moradi:2012:HFS

[480] Amir Moradi and Oliver Mischke. How far should theory be from practice? Lecture Notes in Computer Science, 7428:92–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33027-8_6/.

Debraize:2012:EPS

[481] Blandine Debraize. Efficient and provably secure methods for switching from arithmetic to Boolean masking. Lecture Notes in Computer Science, 7428:107–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33027-8_7/.

Banik:2012:DFA

[482] Subhadeep Banik, Subhamoy Maitra, and Santanu Sarkar. A differential fault attack on the grain family of stream ciphers. Lecture Notes in Computer Science, 7428:122–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33027-8_8/.

Oren:2012:ASC

[483] Yossef Oren, Mathieu Renauld, François-Xavier Standaert, and Avishai Wool. Algebraic side-channel attacks beyond the Hamming weight leakage model. Lecture Notes in Computer Science, 7428:140–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33027-8_9/.

Reparaz:2012:STS

[484] Oscar Reparaz, Benedikt Gierlichs, and Ingrid Verbauwhede. Selecting time samples for multivariate DPA attacks. Lecture Notes in Computer Science, 7428:155–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33027-8_10/.

Gerard:2012:UOL

[485] Benoît Gérard and François-Xavier Standaert. Unified and optimized linear collision attacks and their application in a non-profiled setting. Lecture Notes in Computer Science, 7428:175–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33027-8_11/.

Medwed:2012:TSE

[486] Marcel Medwed, François-Xavier Standaert, and Antoine Joux. Towards super-exponential side-channel security with efficient leakage-resilient PRFs. Lecture Notes in Computer Science, 7428:193–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33027-8_12/.

Faust:2012:PLR

[487] Sebastian Faust, Krzysztof Pietrzak, and Joachim Schipper. Practical leakage-resilient symmetric cryptography. Lecture Notes in Computer Science, 7428:213–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33027-8_13/.

Fei:2012:SMD

[488] Yunsi Fei, Qiasi Luo, and A. Adam Ding. A statistical model for DPA with novel algorithmic confusion analysis. Lecture Notes in Computer Science, 7428:233–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33027-8_14/.

Ruhrmair:2012:PSA

[489] Ulrich Rührmair and Marten van Dijk. Practical security analysis of PUFbased two-player protocols. Lecture Notes in Computer Science, 7428: 251–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33027-8_15/.

vanderLeest:2012:SDE

[490] Vincent van der Leest, Bart Preneel, and Erik van der Sluis. Soft decision error correction for compact memory-based PUFs using a single enrollment. Lecture Notes in Computer Science, 7428:268–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33027-8_16/.

Katzenbeisser:2012:PMF

[491] Stefan Katzenbeisser, Ünal Kocabaş, Vladimir Rožić, and Ahmad-Reza Sadeghi. PUFs: Myth, fact or busted? A security evaluation of physically unclonable Functions (PUFs) cast in silicon. Lecture Notes in Computer Science, 7428:283–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33027-8_17/.

Maes:2012:PFF

[492] Roel Maes, Anthony Van Herrewege, and Ingrid Verbauwhede. PUFKY: a fully functional PUF-based cryptographic key generator. Lecture Notes in Computer Science, 7428:302–319, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33027-8_18/.

Bernstein:2012:NC

[493] Daniel J. Bernstein and Peter Schwabe. NEON crypto. Lecture Notes in Computer Science, 7428:320–339, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33027-8_19/.

Heyse:2012:TOC

[494] Stefan Heyse and Tim Güneysu. Towards one cycle per bit asymmetric encryption: Code-based cryptography on reconfigurable hardware. Lecture Notes in Computer Science, 7428:340–355, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33027-8_20/.

Anonymous:2012:FMy

[495] Anonymous. Front matter. Lecture Notes in Computer Science, 7428: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33027-8/1.

Gordon:2012:CER

[496] Goren Gordon and Ehud Ahissar. A curious emergence of reaching. Lecture Notes in Computer Science, 7429:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32527-4_1/.

Dahl:2012:FDM

[497] Torbjørn S. Dahl and Alexandros Paraschos. A force-distance model of humanoid arm withdrawal reflexes. Lecture Notes in Computer Science, 7429:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32527-4_2/.

Mahyuddin:2012:NAC

[498] Muhammad Nasiruddin Mahyuddin, Guido Herrmann, and Said Ghani Khan. A novel adaptive control algorithm in application to a humanoid robot arm. Lecture Notes in Computer Science, 7429:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32527-4_3/.

Rano:2012:WBV

[499] Iñaki Rañó and Jean-Stephane Jokeit. A wandering Braitenberg vehicle 2b that densely covers a bounded workspace. Lecture Notes in Computer Science, 7429:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32527-4_4/.

Rajruangrabin:2012:AIM

[500] Jartuwat Rajruangrabin, Isura Ranatunga, and Dan O. Popa. Adaptive interface mapping for intuitive teleoperation of multi-DOF robots. Lecture Notes in Computer Science, 7429:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32527-4_5/.

Santos:2012:AST

[501] Carlos Santos, Manuel Mazo Jr., and Felipe Espinosa. Adaptive selftriggered control of a remotely operated robot. Lecture Notes in Computer Science, 7429:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32527-4_6/.

Ranasinghe:2012:AST

[502] Nadeesha Oliver Ranasinghe and Wei-Min Shen. Autonomous surveillance tolerant to interference. Lecture Notes in Computer Science, 7429: 73–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32527-4_7/.

Araiza-Illan:2012:BIA

[503] Dejanira Araiza-Illan and Tony J. Dodd. Bio-inspired autonomous navigation and escape from pursuers with potential functions. Lecture Notes in Computer Science, 7429:84–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32527-4_8/.

Harris:2012:BCB

[504] Christopher M. Harris and Jonathan Waddington. Biomimetics of choice behaviour for autonomous agents. Lecture Notes in Computer Science, 7429:96–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32527-4_9/.

Broun:2012:BKM

[505] Alan Broun, Chris Beck, Tony Pipe, Majid Mirmehdi, and Chris Melhuish. Building a kinematic model of a Robot’s arm with a depth camera. Lecture Notes in Computer Science, 7429:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32527-4_10/.

Kochuvila:2012:CCG

[506] Sreeja Kochuvila, Shikha Tripathi, and Sudarshan T. S. B. Control of a compass gait biped robot based on partial feedback linearization. Lecture Notes in Computer Science, 7429:117–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32527-4_11/.

Vaussard:2012:CEC

[507] Florian Vaussard, Philippe Rétornaz, David Hamel, and Francesco Mondada. Cutting down the energy consumed by domestic robots: Insights from robotic vacuum cleaners. Lecture Notes in Computer Science, 7429: 128–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32527-4_12/.

Andersen:2012:DCE

[508] Nils Axel Andersen and Ole Ravn. Designing competitions for education in robotics. Lecture Notes in Computer Science, 7429:140–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32527-4_13/.

Solanes:2012:DRN

[509] J. Ernesto Solanes, Leopoldo Armesto, Josep Tornero, Pau Muñoz-Benavent, and Vicent Girbés. Dual-rate non-linear high order holds for visual servoing applications. Lecture Notes in Computer Science, 7429: 152–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32527-4_14/.

Byrne:2012:ELS

[510] S. Byrne, W. Naeem, and R. S. Ferguson. Efficient local sampling for motion planning of a robotic manipulator. Lecture Notes in Computer Science, 7429:164–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32527-4_15/.

Tritton:2012:ERW

[511] Terry Tritton, Joanna Hall, Angela Rowe, Sophie Valentine, and Alicja Jedrzejewska. Engaging with robots while giving simple instructions. Lecture Notes in Computer Science, 7429:176–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32527-4_16/.

Aufderheide:2012:ELM

[512] Dominik Aufderheide, Werner Krybus, and Ulf Witkowski. Experiences with LEGO MINDSTORMS as an embedded and robotics plattform within the undergraduate curriculum. Lecture Notes in Computer Science, 7429:185–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32527-4_17/.

Nikandrova:2012:ESB

[513] Ekaterina Nikandrova, Janne Laaksonen, and Ville Kyrki. Explorative sensor-based grasp planning. Lecture Notes in Computer Science, 7429: 197–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32527-4_18/.

Iverach-Brereton:2012:ISH

[514] Chris Iverach-Brereton, Andrew Winton, and Jacky Baltes. Ice skating humanoid robot. Lecture Notes in Computer Science, 7429:209– 219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32527-4_19/.

Burbridge:2012:LGM

[515] Chris Burbridge and Richard Dearden. Learning the geometric meaning of symbolic abstractions for manipulation planning. Lecture Notes in Computer Science, 7429:220–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32527-4_20/.

Anonymous:2012:FMz

[516] Anonymous. Front matter. Lecture Notes in Computer Science, 7429: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32527-4/1.

Resconi:2012:IAS

[517] Germano Resconi and Chris Hinde. Introduction to active sets and unification. Lecture Notes in Computer Science, 7430:1–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34645-3_1/.

Oommen:2012:MTT

[518] B. John Oommen and M. Khaled Hashem. Modeling a teacher in a tutorial- like system using learning automata. Lecture Notes in Computer Science, 7430:37–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34645-3_2/.

Horak:2012:PNR

[519] Aleš Horák, Miroslav Prýmek, and Tadeusz Sikora. Power network reliability computations using multi-agent simulation. Lecture Notes in Computer Science, 7430:63–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34645-3_3/.

Piech:2012:SAR

[520] Henryk Piech. Sequence automata for researching consensus levels. Lecture Notes in Computer Science, 7430:82–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34645-3_4/.

Takahashi:2012:AIF

[521] Hiroshi Takahashi. An analysis of the influence of fundamental indexing on financial markets through agent-based modeling: The fundamentalist and fundamental indexing. Lecture Notes in Computer Science, 7430: 102–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34645-3_5/.

Kiv:2012:ADI

[522] Sodany Kiv, Yves Wautelet, and Manuel Kolp. Agent-driven integration architecture for component-based software development. Lecture Notes in Computer Science, 7430:121–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34645-3_6/.

Latham:2012:ATI

[523] Annabel Latham and Keeley Crockett. Adaptive tutoring in an intelligent conversational agent system. Lecture Notes in Computer Science, 7430: 148–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34645-3_7/.

Treur:2012:BCP

[524] Jan Treur. Biological and computational perspectives on the emergence of social phenomena: Shared understanding and collective power. Lecture Notes in Computer Science, 7430:168–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34645-3_8/.

Szymanezyk:2012:ABC

[525] Olivier Szymanezyk and Tom Duckett. Agent-based crowd simulation in airports using games technology. Lecture Notes in Computer Science, 7430:192–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34645-3_9/.

Maleszka:2012:AAS

[526] Marcin Maleszka and Ngoc Thanh Nguyen. Approximate algorithms for solving O_{1} consensus problems using complex tree structure. Lecture Notes in Computer Science, 7430:214–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34645-3_10/.

Anonymous:2012:BMi

[527] Anonymous. Back matter. Lecture Notes in Computer Science, 7430: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34645-3/1.

Anonymous:2012:FMba

[528] Anonymous. Front matter. Lecture Notes in Computer Science, 7430: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34645-3/1.

Bano:2012:SAW

[529] J. Bano, A. Hostettler, S. A. Nicolau, C. Doignon, H. S. Wu, M. H. Huang, and L. Soler. Simulation of the abdominal wall and its arteries after pneumoperitoneum for guidance of port positioning in laparoscopic surgery. Lecture Notes in Computer Science, 7431:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33179-4_1/.

Masutani:2012:ASF

[530] Yoshitaka Masutani, Mitsutaka Nemoto, Shohei Hanaoka, Naoto Hayashi, and Kuni Ohtomo. Appearance similarity flow for quantification of anatomical landmark uncertainty in medical images. Lecture Notes in Computer Science, 7431:12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33179-4_2/.

Wei:2012:SBT

[531] Zhenwen Wei, Caiming Zhang, Xingqiang Yang, and Xiaofeng Zhang. Segmentation of brain tumors in CT images using level sets. Lecture Notes in Computer Science, 7431:22–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33179-4_3/.

Bakas:2012:FLL

[532] Spyridon Bakas, Andreas Hoppe, Katerina Chatzimichail, and Vasileios Galariotis. Focal liver lesion tracking in CEUS for characterisation based on dynamic behaviour. Lecture Notes in Computer Science, 7431: 32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33179-4_4/.

Hajiesmaeili:2012:SHD

[533] Maryam Hajiesmaeili, Bashir Bagherinakhjavanlo, Jamshid Dehmeshki, and Tim Ellis. Segmentation of the hippocampus for detection of Alzheimer’s disease. Lecture Notes in Computer Science, 7431:42–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33179-4_5/.

Asarnow:2012:SPH

[534] Daniel Asarnow and Rahul Singh. Segmentation of parasites for highcontent screening using phase congruency and grayscale morphology. Lecture Notes in Computer Science, 7431:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33179-4_6/.

Ye:2012:MNB

[535] Jian Ye, Igor Yanovsky, Bin Dong, Rima Gandlin, Achi Brandt, and Stanley Osher. Multigrid narrow band surface reconstruction via level set functions. Lecture Notes in Computer Science, 7431:61–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33179-4_7/.

Chen:2012:RTS

[536] Xiao Chen, Guangming Wang, Ying Zhu, and G. Scott Owen. Real-time simulation of ship motions in waves. Lecture Notes in Computer Science, 7431:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33179-4_8/.

Blanchette:2012:ASM

[537] Damon Blanchette and Emmanuel Agu. Adaptive spectral mapping for real-time dispersive refraction. Lecture Notes in Computer Science, 7431: 81–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33179-4_9/.

Feng:2012:DMC

[538] Powei Feng and Joe Warren. A dual method for constructing multimaterial solids from ray-reps. Lecture Notes in Computer Science, 7431: 92–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33179-4_10/.

Sedlacek:2012:UDR

[539] David Sedlacek and Jiri Zara. User driven 3D reconstruction environment. Lecture Notes in Computer Science, 7431:104–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33179-4_11/.

Amresh:2012:MAL

[540] Ashish Amresh, John Femiani, and Christoph Fünfzig. Methods for approximating loop subdivision using tessellation enabled GPUs. Lecture Notes in Computer Science, 7431:115–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33179-4_12/.

Liu:2012:BAC

[541] Kun Liu and Rhaleb Zayer. Bundle adjustment constrained smoothing for multi-view point cloud data. Lecture Notes in Computer Science, 7431: 126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33179-4_13/.

Rastgar:2012:GSM

[542] Houman Rastgar, Eric Dubois, and Liang Zhang. Guided sampling in multiple view robust motion estimation using regression diagnostics. Lecture Notes in Computer Science, 7431:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33179-4_14/.

Doliotis:2012:HSP

[543] Paul Doliotis, Vassilis Athitsos, Dimitrios Kosmopoulos, and Stavros Perantonis. Hand shape and 3D pose estimation using depth data from a single cluttered frame. Lecture Notes in Computer Science, 7431: 148–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33179-4_15/.

Fortenbury:2012:FLR

[544] Billy Ray Fortenbury and Gutemberg Guerra-Filho. Fusing low-resolution depth maps into high-resolution stereo matching. Lecture Notes in Computer Science, 7431:159–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33179-4_16/.

Galego:2012:ACP

[545] Ricardo Galego, Alexandre Bernardino, and José Gaspar. Autocalibration of pan-tilt cameras including radial distortion and zoom. Lecture Notes in Computer Science, 7431:169–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33179-4_17/.

Fortenbury:2012:RSC

[546] Billy Ray Fortenbury and Gutemberg Guerra-Filho. Robust 2D/3D calibration using RANSAC registration. Lecture Notes in Computer Science, 7431:179–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33179-4_18/.

Catano:2012:KDB

[547] M. A. Cataño and J. Climent. Keypoint detection based on the unimodality test of HOGs. Lecture Notes in Computer Science, 7431: 189–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33179-4_19/.

Li:2012:NRP

[548] Bo Li, Afzal Godil, and Henry Johan. Non-rigid and partial 3D model retrieval using hybrid shape descriptor and meta similarity. Lecture Notes in Computer Science, 7431:199–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33179-4_20/.

Anonymous:2012:FMbb

[549] Anonymous. Front matter. Lecture Notes in Computer Science, 7431: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33179-4/1.

Hofbauer:2012:IRI

[550] Heinz Hofbauer, Christian Rathgeb, Andreas Uhl, and Peter Wild. Iris recognition in image domain: Quality-metric based comparators. Lecture Notes in Computer Science, 7432:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33191-6_1/.

Sedmidubsky:2012:GRB

[551] Jan Sedmidubsky, Jakub Valcik, Michal Balazia, and Pavel Zezula. Gait recognition based on normalized walk cycles. Lecture Notes in Computer Science, 7432:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33191-6_2/.

Kim:2012:INS

[552] Hwi-Gang Kim, Eun Jung Lee, Gang-Joon Yoon, Sung-Dae Yang, Eui Chul Lee, and Sang Min Yoon. Illumination normalization for SIFT based finger vein authentication. Lecture Notes in Computer Science, 7432:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33191-6_3/.

Kotsia:2012:HRS

[553] Irene Kotsia, Weiwei Guo, and Ioannis Patras. Higher rank support tensor machines. Lecture Notes in Computer Science, 7432:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33191-6_4/.

Choi:2012:MSI

[554] Inho Choi and Daijin Kim. Multi-scale integral modified census transform for eye detection. Lecture Notes in Computer Science, 7432:41– 50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33191-6_5/.

Wesley:2012:CAT

[555] Avinash Wesley, Pradeep Buddharaju, Robert Pienta, and Ioannis Pavlidis. A comparative analysis of thermal and visual modalities for automated facial expression recognition. Lecture Notes in Computer Science, 7432:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33191-6_6/.

Nait-Charif:2012:VTL

[556] Hammadi Nait-Charif, Allen Breen, and Paul Thompson. Vertebrae tracking in lumbar spinal video-fluoroscopy using particle filters with semi-automatic initialisation. Lecture Notes in Computer Science, 7432: 61–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33191-6_7/.

Panin:2012:MIM

[557] Giorgio Panin. Mutual information for multi-modal, discontinuitypreserving image registration. Lecture Notes in Computer Science, 7432: 70–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33191-6_8/.

Hussain:2012:MDD

[558] Muhammad Hussain, Salabat Khan, Ghulam Muhammad, and George Bebis. Mass detection in digital mammograms using optimized Gabor filter bank. Lecture Notes in Computer Science, 7432:82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33191-6_9/.

Sukno:2012:CDL

[559] Federico M. Sukno, John L. Waddington, and Paul F. Whelan. Comparing 3D descriptors for local search of craniofacial landmarks. Lecture Notes in Computer Science, 7432:92–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33191-6_10/.

Saha:2012:VBT

[560] Utsab Saha and Rahul Singh. Vision-based tracking of complex macroparasites for high-content phenotypic drug screening. Lecture Notes in Computer Science, 7432:104–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33191-6_11/.

DeVylder:2012:CND

[561] Jonas De Vylder, Jan Aelterman, Mado Vandewoestyne, Trees Lepez, and Dieter Deforce. Cell nuclei detection using globally optimal active contours with shape prior. Lecture Notes in Computer Science, 7432: 115–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33191-6_12/.

Kolawole:2012:NGR

[562] Akintola Kolawole and Alireza Tavakkoli. A novel gait recognition system based on hidden Markov models. Lecture Notes in Computer Science, 7432:125–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33191-6_13/.

Karali:2012:MHS

[563] Abubakrelsedik Karali and Mohamed ElHelw. Motion history of skeletal volumes for human action recognition. Lecture Notes in Computer Science, 7432:135–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33191-6_14/.

Yoon:2012:CM

[564] Sang Min Yoon and Gang-Joon Yoon. Compressive matting. Lecture Notes in Computer Science, 7432:145–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33191-6_15/.

Yatagawa:2012:TBC

[565] Tatsuya Yatagawa and Yasushi Yamaguchi. A template-based completion framework for videos with dynamic backgrounds. Lecture Notes in Computer Science, 7432:155–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33191-6_16/.

Azary:2012:ACU

[566] Sherif Azary and Andreas Savakis. 3D action classification using sparse spatio-temporal feature representations. Lecture Notes in Computer Science, 7432:166–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33191-6_17/.

Kamberov:2012:SDA

[567] George Kamberov, Matt Burlick, Lazaros Karydas, and Olga Koteoglou. SCAR: Dynamic adaptation for person detection and persistence analysis in unconstrained videos. Lecture Notes in Computer Science, 7432: 176–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33191-6_18/.

Galanopoulos:2012:EDR

[568] Georgios Galanopoulos, Constantin Papaodysseus, Dimitiris Arabadjis, and Michael Exarhos. Exploiting 3D digital representations of ancient inscriptions to identify their writer. Lecture Notes in Computer Science, 7432:188–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33191-6_19/.

Alsam:2012:WED

[569] Ali Alsam, Hans Jakob Rivertz, and Puneet Sharma. What the eye did not see — a fusion approach to image coding. Lecture Notes in Computer Science, 7432:199–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33191-6_20/.

Anonymous:2012:FMbc

[570] Anonymous. Front matter. Lecture Notes in Computer Science, 7432: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33191-6/1.

Behsaz:2012:TUS

[571] Bahar Behsaz, Ján Maňuch, and Ladislav Stacho. Turing universality of step-wise and stage assembly at temperature 1. Lecture Notes in Computer Science, 7433:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32208-2_1/.

Brijder:2012:TSD

[572] Robert Brijder and Joris J. M. Gillis. A type system for DNAQL. Lecture Notes in Computer Science, 7433:12–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32208-2_2/.

Chen:2012:DFC

[573] Ho-Lin Chen, David Doty, and David Soloveichik. Deterministic function computation with chemical reaction networks. Lecture Notes in Computer Science, 7433:25–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32208-2_3/.

Condon:2012:RBC

[574] Anne Condon, Bonnie Kirkpatrick, and Ján Maňuch. Reachability bounds for chemical reaction networks and Strand displacement systems. Lecture Notes in Computer Science, 7433:43–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32208-2_4/.

Czeizler:2012:SMT

[575] Eugen Czeizler and Alexandru Popa. Synthesizing minimal tile sets for complex patterns in the framework of patterned DNA self-assembly. Lecture Notes in Computer Science, 7433:58–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32208-2_5/.

Garzon:2012:GAG

[576] Max H. Garzon and Kiran C. Bobba. A geometric approach to Gibbs energy landscapes and optimal DNA codeword design. Lecture Notes in Computer Science, 7433:73–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32208-2_6/.

Kan:2012:DBM

[577] Anton Kan, Koh ichiroh Shohda, and Akira Suyama. A DNA based molecular logic gate capable of a variety of logical operations. Lecture Notes in Computer Science, 7433:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32208-2_7/.

Kari:2012:DWR

[578] Lila Kari and Steffen Kopecki. Deciding whether a regular language is generated by a splicing system. Lecture Notes in Computer Science, 7433: 98–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32208-2_8/.

deMurieta:2012:PRB

[579] Iñaki Sainz de Murieta and Alfonso Rodríguez-Patón. Probabilistic reasoning with a Bayesian DNA device based on Strand displacement. Lecture Notes in Computer Science, 7433:110–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32208-2_9/.

Svaneborg:2012:DSA

[580] Carsten Svaneborg and Harold Fellermann. DNA self-assembly and computation studied with a coarse-grained dynamic bonded model. Lecture Notes in Computer Science, 7433:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32208-2_10/.

Thachuk:2012:SEE

[581] Chris Thachuk and Anne Condon. Space and energy efficient computation with DNA Strand displacement systems. Lecture Notes in Computer Science, 7433:135–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32208-2_11/.

Anonymous:2012:BMj

[582] Anonymous. Back matter. Lecture Notes in Computer Science, 7433: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32208-2/1.

Anonymous:2012:FMbd

[583] Anonymous. Front matter. Lecture Notes in Computer Science, 7433: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32208-2/1.

Bhattacharya:2012:LTA

[584] Binay Bhattacharya and Tsunehiko Kameda. A linear time algorithm for computing minmax regret 1-median on a tree. Lecture Notes in Computer Science, 7434:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32241-9_1/.

Jaiswal:2012:SSB

[585] Ragesh Jaiswal, Amit Kumar, and Sandeep Sen. A simple D^{2}-sampling based PTAS for k-means and other clustering problems. Lecture Notes in Computer Science, 7434:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32241-9_2/.

Bampis:2012:SSM

[586] Evripidis Bampis, Dimitrios Letsios, and Ioannis Milis. Speed scaling for maximum lateness. Lecture Notes in Computer Science, 7434: 25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32241-9_3/.

Floderus:2012:ISI

[587] Peter Floderus, Mirosław Kowaluk, and Andrzej Lingas. Induced subgraph isomorphism: Are some patterns substantially easier than others? Lecture Notes in Computer Science, 7434:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32241-9_4/.

Bezakova:2012:CMS

[588] Ivona Bezáková and Zachary Langley. Contiguous minimum single-source-multi-sink cuts in weighted planar graphs. Lecture Notes in Computer Science, 7434:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32241-9_5/.

Han:2012:OKP

[589] Xin Han, Yasushi Kawase, and Kazuhisa Makino. Online knapsack problem with removal cost. Lecture Notes in Computer Science, 7434: 61–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32241-9_6/.

Xiao:2012:IEA

[590] Mingyu Xiao and Hiroshi Nagamochi. An improved exact algorithm for TSP in degree-4 graphs. Lecture Notes in Computer Science, 7434: 74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32241-9_7/.

Rue:2012:DPM

[591] Juanjo Rué, Ignasi Sau, and Dimitrios M. Thilikos. Dynamic programming for H-minor-free graphs. Lecture Notes in Computer Science, 7434: 86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32241-9_8/.

Mastrolilli:2012:RMM

[592] Monaldo Mastrolilli and Georgios Stamoulis. Restricted max-min fair allocations with inclusion-free intervals. Lecture Notes in Computer Science, 7434:98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32241-9_9/.

Babenko:2012:IAP

[593] Maxim A. Babenko, Kamil Salikhov, and Stepan Artamonov. An improved algorithm for packing T-paths in inner Eulerian networks. Lecture Notes in Computer Science, 7434:109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32241-9_10/.

vanBevern:2012:TOE

[594] René van Bevern. Towards optimal and expressive kernelization for d-hitting set. Lecture Notes in Computer Science, 7434:121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32241-9_11/.

Couturier:2012:MNM

[595] Jean-François Couturier, Pinar Heggernes, and Pim van ’t Hof. Maximum number of minimal feedback vertex sets in chordal graphs and cographs. Lecture Notes in Computer Science, 7434:133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-32241-9_12/.

Peng:2012:LAF

[596] Pan Peng. A local algorithm for finding dense bipartite-like subgraphs. Lecture Notes in Computer Science, 7434:145–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32241-9_13/.

Kloks:2012:ASC

[597] Ton Kloks, Sheung-Hung Poon, Chin-Ting Ung, and Yue-Li Wang. Algorithms for the strong chromatic index of Halin graphs, distance-hereditary graphs and maximal outerplanar graphs. Lecture Notes in Computer Science, 7434:157–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32241-9_14/.

Wang:2012:MDH

[598] Biing-Feng Wang and Chih-Hsuan Li. On the minimum degree hypergraph problem with subset size two and the red-blue set cover problem with the consecutive ones property. Lecture Notes in Computer Science, 7434: 169–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32241-9_15/.

Chandran:2012:RCS

[599] L. Sunil Chandran and Deepak Rajendraprasad. Rainbow colouring of split and threshold graphs. Lecture Notes in Computer Science, 7434: 181–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32241-9_16/.

Popa:2012:ARB

[600] Alexandru Popa. Approximating the rainbow — better lower and upper bounds. Lecture Notes in Computer Science, 7434:193–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32241-9_17/.

Belmonte:2012:RNL

[601] Rémy Belmonte, Pinar Heggernes, Pim van ’t Hof, and Reza Saei. Ramsey numbers for line graphs and perfect graphs. Lecture Notes in Computer Science, 7434:204–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32241-9_18/.

Aichholzer:2012:GOT

[602] Oswin Aichholzer, Matias Korman, and Alexander Pilz. Geodesic order types. Lecture Notes in Computer Science, 7434:216–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32241-9_19/.

Durocher:2012:CPR

[603] Stephane Durocher and Saeed Mehrabi. Computing partitions of rectilinear polygons with minimum stabbing number. Lecture Notes in Computer Science, 7434:228–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32241-9_20/.

Anonymous:2012:FMbe

[604] Anonymous. Front matter. Lecture Notes in Computer Science, 7434: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32241-9/1.

Alam:2012:ABC

[605] Mohammad Shafiul Alam, Md. Monirul Islam, and Kazuyuki Murase. Artificial bee colony algorithm with improved explorations for numerical function optimization. Lecture Notes in Computer Science, 7435: 1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32639-4_1/.

Zangari:2012:EAM

[606] Murilo Zangari, Wesley Romão, and Ademir Aparecido Constantino. Extensions of ant-miner algorithm to deal with class imbalance problem. Lecture Notes in Computer Science, 7435:9–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-32639-4_2/.

Mehboob:2012:MPA

[607] Zareen Mehboob, Hujun Yin, Sophie M. Wuerger, and Laura M. Parkes. Multivoxel pattern analysis using information-preserving EMD. Lecture Notes in Computer Science, 7435:19–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32639-4_3/.

Parpinelli:2012:PRU

[608] Rafael Stubs Parpinelli and Heitor Silvério Lopes. Population resizing using nonlinear dynamics in an ecology-based approach. Lecture Notes in Computer Science, 7435:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32639-4_4/.

Fuangkhon:2012:MCC

[609] Piyabute Fuangkhon and Thitipong Tanprasert. Multi-class contour preserving classification. Lecture Notes in Computer Science, 7435:35–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32639-4_5/.

Davis:2012:ABA

[610] Luke M. Davis, Barry-John Theobald, and Anthony Bagnall. Automated bone age assessment using feature extraction. Lecture Notes in Computer Science, 7435:43–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32639-4_6/.

Silva:2012:HAB

[611] Washington Silva and Ginalber Serra. A hybrid approach based on DCTgenetic-fuzzy inference system for speech recognition. Lecture Notes in Computer Science, 7435:52–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32639-4_7/.

Shi:2012:ARC

[612] Yong Shi and Brian Graham. An approach to reshaping clusters for nearest neighbor search. Lecture Notes in Computer Science, 7435: 60–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32639-4_8/.

Hills:2012:IMF

[613] Jon Hills, Luke M. Davis, and Anthony Bagnall. Interestingness measures for fixed consequent rules. Lecture Notes in Computer Science, 7435:68–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32639-4_9/.

Huo:2012:AED

[614] Jing Huo, Yang Gao, Wanqi Yang, and Hujun Yin. Abnormal event detection via multi-instance dictionary learning. Lecture Notes in Computer Science, 7435:76–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32639-4_10/.

Junior:2012:SPR

[615] Luneque Silva Junior, Nadia Nedjah, and Luiza de Macedo Mourelle. Static packet routing in NoC platform using ACO-based algorithms. Lecture Notes in Computer Science, 7435:84–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32639-4_11/.

R:2012:SMA

[616] Cordeiro F. R., Lima S. M. L., Silva-Filho A. G., and Santos W. P. Segmentation of mammography by applying extreme learning machine in tumor detection. Lecture Notes in Computer Science, 7435: 92–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32639-4_12/.

dosSantos:2012:ASO

[617] Tiago Henrique dos Santos, Alessandro Goedtel, and Sérgio Augusto Oliveira da Silva. An ANN speed observer applied to three-phase induction motor. Lecture Notes in Computer Science, 7435:101–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32639-4_13/.

Fernandez-de-Alba:2012:OSI

[618] José M. Fernández de Alba, Pablo Campillo, Rubén Fuentes-Fernández, and Juan Pavón. Opportunistic sensor interpretation in a virtual smart environment. Lecture Notes in Computer Science, 7435:109–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32639-4_14/.

Jacomini:2012:CPA

[619] Ricardo de Souza Jacomini, Marcelo Zanchetta do Nascimento, and Rogério Daniel Dantas. Comparison of PCA and ANOVA for information selection of CC and MLO views in classification of mammograms. Lecture Notes in Computer Science, 7435:117–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32639-4_15/.

Veroneze:2012:MAA

[620] Rosana Veroneze, André R. Gonçalves, and Fernando J. Von Zuben. A multiobjective analysis of adaptive clustering algorithms for the definition of RBF neural network centers in regression problems. Lecture Notes in Computer Science, 7435:127–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32639-4_16/.

dosSantos:2012:MMT

[621] Michel M. dos Santos, Mêuser J. S. Valença, and Wellington P. dos Santos. Mean multiclass Type I and II errors for training multilayer perceptron with particle swarm in image segmentation. Lecture Notes in Computer Science, 7435:135–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32639-4_17/.

Chidambaram:2012:IAA

[622] Chidambaram Chidambaram, Marlon Subtil Marçal, Leyza Baldo Dorini, and Hugo Vieira Neto. An improved ABC algorithm approach using SURF for face identification. Lecture Notes in Computer Science, 7435: 143–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32639-4_18/.

Costa:2012:MBG

[623] Ariadne A. Costa, Antonio C. Roque, Silvio Morato, and Renato Tinós. A model based on genetic algorithm for investigation of the behavior of rats in the elevated plus-maze. Lecture Notes in Computer Science, 7435: 151–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32639-4_19/.

Esmin:2012:DCU

[624] Ahmed A. A. Esmin and Stan Matwin. Data clustering using hybrid particle swarm optimization. Lecture Notes in Computer Science, 7435: 159–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32639-4_20/.

Anonymous:2012:FMbf

[625] Anonymous. Front matter. Lecture Notes in Computer Science, 7435: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32639-4/1.

Abadi:2012:SSF

[626] Martín Abadi. Software security: a formal perspective. Lecture Notes in Computer Science, 7436:1–5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32759-9_1/.

Degani:2012:FMW

[627] Asaf Degani. Formal methods in the wild: Trains, planes, & automobile. Lecture Notes in Computer Science, 7436:6, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-32759-9_2.

Wassyng:2012:WWW

[628] Alan Wassyng. Who are we, and what are we doing here? Lecture Notes in Computer Science, 7436:7–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32759-9_3/.

Aarts:2012:ALT

[629] Fides Aarts, Faranak Heidarian, Harco Kuppens, and Petur Olsen. Automata learning through counterexample guided abstraction refinement. Lecture Notes in Computer Science, 7436:10–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32759-9_4/.

Alpuente:2012:JTS

[630] María Alpuente, Demis Ballis, Francisco Frechina, and Daniel Romero. Julienne: a trace slicer for conditional rewrite theories. Lecture Notes in Computer Science, 7436:28–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32759-9_5/.

Andre:2012:ITA

[631] Étienne André, Laurent Fribourg, Ulrich Kühne, and Romain Soulat. IMITATOR 2.5: a tool for analyzing robustness in scheduling problems. Lecture Notes in Computer Science, 7436:33–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32759-9_6/.

Aponte:2012:MCP

[632] Virginia Aponte, Pierre Courtieu, Yannick Moy, and Marc Sango. Maximal and compositional pattern-based loop invariants. Lecture Notes in Computer Science, 7436:37–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32759-9_7/.

Asplund:2012:FAA

[633] Mikael Asplund, Atif Manzoor, Mélanie Bouroche, and Siobhàn Clarke. A formal approach to autonomous vehicle coordination. Lecture Notes in Computer Science, 7436:52–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32759-9_8/.

Barringer:2012:QEA

[634] Howard Barringer, Yliès Falcone, Klaus Havelund, and Giles Reger. Quantified event automata: Towards expressive and efficient runtime monitors. Lecture Notes in Computer Science, 7436:68–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32759-9_9/.

Bauer:2012:DLM

[635] Andreas Bauer and Yliès Falcone. Decentralised LTL monitoring. Lecture Notes in Computer Science, 7436:85–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32759-9_10/.

Benkirane:2012:MEP

[636] Soufiene Benkirane, Rachel Norman, Erin Scott, and Carron Shankland. Measles epidemics and PEPA: An exploration of historic disease dynamics using process algebra. Lecture Notes in Computer Science, 7436: 101–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32759-9_11/.

Carlier:2012:CCS

[637] Matthieu Carlier, Catherine Dubois, and Arnaud Gotlieb. A certified constraint solver over finite domains. Lecture Notes in Computer Science, 7436:116–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32759-9_12/.

Christakis:2012:CVT

[638] Maria Christakis, Peter Müller, and Valentin Wüstholz. Collaborative verification and testing with explicit assumptions. Lecture Notes in Computer Science, 7436:132–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32759-9_13/.

Cousineau:2012:TP

[639] Denis Cousineau, Damien Doligez, Leslie Lamport, and Stephan Merz. TLA^{+} proofs. Lecture Notes in Computer Science, 7436:147–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32759-9_14/.

DIppolito:2012:MTS

[640] Nicolás D’Ippolito, Victor Braberman, and Nir Piterman. The modal transition system control problem. Lecture Notes in Computer Science, 7436:155–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32759-9_15/.

Dormoy:2012:WSR

[641] Julien Dormoy, Olga Kouchnarenko, and Arnaud Lanoix. When structural refinement of components keeps temporal properties over reconfigurations. Lecture Notes in Computer Science, 7436:171–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32759-9_16/.

Ermis:2012:EI

[642] Evren Ermis, Martin Schäf, and Thomas Wies. Error invariants. Lecture Notes in Computer Science, 7436:187–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32759-9_17/.

Giorgino:2012:CPM

[643] Mathieu Giorgino and Martin Strecker. Correctness of pointer manipulating algorithms illustrated by a verified BDD construction. Lecture Notes in Computer Science, 7436:202–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32759-9_18/.

Heather:2012:FFM

[644] James Heather and Steve Schneider. A formal framework for modelling coercion resistance and receipt freeness. Lecture Notes in Computer Science, 7436:217–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32759-9_19/.

Hierons:2012:UTA

[645] Robert M. Hierons, Mercedes G. Merayo, and Manuel Núñez. Using time to add order to distributed testing. Lecture Notes in Computer Science, 7436:232–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32759-9_20/.

Anonymous:2012:FMbg

[646] Anonymous. Front matter. Lecture Notes in Computer Science, 7436: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32759-9/1.

Abid:2012:RTS

[647] Nouha Abid and Silvano Dal Zilio. Real-time specification patterns and tools. Lecture Notes in Computer Science, 7437:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32469-7_1/.

Albert:2012:AEA

[648] Elvira Albert and Bjarte M. Østvold. Automated extraction of abstract behavioural models from JMS applications. Lecture Notes in Computer Science, 7437:16–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32469-7_2/.

Ayache:2012:CRC

[649] Nicolas Ayache and Roberto M. Amadio. Certifying and reasoning on cost annotations in C programs. Lecture Notes in Computer Science, 7437:32–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32469-7_3/.

Baier:2012:WLH

[650] Christel Baier, Marcus Daum, and Benjamin Engel. Waiting for locks: How Long does it usually take? Lecture Notes in Computer Science, 7437:47–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32469-7_4/.

Bandur:2012:MAS

[651] Victor Bandur, Wolfram Kahl, and Alan Wassyng. Microcontroller assembly synthesis from timed automaton task specifications. Lecture Notes in Computer Science, 7437:63–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32469-7_5/.

Barnat:2012:TCS

[652] Jiri Barnat, Jan Beran, and Lubos Brim. Tool chain to support automated formal verification of avionics Simulink designs. Lecture Notes in Computer Science, 7437:78–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32469-7_6/.

Barrett:2012:RAB

[653] Edd Barrett and Andy King. Range analysis of binaries with minimal effort. Lecture Notes in Computer Science, 7437:93–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32469-7_7/.

Correnson:2012:CAC

[654] Loïc Correnson and Julien Signoles. Combining analyses for C program verification. Lecture Notes in Computer Science, 7437:108–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32469-7_8/.

Cranen:2012:MCF

[655] Sjoerd Cranen. Model checking the FlexRay startup phase. Lecture Notes in Computer Science, 7437:131–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32469-7_9/.

Droste:2012:MBR

[656] Rainer Droste and Christoph Läsche. Model-based risk assessment supporting development of HSE plans for safe offshore operations. Lecture Notes in Computer Science, 7437:146–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32469-7_10/.

Ferrucci:2012:MAV

[657] Luca Ferrucci and Dino Mandrioli. Modular automated verification of flexible manufacturing systems with metric temporal logic and nonstandard analysis. Lecture Notes in Computer Science, 7437:162–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32469-7_11/.

Popa:2012:ORS

[658] Vlad Popa and Wolfgang Schwitzer. Optimizing the robustness of software against communication latencies in distributed reactive embedded systems. Lecture Notes in Computer Science, 7437:177–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32469-7_12/.

Siek:2012:FDT

[659] Konrad Siek and Paweł T. Wojciechowski. A formal design of a tool for static analysis of upper bounds on object calls in Java. Lecture Notes in Computer Science, 7437:192–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32469-7_13/.

Slaby:2012:CPD

[660] Jiří Slabý, Jan Strejček, and Marek Trtík. Checking properties described by state machines: On synergy of instrumentation, slicing, and symbolic execution. Lecture Notes in Computer Science, 7437:207–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32469-7_14/.

Anonymous:2012:BMk

[661] Anonymous. Back matter. Lecture Notes in Computer Science, 7437: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32469-7/1.

Anonymous:2012:FMbh

[662] Anonymous. Front matter. Lecture Notes in Computer Science, 7437: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32469-7/1.

Kowalski:2012:LBF

[663] Robert Kowalski and Fariba Sadri. A logic-based framework for reactive systems. Lecture Notes in Computer Science, 7438:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32689-9_1/.

Mugnier:2012:OBQ

[664] Marie-Laure Mugnier. Ontology-based query answering with existential rules. Lecture Notes in Computer Science, 7438:16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32689-9_2.

Cosentino:2012:MDR

[665] Valerio Cosentino, Jordi Cabot, Patrick Albert, Philippe Bauquel, and Jacques Perronnet. A model driven reverse engineering framework for extracting business rules out of a Java application. Lecture Notes in Computer Science, 7438:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32689-9_3/.

Hashmi:2012:BPD

[666] Mustafa Hashmi, Guido Governatori, and Moe Thandar Wynn. Business process data compliance. Lecture Notes in Computer Science, 7438: 32–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32689-9_4/.

Silva:2012:FBR

[667] Bruno Berstel-Da Silva. Formalizing both refraction-based and sequential executions of production rule programs. Lecture Notes in Computer Science, 7438:47–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32689-9_5/.

ElGhali:2012:BOO

[668] Adil El Ghali, Amina Chniti, and Hugues Citeau. Bringing OWL ontologies to the business rules users. Lecture Notes in Computer Science, 7438:62–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32689-9_6/.

Guisse:2012:RTB

[669] Abdooulaye Guissé, François Lévy, and Adeline Nazarenko. From regulatory texts to BRMS: How to guide the acquisition of business rules? Lecture Notes in Computer Science, 7438:77–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32689-9_7/.

Malik:2012:RBA

[670] Saleem Malik and Imran Sarwar Bajwa. A rule based approach for business rule generation from business process models. Lecture Notes in Computer Science, 7438:92–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32689-9_8/.

Paschke:2012:RRS

[671] Adrian Paschke, Harold Boley, Zhili Zhao, Kia Teymourian, and Tara Athan. Reaction RuleML 1.0: Standardized semantic reaction rules. Lecture Notes in Computer Science, 7438:100–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32689-9_9/.

Patkos:2012:PRB

[672] Theodore Patkos, Abdelghani Chibani, Dimitris Plexousakis, and Yacine Amirat. A production rule-based framework for causal and epistemic reasoning. Lecture Notes in Computer Science, 7438:120–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32689-9_10/.

Ksystra:2012:ASR

[673] Katerina Ksystra, Nikolaos Triantafyllou, and Petros Stefaneas. On the algebraic semantics of reactive rules. Lecture Notes in Computer Science, 7438:136–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32689-9_11/.

Bragaglia:2012:RBC

[674] Stefano Bragaglia, Federico Chesani, Paola Mello, and Davide Sottara. A rule-based calculus and processing of complex events. Lecture Notes in Computer Science, 7438:151–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32689-9_12/.

Costantini:2012:CRP

[675] Stefania Costantini and Giovanni De Gasperis. Complex reactivity with preferences in rule-based agents. Lecture Notes in Computer Science, 7438: 167–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32689-9_13/.

Weigand:2012:PWP

[676] Hans Weigand and Adrian Paschke. The pragmatic web: Putting rules in context. Lecture Notes in Computer Science, 7438:182–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32689-9_14/.

Kravari:2012:HHR

[677] Kalliopi Kravari and Nick Bassiliades. HARM: a hybrid rule-based agent reputation model based on temporal defeasible logic. Lecture Notes in Computer Science, 7438:193–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32689-9_15/.

Jailly:2012:WIE

[678] Benjamin Jailly, Christophe Gravier, Julien Subercaze, Marius Preda, and Jacques Fayolle. Whispering interactions to the end user using rules. Lecture Notes in Computer Science, 7438:208–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32689-9_16/.

Viktoratos:2012:PLI

[679] Iosif Viktoratos, Athanasios Tsadiras, and Nick Bassiliades. Personalizing location information through rule-based policies. Lecture Notes in Computer Science, 7438:215–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32689-9_17/.

Demey:2012:UDF

[680] Yan Tang Demey and Trung-Kien Tran. Using SOIQ(D) to formalize semantics within a semantic decision table. Lecture Notes in Computer Science, 7438:224–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32689-9_18/.

Batsakis:2012:IRT

[681] Sotiris Batsakis and Euripides G. M. Petrakis. Imposing restrictions over temporal properties in OWL: a rule-based approach. Lecture Notes in Computer Science, 7438:240–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32689-9_19/.

Almendros-Jimenez:2012:ORL

[682] Jesús M. Almendros-Jiménez. OWL RL in logic programming: Querying, reasoning and inconsistency explanations. Lecture Notes in Computer Science, 7438:248–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32689-9_20/.

Anonymous:2012:FMbi

[683] Anonymous. Front matter. Lecture Notes in Computer Science, 7438: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32689-9/1.

Nishida:2012:ADP

[684] Kazufumi Nishida, Koji Nakano, and Yasuaki Ito. Accelerating the dynamic programming for the optimal polygon triangulation on the GPU. Lecture Notes in Computer Science, 7439:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33078-0_1/.

Huang:2012:SCR

[685] Xu Huang, Dharmendra Sharma, and Muhammad Ahmed. Security computing for the resiliency of protecting from internal attacks in distributed wireless sensor networks. Lecture Notes in Computer Science, 7439:16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33078-0_2/.

Lin:2012:PAN

[686] Shin-Yeu Lin and Xian-Chang Guo. Parallel algorithm for nonlinear network optimization problems and real-time applications. Lecture Notes in Computer Science, 7439:30–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33078-0_3/.

Schneider:2012:OSR

[687] Max Schneider, Nikola Belic, Christoph Sambale, Ulrich Hofmann, and Dietmar Fey. Optimization of a short-range proximity effect correction algorithm in E-beam lithography using GPGPUs. Lecture Notes in Computer Science, 7439:41–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33078-0_4/.

Alachiotis:2012:EMG

[688] Nikolaos Alachiotis, Pavlos Pavlidis, and Alexandros Stamatakis. Exploiting multi-grain parallelism for efficient selective sweep detection. Lecture Notes in Computer Science, 7439:56–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33078-0_5/.

Marinho:2012:VAQ

[689] Eraldo P. Marinho and Alexandro Baldassin. Vectorized algorithms for quadtree construction and descent. Lecture Notes in Computer Science, 7439:69–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33078-0_6/.

Nguyen:2012:MLM

[690] Vu Thien Nga Nguyen, Raimund Kirner, and Frank Penczek. A multi-level monitoring framework for stream-based coordination programs. Lecture Notes in Computer Science, 7439:83–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33078-0_7/.

Nakano:2012:OPP

[691] Koji Nakano. An optimal parallel prefix-sums algorithm on the memory machine models for GPUs. Lecture Notes in Computer Science, 7439: 99–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33078-0_8/.

Schaetz:2012:MGP

[692] Sebastian Schaetz and Martin Uecker. A multi-GPU programming library for real-time applications. Lecture Notes in Computer Science, 7439: 114–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33078-0_9/.

Venugopalan:2012:OLP

[693] Sarad Venugopalan and Oliver Sinnen. Optimal linear programming solutions for multiprocessor scheduling with communication delays. Lecture Notes in Computer Science, 7439:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33078-0_10/.

Ichinomiya:2012:BRT

[694] Yoshihiro Ichinomiya, Motoki Amagasaki, Masahiro Iida, and Morihiro Kuga. A bitstream relocation technique to improve flexibility of partial reconfiguration. Lecture Notes in Computer Science, 7439:139–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33078-0_11/.

Wang:2012:HHG

[695] Chuan Wang, Jianhua Gu, Yunlan Wang, and Tianhai Zhao. A hybrid heuristic-genetic algorithm for task scheduling in heterogeneous multi-core system. Lecture Notes in Computer Science, 7439:153–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33078-0_12/.

Wang:2012:ETA

[696] Li Wang, Jing Liu, Jingtong Hu, Qingfeng Zhuge, Duo Liu, and Edwin H.-M. Sha. Efficient task assignment on heterogeneous multicore systems considering communication overhead. Lecture Notes in Computer Science, 7439:171–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33078-0_13/.

Caron:2012:BCR

[697] Eddy Caron, Frédéric Desprez, Adrian Muresan, and Frédéric Suter. Budget constrained resource allocation for non-deterministic workflows on an IaaS cloud. Lecture Notes in Computer Science, 7439:186–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33078-0_14/.

Nakahara:2012:WFF

[698] Hiroki Nakahara, Hiroyuki Nakanishi, and Tsutomu Sasao. On a wideband fast Fourier transform using piecewise linear approximations: Application to a radio telescope spectrometer. Lecture Notes in Computer Science, 7439:202–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33078-0_15/.

Bousdira:2012:VLA

[699] Wadoud Bousdira, Frédéric Loulergue, and Julien Tesson. A verified library of algorithmic skeletons on evenly distributed arrays. Lecture Notes in Computer Science, 7439:218–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33078-0_16/.

Umeda:2012:PMP

[700] Takayuki Umeda and Keiichiro Fukazawa. Performance measurement of parallel Vlasov code for space plasma on various scalar-type supercomputer systems. Lecture Notes in Computer Science, 7439:233–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33078-0_17/.

Haigh:2012:USC

[701] Andrew A. Haigh, Bradley E. Treeby, and Eric C. McCreath. Ultrasound simulation on the Cell Broadband Engine using the Westervelt equation. Lecture Notes in Computer Science, 7439:241–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33078-0_18/.

Yao:2012:SDF

[702] Nianmin Yao, Yong Han, Shaobin Cai, and Qilong Han. Study on the data flow balance in NFS server with iSCSI. Lecture Notes in Computer Science, 7439:253–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33078-0_19/.

Kirsch:2012:PSS

[703] Christoph M. Kirsch, Hannes Payer, Harald Röck, and Ana Sokolova. Performance, scalability, and semantics of concurrent FIFO queues. Lecture Notes in Computer Science, 7439:273–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33078-0_20/.

Anonymous:2012:FMbj

[704] Anonymous. Front matter. Lecture Notes in Computer Science, 7439: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33078-0/1.

Chiu:2012:AMM

[705] Jih-Ching Chiu, Kai-Ming Yang, and Chen-Ang Wong. Analytical modeling for multi-transaction bus on distributed systems. Lecture Notes in Computer Science, 7440:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33065-0_1/.

Shu:2012:LST

[706] LihChyun Shu, Ying-Cheng Su, Chang-Ming Tasi, and Huey-Min Sun. Leveraging the strengths of transactional memory while maintaining system performance for a multiplayer gaming application. Lecture Notes in Computer Science, 7440:10–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33065-0_2/.

Kim:2012:EPD

[707] Il-Young Kim and Jong-Kook Kim. Enhancing the performance of a distributed mobile computing environment by topology construction. Lecture Notes in Computer Science, 7440:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33065-0_3/.

Dai:2012:FER

[708] Yi Dai, Ke fei Wang, Wei xia Xu, He ying Zhang, and Shao gang Wang. Frame error rate testing for high speed optical interconnect. Lecture Notes in Computer Science, 7440:31–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33065-0_4/.

Atoofian:2012:MCS

[709] Ehsan Atoofian and Amir Ghanbari Bavarsad. Maintaining consistency in software transactional memory through dynamic versioning tuning. Lecture Notes in Computer Science, 7440:40–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33065-0_5/.

Chen:2012:SCP

[710] Da-Ren Chen, Chiun-Chieh Hsu, You-Shyang Chen, and Jun-Fu Guo. Solving a 2-covered path problem with variable radii for wireless sensor networks. Lecture Notes in Computer Science, 7440:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33065-0_6/.

Ohnishi:2012:DAD

[711] Yoshimasa Ohnishi and Takaichi Yoshida. Design of an applicationdependent static-based shared memory network. Lecture Notes in Computer Science, 7440:62–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33065-0_7/.

Nguyen:2012:DQB

[712] Anh P. Nguyen and Thuc D. Nguyen. Determining quality of S-boxes using pseudo random sequences generated from stream ciphers. Lecture Notes in Computer Science, 7440:72–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33065-0_8/.

Ji:2012:EOL

[713] Weixing Ji, Yizhuo Wang, Zhi Huang, Junqing Zhao, and Xi Li. Exploring object-level parallelism on chip multi-processors. Lecture Notes in Computer Science, 7440:80–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33065-0_9/.

Wang:2012:HWR

[714] Yi-Chun Wang and Justie Su-Tzu Juan. The Hamiltonicity of WKrecursive pyramid. Lecture Notes in Computer Science, 7440:90–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33065-0_10/.

Dib:2012:TML

[715] Djawida Dib, Nikos Parlavantzas, and Christine Morin. Towards multilevel adaptation for distributed operating systems and applications. Lecture Notes in Computer Science, 7440:100–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33065-0_11/.

Wei:2012:FCA

[716] Chih-Chiang Wei. FIDs classifier for artificial intelligence and its application. Lecture Notes in Computer Science, 7440:110–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33065-0_12/.

Noor:2012:AAF

[717] Ahmad Shukri Mohd Noor, Mustafa Mat Deris, Tutut Herawan, and Mohamad Nor Hassan. On affirmative adaptive failure detection. Lecture Notes in Computer Science, 7440:120–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33065-0_13/.

Lee:2012:NLL

[718] Wen-Ta Lee, Min-Sheng Chang, and Wei-Chieh Shen. A new low latency parallel turbo decoder employing parallel phase decoding method. Lecture Notes in Computer Science, 7440:130–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33065-0_14/.

Chedid:2012:RGQ

[719] Fouad B. Chedid. A regular group quorum system of degree ⌈\sqrt{n/2}⌉. Lecture Notes in Computer Science, 7440:140–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33065-0_15/.

Chedid:2012:NDO

[720] Fouad B. Chedid. A note on developing optimal and scalable parallel two-list algorithms. Lecture Notes in Computer Science, 7440: 148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33065-0_16/.

Peng:2012:SBO

[721] Kai Peng, Hua Zou, Rongheng Lin, and Fangchun Yang. Small businessoriented index construction of cloud data. Lecture Notes in Computer Science, 7440:156–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33065-0_17/.

Liu:2012:HPM

[722] Jie Liu, Lihua Chi, Chunye Gong, Han Xu, Jie Jiang, Yihui Yan, and Qingfeng Hu. High-performance matrix multiply on a massively multithreaded fiteng1000 processor. Lecture Notes in Computer Science, 7440: 166–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33065-0_18/.

Yu:2012:BBP

[723] Dongjin Yu, Wei Wu, Suhang Zheng, and Zhixiang Zhu. BIDE-based parallel mining of frequent closed sequences with MapReduce. Lecture Notes in Computer Science, 7440:177–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33065-0_19/.

Kim:2012:ETS

[724] Sung Il Kim, Jong-Kook Kim, Hyoung Uk Ha, Tae Ho Kim, and Kyu Hyun Choi. Efficient task scheduling for hard real-time tasks in asymmetric multicore processors. Lecture Notes in Computer Science, 7440: 187–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33065-0_20/.

Anonymous:2012:FMbk

[725] Anonymous. Front matter. Lecture Notes in Computer Science, 7440: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33065-0/1.

Thomas:2012:OAB

[726] A. Thomas and B. John Oommen. Optimal “anti-Bayesian” parametric pattern classification using order statistics criteria. Lecture Notes in Computer Science, 7441:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33275-3_1/.

Fischer:2012:IRB

[727] Asja Fischer and Christian Igel. An introduction to restricted Boltzmann machines. Lecture Notes in Computer Science, 7441:14–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33275-3_2/.

Liu:2012:HAR

[728] Zicheng Liu. Human activity recognition with 2D and 3D cameras. Lecture Notes in Computer Science, 7441:37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-33275-3_3.

Taubin:2012:SSD

[729] Gabriel Taubin. Smooth signed distance surface reconstruction and applications. Lecture Notes in Computer Science, 7441:38–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33275-3_4/.

Wachs:2012:RPM

[730] Juan P. Wachs. Robot, pass me the scissors! how robots can assist us in the operating Room. Lecture Notes in Computer Science, 7441: 46–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33275-3_5/.

Rodriguez-Serrano:2012:PRT

[731] José Antonio Rodriguez-Serrano. Pattern recognition in transportation. Lecture Notes in Computer Science, 7441:58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-33275-3_6.

Bottino:2012:IDA

[732] Andrea Bottino and Aldo Laurentini. The intrinsic dimensionality of attractiveness: a study in face profiles. Lecture Notes in Computer Science, 7441:59–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33275-3_7/.

Jaiswal:2012:HPC

[733] Ajay Jaiswal, Nitin Kumar, and R. K. Agrawal. A hybrid of principal component analysis and partial least squares for face recognition across pose. Lecture Notes in Computer Science, 7441:67–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33275-3_8/.

Ramon-Balmaseda:2012:GCL

[734] Enrique Ramón-Balmaseda, Javier Lorenzo-Navarro, and Modesto Castrillón-Santana. Gender classification in large databases. Lecture Notes in Computer Science, 7441:74–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33275-3_9/.

Castrillon-Santana:2012:CFF

[735] Modesto Castrillón-Santana, Daniel Hernández-Sosa, and Javier Lorenzo-Navarro. Combining face and facial feature detectors for face detection performance improvement. Lecture Notes in Computer Science, 7441: 82–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33275-3_10/.

Andres:2012:RFR

[736] Ariel Morelli Andres, Sebastian Padovani, Mariano Tepper, Marta Mejail, and Julio Jacobo. Randomized face recognition on partially occluded images. Lecture Notes in Computer Science, 7441:90–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33275-3_11/.

Hernandez:2012:FRW

[737] Noslen Hernández, Yoanna Martínez-Díaz, Dania Porro-Muñoz, and Heydi Méndez-Vázquez. Face recognition: Would going back to functional nature be a good idea? Lecture Notes in Computer Science, 7441: 98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33275-3_12/.

Martinez-Diaz:2012:DRB

[738] Yoanna Martínez-Díaz, Heydi Méndez-Vázquez, and Yenisel Plasencia-Calaña. Dissimilarity representations based on multi-block LBP for face detection. Lecture Notes in Computer Science, 7441:106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33275-3_13/.

Gomez-Barrero:2012:VIB

[739] Marta Gomez-Barrero, Javier Galbally, Pedro Tome, and Julian Fierrez. On the vulnerability of iris-based systems to a software attack based on a genetic algorithm. Lecture Notes in Computer Science, 7441: 114–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33275-3_14/.

Gonzalez:2012:RKB

[740] Fabio A. González, David Bermeo, Laura Ramos, and Olfa Nasraoui. On the robustness of kernel-based clustering. Lecture Notes in Computer Science, 7441:122–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33275-3_15/.

Peluffo-Ordonez:2012:IMC

[741] Diego Hernán Peluffo-Ordóñez and Carlos Daniel Acosta-Medina. An improved multi-class spectral clustering based on normalized cuts. Lecture Notes in Computer Science, 7441:130–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33275-3_16/.

Ahumada:2012:SHM

[742] Hernán C. Ahumada and Pablo M. Granitto. A simple hybrid method for semi-supervised learning. Lecture Notes in Computer Science, 7441: 138–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33275-3_17/.

Ryazanov:2012:CID

[743] Vladimir V. Ryazanov. Clustering of incomplete data and evaluation of clustering quality. Lecture Notes in Computer Science, 7441:146– 153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33275-3_18/.

Duval-Poo:2012:NCC

[744] Miguel A. Duval-Poo, Joan Sosa-García, Alejandro Guerra-Gandón, and Sandro Vega-Pons. A new classifier combination scheme using clustering ensemble. Lecture Notes in Computer Science, 7441:154–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33275-3_19/.

Duarte-Villasenor:2012:NDB

[745] Miriam Mónica Duarte-Villaseñor and Jesús Ariel Carrasco-Ochoa. Nested dichotomies based on clustering. Lecture Notes in Computer Science, 7441:162–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33275-3_20/.

Anonymous:2012:FMbl

[746] Anonymous. Front matter. Lecture Notes in Computer Science, 7441: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33275-3/1.

Abramov:2012:PSL

[747] S. A. Abramov and M. Petkovšek. On polynomial solutions of linear partial differential and (q-)difference equations. Lecture Notes in Computer Science, 7442:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32973-9_1/.

Ananth:2012:ACR

[748] Prabhanjan Ananth and Ambedkar Dukkipati. An algebraic characterization of rainbow connectivity. Lecture Notes in Computer Science, 7442: 12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32973-9_2/.

Batkhin:2012:AMA

[749] Alexander Batkhin. Application of the method of asymptotic solution to one multi-parameter problem. Lecture Notes in Computer Science, 7442:22–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32973-9_3/.

Bodrato:2012:NAL

[750] Marco Bodrato and Alberto Zanoni. A new algorithm for Long integer cube computation with some insight into higher powers. Lecture Notes in Computer Science, 7442:34–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32973-9_4/.

Bourdykine:2012:LAM

[751] Pavel Bourdykine and Stephen M. Watt. Lightweight abstraction for mathematical computation in Java. Lecture Notes in Computer Science, 7442:47–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32973-9_5/.

Bruno:2012:CNF

[752] Alexander D. Bruno and Victor F. Edneral. Calculation of normal forms of the Euler–Poisson equations. Lecture Notes in Computer Science, 7442: 60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32973-9_6/.

Budzko:2012:SEP

[753] Dzmitry A. Budzko and Alexander N. Prokopenya. Stability of equilibrium positions in the spatial circular restricted four-body problem. Lecture Notes in Computer Science, 7442:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32973-9_7/.

Errami:2012:CHB

[754] Hassan Errami and Werner M. Seiler. Computing Hopf bifurcations in chemical reaction networks using reaction coordinates. Lecture Notes in Computer Science, 7442:84–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32973-9_8/.

Gerdt:2012:CIS

[755] Vladimir Gerdt and Amir Hashemi. Comprehensive involutive systems. Lecture Notes in Computer Science, 7442:98–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32973-9_9/.

Giesbrecht:2012:PTA

[756] Mark Giesbrecht and Albert Heinle. A polynomial-time algorithm for the Jacobson form of a matrix of Ore polynomials. Lecture Notes in Computer Science, 7442:117–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32973-9_10/.

Gine:2012:RCP

[757] Jaume Giné, Colin Christopher, and Mateja Prešern. The resonant center problem for a 2:-3 resonant cubic Lotka–Volterra system. Lecture Notes in Computer Science, 7442:129–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32973-9_11/.

Grigoriev:2012:CSS

[758] Dima Grigoriev and Andreas Weber. Complexity of solving systems with few independent monomials and applications to mass-action kinetics. Lecture Notes in Computer Science, 7442:143–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32973-9_12/.

Gusev:2012:SNC

[759] Alexander Gusev and Sergue Vinitsky. Symbolic-numerical calculations of high-|m| Rydberg states and decay rates in strong magnetic fields. Lecture Notes in Computer Science, 7442:155–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32973-9_13/.

Hashemi:2012:QSV

[760] Amir Hashemi and Michael Schweinfurter. Quasi-stability versus genericity. Lecture Notes in Computer Science, 7442:172–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32973-9_14/.

Kemper:2012:ITA

[761] Gregor Kemper. Invariant theory: Applications and computations. Lecture Notes in Computer Science, 7442:185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-32973-9_15.

Li:2012:LGP

[762] Jia Li, Jin-San Cheng, and Elias P. Tsigaridas. Local generic position for root isolation of zero-dimensional triangular polynomial systems. Lecture Notes in Computer Science, 7442:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32973-9_16/.

Marcus:2012:FAC

[763] Steffen Marcus, Marc Moreno Maza, and Paul Vrbik. On Fulton’s algorithm for computing intersection multiplicities. Lecture Notes in Computer Science, 7442:198–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32973-9_17/.

Mezzarobba:2012:NSC

[764] Marc Mezzarobba. A note on the space complexity of fast D-finite function evaluation. Lecture Notes in Computer Science, 7442:212–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32973-9_18/.

Maza:2012:IMZ

[765] Marc Moreno Maza, Éric Schost, and Paul Vrbik. Inversion modulo zerodimensional regular chains. Lecture Notes in Computer Science, 7442: 224–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32973-9_19/.

Monagan:2012:SPP

[766] Michael Monagan and Roman Pearce. Sparse polynomial powering using heaps. Lecture Notes in Computer Science, 7442:236–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32973-9_20/.

Anonymous:2012:FMbm

[767] Anonymous. Front matter. Lecture Notes in Computer Science, 7442: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32973-9/1.

Wahid:2012:CSR

[768] Fathul Wahid. The current state of research on eGovernment in developing countries: a literature review. Lecture Notes in Computer Science, 7443:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33489-4_1/.

Nurdin:2012:IEI

[769] Nurdin Nurdin, Rosemary Stockdale, and Helana Scheepers. The influence of external institutional pressures on local E-government adoption and implementation: a coercive perspective within an Indonesian local E-government context. Lecture Notes in Computer Science, 7443: 13–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33489-4_2/.

vandenBoer:2012:ISC

[770] Yvon van den Boer, Lidwien van de Wijngaert, Willem Pieterson, and Rex Arendsen. On the interaction of source and channel choice in the government-to-business context. Lecture Notes in Computer Science, 7443:27–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33489-4_3/.

Alawadhi:2012:BUS

[771] Suha Alawadhi, Armando Aldama-Nalda, Hafedh Chourabi, J. Ramon Gil-Garcia, and Sofia Leung. Building understanding of smart city initiatives. Lecture Notes in Computer Science, 7443:40–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33489-4_4/.

Maarse:2012:NAL

[772] Nicole Maarse and Marijn Janssen. The need to adjust lean to the public sector. Lecture Notes in Computer Science, 7443:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33489-4_5/.

Gidlund:2012:DDD

[773] Katarina L. Gidlund. ‘demand driven development of public e-services’. Lecture Notes in Computer Science, 7443:66–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33489-4_6/.

Falcioni:2012:IPB

[774] Damiano Falcioni, Andrea Polini, Alberto Polzonetti, and Barbara Re. Improving PA business processes through modeling, analysis, and reengineering. Lecture Notes in Computer Science, 7443:78–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33489-4_7/.

Zuiderwijk:2012:IGP

[775] Anneke Zuiderwijk, Marijn Janssen, Ronald Meijer, Sunil Choenni, and Yannis Charalabidis. Issues and guiding principles for opening governmental judicial research data. Lecture Notes in Computer Science, 7443: 90–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33489-4_8/.

Panagiotopoulos:2012:DAH

[776] Panagiotis Panagiotopoulos, Alinaghi Ziaee Bigdeli, and Steven Sams. ‘5 Days in August’ — how London local authorities used Twitter during the 2011 Riots. Lecture Notes in Computer Science, 7443:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33489-4_9/.

Scannell:2012:SSI

[777] Mark Scannell and Frank Bannister. Shared services in Irish local government. Lecture Notes in Computer Science, 7443:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33489-4_10/.

Andrade:2012:ITE

[778] André Andrade and Luiz Antonio Joia. Information technology and the efficiency of the Brazilian judiciary system. Lecture Notes in Computer Science, 7443:126–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33489-4_11/.

deVries:2012:TBI

[779] Walter T. de Vries and Gianluca Miscione. The tool that has to build itself: The case of Dutch geo-data. Lecture Notes in Computer Science, 7443: 137–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33489-4_12/.

Zwattendorfer:2012:CBL

[780] Bernd Zwattendorfer, Arne Tauber, Klaus Stranacher, and Peter Reichstädter. Cross-border legal identity management. Lecture Notes in Computer Science, 7443:149–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33489-4_13/.

Lemmetti:2012:UEA

[781] Juha Lemmetti and Samuli Pekkola. Understanding enterprise architecture: Perceptions by the Finnish public sector. Lecture Notes in Computer Science, 7443:162–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33489-4_14/.

Jansen:2012:UIP

[782] Arild Jansen. The understanding of ICTs in public sector and its impact on governance. Lecture Notes in Computer Science, 7443:174–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33489-4_15/.

vanVeenstra:2012:IOG

[783] Anne Fleur van Veenstra and Marijn Janssen. Investigating outcomes of T-government using a public value management approach. Lecture Notes in Computer Science, 7443:187–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33489-4_16/.

Saebo:2012:GTC

[784] Øystein Sæbø. E-government in Tanzania: Current status and future challenges. Lecture Notes in Computer Science, 7443:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33489-4_17/.

Solar:2012:MAO

[785] Mauricio Solar, Gastón Concha, and Luis Meijueiro. A model to assess open government data in public agencies. Lecture Notes in Computer Science, 7443:210–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33489-4_18/.

Krogstie:2012:CPP

[786] John Krogstie. Comparing private and public sector on information systems development and maintenance efficiency. Lecture Notes in Computer Science, 7443:222–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33489-4_19/.

Stanimirovic:2012:AME

[787] Dalibor Stanimirovic, Tina Jukic, Janja Nograsek, and Mirko Vintar. Analysis of the methodologies for evaluation of E-government policies. Lecture Notes in Computer Science, 7443:234–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33489-4_20/.

Anonymous:2012:FMbn

[788] Anonymous. Front matter. Lecture Notes in Computer Science, 7443: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33489-4/1.

Dahlgren:2012:SMC

[789] Peter Dahlgren. Social media and counter-democracy: The contingences of participation. Lecture Notes in Computer Science, 7444:1– 12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33250-0_1/.

Albrecht:2012:CRC

[790] Steffen Albrecht. E-consultations: a review of current practice and a proposal for opening up the process. Lecture Notes in Computer Science, 7444:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33250-0_2/.

Johannessen:2012:CRM

[791] Marius Rohde Johannessen, Leif Skiftenes Flak, and Øystein Sæbø. Choosing the right medium for municipal eParticipation based on stakeholder expectations. Lecture Notes in Computer Science, 7444:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33250-0_3/.

Gronlund:2012:CGP

[792] Åke Grönlund and Iryna Susha. A communication genre perspective on epetitioning: The case of the citizens’ initiative. Lecture Notes in Computer Science, 7444:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33250-0_4/.

Panopoulou:2012:CEO

[793] Eleni Panopoulou, Efpraxia Dalakiouridou, Efthimios Tambouris, and Konstantinos Tarabanis. Citizens’ evaluation of an online argument visualisation platform for eParticipation. Lecture Notes in Computer Science, 7444:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33250-0_5/.

Benn:2012:PVA

[794] Neil Benn and Ann Macintosh. PolicyCommons — visualizing arguments in policy consultation. Lecture Notes in Computer Science, 7444: 61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33250-0_6/.

Shiramatsu:2012:SJR

[795] Shun Shiramatsu, Robin M. E. Swezey, Hiroyuki Sano, Norifumi Hirata, and Tadachika Ozono. Structuring Japanese regional information gathered from the Web as linked open data for use in concern assessment. Lecture Notes in Computer Science, 7444:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33250-0_7/.

Aichholzer:2012:UOC

[796] Georg Aichholzer, Doris Allhutter, and Stefan Strauß. Using online carbon calculators for participation in local climate initiatives. Lecture Notes in Computer Science, 7444:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33250-0_8/.

Scherer:2012:RPM

[797] Sabrina Scherer and Maria A. Wimmer. Reference process model for participatory budgeting in Germany. Lecture Notes in Computer Science, 7444:97–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33250-0_9/.

Glassey:2012:GTP

[798] Olivier Glassey and Camille-Angelo Aglione. Getting teenagers to participate: a survey of the Youth Council of the city of Lausanne. Lecture Notes in Computer Science, 7444:112–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33250-0_10/.

Wyner:2012:MBC

[799] Adam Z. Wyner, Katie Atkinson, and Trevor Bench-Capon. Model based critique of policy proposals. Lecture Notes in Computer Science, 7444: 120–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33250-0_11/.

Klinger:2012:OCS

[800] Roman Klinger, Philipp Senger, Sumit Madan, and Michal Jacovi. Online communities support policy-making: The need for data analysis. Lecture Notes in Computer Science, 7444:132–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33250-0_12/.

Slaviero:2012:EEU

[801] Cleyton Slaviero, Ana Cristina Bicharra Garcia, and Cristiano Maciel. Exploiting eParticipation using an ontological approach. Lecture Notes in Computer Science, 7444:144–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33250-0_13/.

Charalabidis:2012:PPF

[802] Yannis Charalabidis, Anna Triantafillou, Vangelis Karkaletsis, and Euripidis Loukis. Public policy formulation through non moderated crowdsourcing in social media. Lecture Notes in Computer Science, 7444: 156–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33250-0_14/.

Anonymous:2012:BMl

[803] Anonymous. Back matter. Lecture Notes in Computer Science, 7444: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33250-0/1.

Anonymous:2012:FMbo

[804] Anonymous. Front matter. Lecture Notes in Computer Science, 7444: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33250-0/1.

Bonizzoni:2012:HGF

[805] Paola Bonizzoni. The holy grail: Finding the genetic bases of phenotypic characters. Lecture Notes in Computer Science, 7445:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32894-7_1.

Calude:2012:ICP

[806] Cristian S. Calude, Elena Calude, and Melissa S. Queen. Inductive complexity of P versus NP problem. Lecture Notes in Computer Science, 7445:2–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32894-7_2/.

Doursat:2012:AEE

[807] René Doursat. Advances in embryomorphic engineering. Lecture Notes in Computer Science, 7445:10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-32894-7_3.

Lutz:2012:RT

[808] Jack H. Lutz. Reasoning as though. Lecture Notes in Computer Science, 7445:11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-32894-7_4.

Margenstern:2012:UHP

[809] Maurice Margenstern. Universality and the halting problem for cellular automata in hyperbolic spaces: The side of the halting problem. Lecture Notes in Computer Science, 7445:12–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32894-7_5/.

Patitz:2012:ITB

[810] Matthew J. Patitz. An introduction to tile-based self-assembly. Lecture Notes in Computer Science, 7445:34–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32894-7_6/.

Spicher:2012:SCM

[811] Antoine Spicher, Olivier Michel, and Jean-Louis Giavitto. Spatial computing in MGS. Lecture Notes in Computer Science, 7445:63–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32894-7_7/.

Csuhaj-Varju:2012:PSC

[812] Erzsébet Csuhaj-Varjú, Marian Gheorghe, and Mike Stannett. P systems controlled by general topologies. Lecture Notes in Computer Science, 7445:70–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32894-7_8/.

Freund:2012:PSM

[813] Rudolf Freund, Yurii Rogozhin, and Sergey Verlan. P systems with minimal left and right insertion and deletion. Lecture Notes in Computer Science, 7445:82–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32894-7_9/.

Goliaei:2012:LBC

[814] Sama Goliaei and Mohammad-Hadi Foroughmand-Araabi. Lower bounds on the complexity of the wavelength-based machine. Lecture Notes in Computer Science, 7445:94–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32894-7_10/.

Grozea:2012:SMI

[815] Cristian Grozea, Florin Manea, Mike Müller, and Dirk Nowotka. String matching with involutions. Lecture Notes in Computer Science, 7445: 106–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32894-7_11/.

Gruau:2012:DEA

[816] Frederic Gruau and Luidnel Maignan. Distributed execution of automata networks on a computing medium: Introducing IfAny machines. Lecture Notes in Computer Science, 7445:118–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32894-7_12/.

Gruenert:2012:SRE

[817] Gerd Gruenert, Gabi Escuela, and Peter Dittrich. Symbol representations in evolving droplet computers. Lecture Notes in Computer Science, 7445: 130–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32894-7_13/.

Hertel:2012:ICG

[818] Joachim Hertel. Inductive complexity of Goodstein’s theorem. Lecture Notes in Computer Science, 7445:141–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32894-7_14/.

Lakin:2012:TBL

[819] Matthew R. Lakin, Amanda Minnich, Terran Lane, and Darko Stefanovic. Towards a biomolecular learning machine. Lecture Notes in Computer Science, 7445:152–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32894-7_15/.

Milici:2012:TMM

[820] Pietro Milici. Tractional motion machines: Tangent-managing planar mechanisms as analog computers and educational artifacts. Lecture Notes in Computer Science, 7445:164–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32894-7_16/.

Montoya:2012:CSC

[821] J. Andres Montoya. Computing with sand: On the complexity of recognizing two-dimensional sandpile critical configurations. Lecture Notes in Computer Science, 7445:174–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32894-7_17/.

Nichele:2012:GPI

[822] Stefano Nichele and Gunnar Tufte. Genome parameters as information to forecast emergent developmental behaviors. Lecture Notes in Computer Science, 7445:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32894-7_18/.

Stepney:2012:HCE

[823] Susan Stepney, Samson Abramsky, Matthias Bechmann, Jerzy Gorecki, and Viv Kendon. Heterotic computing examples with optics, bacteria, and chemicals. Lecture Notes in Computer Science, 7445:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32894-7_19/.

Torbey:2012:RNP

[824] Sami Torbey and Selim G. Akl. Reliable node placement in wireless sensor networks using cellular automata. Lecture Notes in Computer Science, 7445:210–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32894-7_20/.

Anonymous:2012:FMbp

[825] Anonymous. Front matter. Lecture Notes in Computer Science, 7445: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32894-7/1.

Furche:2012:DDD

[826] Tim Furche, Georg Gottlob, and Christian Schallhart. DIADEM: Domains to databases. Lecture Notes in Computer Science, 7446:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32600-4_1/.

Ait-Ameur:2012:SDF

[827] Yamine Ait-Ameur and Idir Ait-Sadoune. Stepwise development of formal models for Web services compositions: Modelling and property verification. Lecture Notes in Computer Science, 7446:9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-32600-4_2.

Wu:2012:HAG

[828] Huayu Wu, Ruiming Tang, Tok Wang Ling, Yong Zeng, and Stéphane Bressan. A hybrid approach for general XML query processing. Lecture Notes in Computer Science, 7446:10–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32600-4_3/.

OConnor:2012:SCS

[829] Martin F. O’Connor and Mark Roantree. SCOOTER: a compact and scalable dynamic labeling scheme for XML updates. Lecture Notes in Computer Science, 7446:26–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32600-4_4/.

Zhuang:2012:RDL

[830] Canwei Zhuang and Shaorong Feng. Reuse the deleted labels for vector order-based dynamic XML labeling schemes. Lecture Notes in Computer Science, 7446:41–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32600-4_5/.

Ou:2012:TEF

[831] Yi Ou, Jianliang Xu, and Theo Härder. Towards an efficient flashbased mid-tier cache. Lecture Notes in Computer Science, 7446:55–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32600-4_6/.

Bhushan:2012:EPL

[832] Alka Bhushan, Nandlal L. Sarda, and P. V. Rami Reddy. Evacuation planning of large buildings using ladders. Lecture Notes in Computer Science, 7446:71–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32600-4_7/.

Vamsikrishna:2012:WEP

[833] Meduri Venkata Vamsikrishna, Zhan Su, and Kian-Lee Tan. A write efficient PCM-aware sort. Lecture Notes in Computer Science, 7446: 86–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32600-4_8/.

Ferrarotti:2012:PAA

[834] Flavio Ferrarotti, Sven Hartmann, Sebastian Link, Mauricio Marin, and Emir Muñoz. Performance analysis of algorithms to reason about XML keys. Lecture Notes in Computer Science, 7446:101–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32600-4_9/.

Ikeda:2012:FTC

[835] Kosetsu Ikeda and Nobutaka Suzuki. Finding top-K correct XPath queries of user’s incorrect XPath query. Lecture Notes in Computer Science, 7446: 116–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32600-4_10/.

Bruhathi:2012:APD

[836] H. S. Bruhathi and Jayant R. Haritsa. Analyzing plan diagrams of XQuery optimizers. Lecture Notes in Computer Science, 7446:131–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32600-4_11/.

Chatvichienchai:2012:SME

[837] Somchai Chatvichienchai. Spreadsheet metadata extraction: a layoutbased approach. Lecture Notes in Computer Science, 7446:147–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32600-4_12/.

Atapattu:2012:AES

[838] Thushari Atapattu, Katrina Falkner, and Nickolas Falkner. Automated extraction of semantic concepts from semi-structured data: Supporting computer-based education through the analysis of lecture notes. Lecture Notes in Computer Science, 7446:161–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32600-4_13/.

Oliveira:2012:CWM

[839] Hilário Oliveira, Rinaldo Lima, João Gomes, Rafael Ferreira, and Fred Freitas. A confidence–weighted metric for unsupervised ontology population from Web texts. Lecture Notes in Computer Science, 7446: 176–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32600-4_14/.

Sassi:2012:SAU

[840] Imen Ben Sassi, Chiraz Trabelsi, Amel Bouzeghoub, and Sadok Ben Yahia. Situation-aware user’s interests prediction for query enrichment. Lecture Notes in Computer Science, 7446:191–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32600-4_15/.

Abdulahhad:2012:ERL

[841] Karam Abdulahhad, Jean-Pierre Chevallet, and Catherine Berrut. The effective relevance link between a document and a query. Lecture Notes in Computer Science, 7446:206–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32600-4_16/.

Bouadi:2012:ICS

[842] Tassadit Bouadi, Marie-Odile Cordier, and René Quiniou. Incremental computation of skyline queries with dynamic preferences. Lecture Notes in Computer Science, 7446:219–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32600-4_17/.

Kiran:2012:EDC

[843] R. Uday Kiran and Masaru Kitsuregawa. Efficient discovery of correlated patterns in transactional databases using items’ support intervals. Lecture Notes in Computer Science, 7446:234–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32600-4_18/.

Tort:2012:CEC

[844] Albert Tort, Antoni Olivé, and Maria-Ribera Sancho. On checking executable conceptual schema validity by testing. Lecture Notes in Computer Science, 7446:249–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32600-4_19/.

Huo:2012:QTT

[845] Wenyu Huo and Vassilis J. Tsotras. Querying transaction–time databases under branched schema evolution. Lecture Notes in Computer Science, 7446:265–280, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32600-4_20/.

Anonymous:2012:FMbq

[846] Anonymous. Front matter. Lecture Notes in Computer Science, 7446: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32600-4/1.

Deagustini:2012:CQA

[847] Cristhian A. D. Deagustini, Santiago E. Fulladoza Dalibón, and Sebastián Gottifredi. Consistent query answering using relational databases through argumentation. Lecture Notes in Computer Science, 7447:1– 15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32597-7_1/.

Jenkins:2012:ADL

[848] John Jenkins, Isha Arkatkar, Sriram Lakshminarasimhan, Neil Shah, and Eric R. Schendel. Analytics-driven lossless data compression for rapid insitu indexing, storing, and querying. Lecture Notes in Computer Science, 7447:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32597-7_2/.

Yu:2012:PWU

[849] Xiuming Yu, Meijing Li, Incheon Paik, and Keun Ho Ryu. Prediction of Web user behavior by discovering temporal relational rules from Web log data. Lecture Notes in Computer Science, 7447:31–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32597-7_3/.

Navarro-Galindo:2012:HAT

[850] José Luis Navarro-Galindo, José Samos, and M. José Muñoz-Alférez. A hybrid approach to text categorization applied to semantic annotation. Lecture Notes in Computer Science, 7447:39–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32597-7_4/.

Loglisci:2012:UFT

[851] Corrado Loglisci, Dino Ienco, Mathieu Roche, Maguelonne Teisseire, and Donato Malerba. An unsupervised framework for topological relations extraction from geographic documents. Lecture Notes in Computer Science, 7447:48–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32597-7_5/.

Nieves:2012:CML

[852] Javier Nieves, Igor Santos, and Pablo G. Bringas. Combination of machine-learning algorithms for fault prediction in high-precision foundries. Lecture Notes in Computer Science, 7447:56–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32597-7_6/.

Tang:2012:FCU

[853] Ruiming Tang, Reynold Cheng, Huayu Wu, and Stéphane Bressan. A framework for conditioning uncertain relational data. Lecture Notes in Computer Science, 7447:71–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32597-7_7/.

Awano:2012:CAN

[854] Yuki Awano, Qiang Ma, and Masatoshi Yoshikawa. Cause analysis of new incidents by using failure knowledge database. Lecture Notes in Computer Science, 7447:88–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32597-7_8/.

Khefifi:2012:MQC

[855] Rania Khéfifi, Pascal Poizat, and Fatiha Saïs. Modeling and querying context-aware personal information spaces. Lecture Notes in Computer Science, 7447:103–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32597-7_9/.

Bahmani:2012:OBR

[856] Amir Bahmani, Sahra Sedigh, and Ali Hurson. Ontology-based recommendation algorithms for personalized education. Lecture Notes in Computer Science, 7447:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32597-7_10/.

Ahmed:2012:TQC

[857] Eya Ben Ahmed, Ahlem Nabli, and Faïez Gargouri. Towards quantitative constraints ranking in data clustering. Lecture Notes in Computer Science, 7447:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32597-7_11/.

Kang:2012:TOA

[858] Kyu-Hwang Kang, Seung-Hwan Lim, Sang-Wook Kim, Min-Hee Jang, and Byeong-Soo Jeong. A topic-oriented analysis of information diffusion in a blogosphere. Lecture Notes in Computer Science, 7447: 129–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32597-7_12/.

Hasegawa:2012:TTS

[859] Keisuke Hasegawa, Qiang Ma, and Masatoshi Yoshikawa. Trip tweets search by considering spatio-temporal continuity of user behavior. Lecture Notes in Computer Science, 7447:141–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32597-7_13/.

Vermaas:2012:ICC

[860] Raymond Vermaas, Damir Vandic, and Flavius Frasincar. Incremental cosine computations for search and exploration of tag spaces. Lecture Notes in Computer Science, 7447:156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32597-7_14/.

Kurabayashi:2012:IAV

[861] Shuichi Kurabayashi and Yasushi Kiyoki. Impression-aware video stream retrieval system with temporal color-sentiment analysis and visualization. Lecture Notes in Computer Science, 7447:168–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32597-7_15/.

Liroz-Gistau:2012:DWB

[862] Miguel Liroz-Gistau, Reza Akbarinia, Esther Pacitti, Fabio Porto, and Patrick Valduriez. Dynamic workload-based partitioning for largescale databases. Lecture Notes in Computer Science, 7447:183–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32597-7_16/.

Rodriguez:2012:DVP

[863] Lisbeth Rodríguez and Xiaoou Li. Dynamic vertical partitioning of multimedia databases using active rules. Lecture Notes in Computer Science, 7447:191–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32597-7_17/.

Jedrzejczak:2012:RBB

[864] Jacek Jedrzejczak, Tomasz Koszlajda, and Robert Wrembel. RTDWbench: Benchmark for testing refreshing performance of real-time data warehouse. Lecture Notes in Computer Science, 7447:199–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32597-7_18/.

Furtado:2012:MLS

[865] Pedro Furtado. Middleware and language for sensor streams. Lecture Notes in Computer Science, 7447:207–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32597-7_19/.

Correndo:2012:SAO

[866] Gianluca Correndo, Antonio Penta, Nicholas Gibbins, and Nigel Shadbolt. Statistical analysis of the owl:sameAs network for aligning concepts in the Linking Open Data Cloud. Lecture Notes in Computer Science, 7447: 215–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32597-7_20/.

Anonymous:2012:FMbr

[867] Anonymous. Front matter. Lecture Notes in Computer Science, 7447: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32597-7/1.

ElAkkaoui:2012:BBC

[868] Zineb El Akkaoui, José-Norberto Mazón, Alejandro Vaisman, and Esteban Zimányi. BPMN-based conceptual modeling of ETL processes. Lecture Notes in Computer Science, 7448:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32584-7_1/.

Cuzzocrea:2012:ECE

[869] Alfredo Cuzzocrea and Robson do N. Fidalgo. Enhancing coverage and expressive power of spatial data warehousing modeling: The SDWM approach. Lecture Notes in Computer Science, 7448:15–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32584-7_2/.

Golfarelli:2012:SPO

[870] Matteo Golfarelli, Stefano Rizzi, and Elisa Turricchia. Sprint planning optimization in agile data warehouse design. Lecture Notes in Computer Science, 7448:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32584-7_3/.

Atigui:2012:UOA

[871] Faten Atigui, Franck Ravat, Olivier Teste, and Gilles Zurfluh. Using OCL for automatically producing multidimensional models and ETL processes. Lecture Notes in Computer Science, 7448:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32584-7_4/.

Benker:2012:CSM

[872] Thomas Benker and Carsten Jürck. A case study on model-driven data warehouse development. Lecture Notes in Computer Science, 7448: 54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32584-7_5/.

Jovanovic:2012:IEP

[873] Petar Jovanovic, Oscar Romero, Alkis Simitsis, and Alberto Abelló. Integrating ETL processes from information requirements. Lecture Notes in Computer Science, 7448:65–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32584-7_6/.

Alrayes:2012:ATM

[874] Norah Alrayes and Wo-Shun Luk. Automatic transformation of multidimensional Web tables into data cubes. Lecture Notes in Computer Science, 7448:81–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32584-7_7/.

Hassan:2012:DMA

[875] Ali Hassan, Franck Ravat, Olivier Teste, Ronan Tournier, and Gilles Zurfluh. Differentiated multiple aggregations in multidimensional databases. Lecture Notes in Computer Science, 7448:93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-32584-7_8/.

Fuad:2012:GAB

[876] Muhammad Marwan Muhammad Fuad. Genetic algorithms-based symbolic aggregate approximation. Lecture Notes in Computer Science, 7448: 105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32584-7_9/.

Abdelbaki:2012:NBA

[877] Wiem Abdelbaki, Riadh Ben Messaoud, and Sadok Ben Yahia. A neuralbased approach for extending OLAP to prediction. Lecture Notes in Computer Science, 7448:117–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32584-7_10/.

Gomes:2012:MAR

[878] João Bártolo Gomes, Shonali Krishnaswamy, Mohamed M. Gaber, and Pedro A. C. Sousa. Mobile activity recognition using ubiquitous data stream mining. Lecture Notes in Computer Science, 7448:130–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32584-7_11/.

Groger:2012:WMD

[879] Christoph Gröger, Johannes Schlaudraff, Florian Niedermann, and Bernhard Mitschang. Warehousing manufacturing data. Lecture Notes in Computer Science, 7448:142–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32584-7_12/.

Kerkad:2012:QBQ

[880] Amira Kerkad, Ladjel Bellatreche, and Dominique Geniet. Queen-Bee: Query interaction-aware for buffer allocation and scheduling problem. Lecture Notes in Computer Science, 7448:156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32584-7_13/.

Lee:2012:EDP

[881] Suan Lee, Jinho Kim, Yang-Sae Moon, and Wookey Lee. Efficient distributed parallel top-down computation of ROLAP data cube using MapReduce. Lecture Notes in Computer Science, 7448:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32584-7_14/.

Narita:2012:LJH

[882] Kazuyo Narita, Shinji Nakadai, and Takuya Araki. Landmark-Join: Hashjoin based string similarity joins with edit distance constraints. Lecture Notes in Computer Science, 7448:180–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32584-7_15/.

Oguz:2012:IIM

[883] Damla Oguz and Belgin Ergenc. Incremental itemset mining based on matrix apriori algorithm. Lecture Notes in Computer Science, 7448: 192–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32584-7_16/.

Qu:2012:FAF

[884] Jun-Feng Qu and Mengchi Liu. A fast algorithm for frequent itemset mining using patricia* structures. Lecture Notes in Computer Science, 7448:205–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32584-7_17/.

Yang:2012:MOO

[885] Hang Yang, Simon Fong, and Yain-Whar Si. Multi-objective optimization for incremental decision tree learning. Lecture Notes in Computer Science, 7448:217–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32584-7_18/.

deAmo:2012:MCP

[886] Sandra de Amo, Mouhamadou Saliou Diallo, Cheikh Talibouya Diop, and Arnaud Giacometti. Mining contextual preference rules for building user profiles. Lecture Notes in Computer Science, 7448:229–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32584-7_19/.

Simovici:2012:PAM

[887] Dan A. Simovici, Paul Fomenky, and Werner Kunz. Polarities, axiallities and marketability of items. Lecture Notes in Computer Science, 7448: 243–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32584-7_20/.

Anonymous:2012:FMbs

[888] Anonymous. Front matter. Lecture Notes in Computer Science, 7448: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32584-7/1.

Hirsch:2012:HMN

[889] Tobias Hirsch, Luigi Lo Iacono, and Ina Wechsung. How much network security must be visible in Web browsers? Lecture Notes in Computer Science, 7449:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32287-7_1/.

Braun:2012:ULA

[890] Bastian Braun, Stefan Kucher, Martin Johns, and Joachim Posegga. A user-level authentication scheme to mitigate Web session-based vulnerabilities. Lecture Notes in Computer Science, 7449:17–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32287-7_2/.

Casalino:2012:ACC

[891] Matteo Maria Casalino, Romuald Thion, and Mohand-Said Hacid. Access control configuration for J2EE Web applications: a formal perspective. Lecture Notes in Computer Science, 7449:30–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32287-7_3/.

Delport:2012:CSS

[892] Waldo Delport and Martin S. Olivier. Cloud separation: Stuck inside the cloud. Lecture Notes in Computer Science, 7449:36–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32287-7_4/.

Syed:2012:PCP

[893] Kashif Syed, Pavol Zavarsky, Dale Lindskog, Ron Ruhl, and Shaun Aghili. Proposed control procedure to mitigate the risks of strategic information outflow in the recruitment process. Lecture Notes in Computer Science, 7449:50–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32287-7_5/.

Netter:2012:ASW

[894] Michael Netter, Sabri Hassan, and Günther Pernul. An autonomous social Web privacy infrastructure with context-aware access control. Lecture Notes in Computer Science, 7449:65–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32287-7_6/.

Lazouski:2012:PEU

[895] Aliaksandr Lazouski, Fabio Martinelli, and Paolo Mori. A prototype for enforcing usage control policies based on XACML. Lecture Notes in Computer Science, 7449:79–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32287-7_7/.

Moyano:2012:CFT

[896] Francisco Moyano, Carmen Fernandez-Gago, and Javier Lopez. A conceptual framework for trust models. Lecture Notes in Computer Science, 7449:93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32287-7_8/.

Bukac:2012:ACS

[897] Vit Bukac, Pavel Tucek, and Martin Deutsch. Advances and challenges in standalone host-based intrusion detection systems. Lecture Notes in Computer Science, 7449:105–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32287-7_9/.

Gorawski:2012:EAS

[898] Marcin Gorawski, Michal Lorek, and Michal Gorawski. Encrypted adaptive storage model — analysis and performance tests. Lecture Notes in Computer Science, 7449:118–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32287-7_10/.

Dorn:2012:ECE

[899] Michael Dorn, Peter Wackersreuther, and Christian Böhm. Efficient comparison of encrypted biometric templates. Lecture Notes in Computer Science, 7449:129–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32287-7_11/.

Behnia:2012:SEI

[900] Rouzbeh Behnia, Swee-Huay Heng, and Che-Sheng Gan. Short and efficient identity-based undeniable signature scheme. Lecture Notes in Computer Science, 7449:143–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32287-7_12/.

DAcquisto:2012:DSM

[901] Giuseppe D’Acquisto, Marta Flamini, and Maurizio Naldi. Damage sharing may not be enough: An analysis of an ex-ante regulation policy for data breaches. Lecture Notes in Computer Science, 7449:149– 160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32287-7_13/.

Langweg:2012:FRP

[902] Hanno Langweg and Lisa Rajbhandari. Flexible regulation with privacy points. Lecture Notes in Computer Science, 7449:161–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32287-7_14/.

Schick:2012:SNR

[903] Rainer Schick and Christoph Ruland. On the security of the nonrepudiation of forwarding service. Lecture Notes in Computer Science, 7449:167–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-32287-7_15/.

Elshaafi:2012:PCM

[904] Hisain Elshaafi, Jimmy McGibney, and Dmitri Botvich. Profitability and cost management of trustworthy composite services. Lecture Notes in Computer Science, 7449:179–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32287-7_16/.

Thong:2012:QAP

[905] Ta Vinh Thong and Levente Buttyán. Query auditing for protecting max/ min values of sensitive attributes in statistical databases. Lecture Notes in Computer Science, 7449:192–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32287-7_17/.

Salem:2012:VSC

[906] Ali Salem, Salah Triki, Hanêne Ben-Abdallah, Nouria Harbi, and Omar Boussaid. Verification of security coherence in data warehouse designs. Lecture Notes in Computer Science, 7449:207–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32287-7_18/.

Vigano:2012:TSP

[907] Luca Viganò. Towards the secure provision and consumption in the Internet of Services. Lecture Notes in Computer Science, 7449:214–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-32287-7_19.

Johns:2012:WSD

[908] Martin Johns and Joachim Posegga. WebSand: Server-driven outbound Web-application sandboxing. Lecture Notes in Computer Science, 7449: 216–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-32287-7_20.

Anonymous:2012:FMbt

[909] Anonymous. Front matter. Lecture Notes in Computer Science, 7449: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32287-7/1.

Jensen:2012:TPN

[910] Kurt Jensen, Wil M. van der Aalst, Marco Ajmone Marsan, Giuliana Franceschinis, Jetty Kleijn, and Lars Michael Kristensen, editors. Transactions on Petri Nets and Other Models of Concurrency VI, volume 7400 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35178-6 (print), 3-642-35179-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
35179-2.

Hao:2012:AEI

[911] Jin-Kao Hao, Pierrick Legrand, Pierre Collet, Nicolas Monmarché, Evelyne Lutton, and Marc Schoenauer, editors. Artificial Evolution: 10th International Conference, Evolution Artificielle, EA 2011, Angers, France, October 24-26, 2011, Revised Selected Papers, volume 7401 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35532-3 (print), 3-642-35533-1 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-35533-2.

Lin:2012:COA

[912] Guohui Lin, editor. Combinatorial Optimization and Applications: 6th International Conference, COCOA 2012, Banff, AB, Canada, August 5-9, 2012. Proceedings, volume 7402 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31769-3 (print), 3-642-31770-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31770-5.

Navarro:2012:SSA

[913] Gonzalo Navarro and Vladimir Pestov, editors. Similarity Search and Applications: 5th International Conference, SISAP 2012, Toronto, ON, Canada, August 9-10, 2012. Proceedings, volume 7404 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32152-6 (print), 3-642-32153-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32153-5.

Wang:2012:WAS

[914] Xinbing Wang, Rong Zheng, Tao Jing, and Kai Xing, editors. Wireless Algorithms, Systems, and Applications: 7th International Conference, WASA 2012, Yellow Mountains, China, August 8-10, 2012. Proceedings, volume 7405 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31868-1 (print), 3-642-31869-X (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31869-6.

Beringer:2012:ITP

[915] Lennart Beringer and Amy Felty, editors. Interactive Theorem Proving: Third International Conference, ITP 2012, Princeton, NJ, USA, August 13-15, 2012. Proceedings, volume 7406 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32346-4 (print), 3-642-32347-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32347-8.

Gupta:2012:ARC

[916] Anupam Gupta, Klaus Jansen, José Rolim, and Rocco Servedio, editors. Approximation, Randomization, and Combinatorial Optimization. Algorithms and Techniques: 15th International Workshop, APPROX 2012, and 16th International Workshop, RANDOM 2012, Cambridge, MA, USA, August 15-17, 2012. Proceedings, volume 7408 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-32511-4 (print), 3-642-32512-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-32512-0.

deSouto:2012:ABC

[917] Marcilio C. de Souto and Maricel G. Kann, editors. Advances in Bioinformatics and Computational Biology: 7th Brazilian Symposium on Bioinformatics, BSB 2012, Campo Grande, Brazil, August 15-17, 2012. Proceedings, volume 7409 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31926-2 (print), 3-642-31927-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31927-3.

Yen:2012:DLT

[918] Hsu-Chun Yen and Oscar H. Ibarra, editors. Developments in Language Theory: 16th International Conference, DLT 2012, Taipei, Taiwan, August 14-17, 2012. Proceedings, volume 7410 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31652-2 (print), 3-642-31653-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31653-1.

Cheung:2012:HLI

[919] Simon K. S. Cheung, Joseph Fong, Lam-For Kwok, Kedong Li, and Reggie Kwan, editors. Hybrid Learning: 5th International Conference, ICHL 2012, Guangzhou, China, August 13-15, 2012. Proceedings, volume 7411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32017-1 (print), 3-642-32018-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32018-7.

Smith:2012:ITS

[920] Adam Smith, editor. Information Theoretic Security: 6th International Conference, ICITS 2012, Montreal, QC, Canada, August 15-17, 2012. Proceedings, volume 7412 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32283-2 (print), 3-642-32284-0 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-32284-6.

Yao:2012:RSC

[921] JingTao Yao, Yan Yang, Roman Słowiński, Salvatore Greco, Huaxiong Li, Sushmita Mitra, and Lech Polkowski, editors. Rough Sets and Current Trends in Computing: 8th International Conference, RSCTC 2012, Chengdu, China, August 17-20, 2012.Proceedings, volume 7413 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32114-3 (print), 3-642-32115-1 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32115-3.

Li:2012:RSK

[922] Tianrui Li, Hung Son Nguyen, Guoyin Wang, Jerzy Grzymala-Busse, Ryszard Janicki, Aboul Ella Hassanien, and Hong Yu, editors. Rough Sets and Knowledge Technology: 7th International Conference, RSKT 2012, Chengdu, China, August 17-20, 2012. Proceedings, volume 7414 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31899-1 (print), 3-642-31900-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31900-6.

Lassiter:2012:NDL

[923] Daniel Lassiter and Marija Slavkovik, editors. New Directions in Logic, Language and Computation: ESSLLI 2010 and ESSLLI 2011 Student Sessions. Selected Papers, volume 7415 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31466-X (print), 3-642-31467-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31467-4.

Rofer:2012:RRS

[924] Thomas Röfer, N. Michael Mayer, Jesus Savage, and Uluç Saranlı, editors. RoboCup 2011: Robot Soccer World Cup XV, volume 7416 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32059-7 (print), 3-642-32060-0 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32060-6.

Safavi-Naini:2012:ACC

[925] Reihaneh Safavi-Naini and Ran Canetti, editors. Advances in Cryptology — CRYPTO 2012: 32nd Annual Cryptology Conference, Santa Barbara, CA, USA, August 19-23, 2012. Proceedings, volume 7417 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32008-2 (print), 3-642-32009-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32009-5.

Gao:2012:WAI

[926] Hong Gao, Lipyeow Lim, Wei Wang, Chuan Li, and Lei Chen, editors. Web-Age Information Management: 13th International Conference, WAIM 2012, Harbin, China, August 18-20, 2012. Proceedings, volume 7418 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32280-8 (print), 3-642-32281-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32281-5.

Bao:2012:WAI

[927] Zhifeng Bao, Yunjun Gao, Yu Gu, Longjiang Guo, Yingshu Li, Jiaheng Lu, Zujie Ren, Chaokun Wang, and Xiao Zhang, editors. Web-Age Information Management: WAIM 2012 International Workshops: GDMM, IWSN, MDSP, USDM, and XMLDM Harbin, China, August 18-20, 2012 Proceedings, volume 7419 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33049-5 (print), 3-642-33050-9 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33050-6.

Beckert:2012:FVO

[928] Bernhard Beckert, Ferruccio Damiani, and Dilian Gurov, editors. Formal Verification of Object-Oriented Software: International Conference, FoVeOOS 2011, Turin, Italy, October 5-7, 2011, Revised Selected Papers, volume 7421 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31761-8 (print), 3-642-31762-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31762-0.

Mahjoub:2012:COS

[929] A. Ridha Mahjoub, Vangelis Markakis, Ioannis Milis, and Vangelis Th. Paschos, editors. Combinatorial Optimization: Second International Symposium, ISCO 2012, Athens, Greece, April 19-21, 2012, Revised Selected Papers, volume 7422 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32146-1 (print), 3-642-32147-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32147-4.

Geva:2012:FRC

[930] Shlomo Geva, Jaap Kamps, and Ralf Schenkel, editors. Focused Retrieval of Content and Structure: 10th International Workshop of the Initiative for the Evaluation of XML Retrieval, INEX 2011, Saarbrücken, Germany, December 12-14, 2011, Revised Selected Papers, volume 7424 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35733-4 (print), 3-642-35734-2 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-35734-3.

Lee:2012:CHI

[931] Geuk Lee, Daniel Howard, Jeong Jin Kang, and Dominik Ślezak, editors. Convergence and Hybrid Information Technology: 6th International Conference, ICHIT 2012, Daejeon, Korea, August 23-25, 2012. Proceedings, volume 7425 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32644-7 (print), 3-642-32645-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-32645-5.

Ziemke:2012:AAI

[932] Tom Ziemke, Christian Balkenius, and John Hallam, editors. From Animals to Animats 12: 12th International Conference on Simulation of Adaptive Behavior, SAB 2012, Odense, Denmark, August 27-30, 2012. Proceedings, volume 7426 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33092-4 (print), 3-642-33093-2 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33093-3.

Kuhn:2012:CNL

[933] Tobias Kuhn and Norbert E. Fuchs, editors. Controlled Natural Language: Third International Workshop, CNL 2012, Zurich, Switzerland, August 29-31, 2012. Proceedings, volume 7427 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32611-0 (print), 3-642-32612-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32612-7.

Prouff:2012:CHE

[934] Emmanuel Prouff and Patrick Schaumont, editors. Cryptographic Hardware and Embedded Systems — CHES 2012: 14th International Workshop, Leuven, Belgium, September 9-12, 2012. Proceedings, volume 7428 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33026-6 (print), 3-642-33027-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33027-8.

Herrmann:2012:AAR

[935] Guido Herrmann, Matthew Studley, Martin Pearson, Andrew Conn, Chris Melhuish, Mark Witkowski, Jong-Hwan Kim, and Prahlad Vadakkepat, editors. Advances in Autonomous Robotics: Joint Proceedings of the 13th Annual TAROS Conference and the 15th Annual FIRA RoboWorld Congress, Bristol, UK, August 20-23, 2012, volume 7429 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32526-2 (print), 3-642-32527-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32527-4.

Nguyen:2012:TCC

[936] Ngoc-Thanh Nguyen, editor. Transactions on Computational Collective Intelligence VIII, volume 7430 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34644-8 (print), 3-642-34645-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34645-3.

Bebis:2012:AVCa

[937] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Charless Fowlkes, Sen Wang, Min-Hyung Choi, Stephan Mantler, Jürgen Schulze, Daniel Acevedo, Klaus Mueller, and Michael Papka, editors. Advances in Visual Computing: 8th International Symposium, ISVC 2012, Rethymnon, Crete, Greece, July 16-18, 2012, Revised Selected Papers, Part I, volume 7431 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33178-5 (print), 3-642-33179-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33179-4.

Bebis:2012:AVCb

[938] George Bebis, Richard Boyle, Bahram Parvin, Darko Koracin, Charless Fowlkes, Sen Wang, Min-Hyung Choi, Stephan Mantler, Jürgen Schulze, Daniel Acevedo, Klaus Mueller, and Michael Papka, editors. Advances in Visual Computing: 8th International Symposium, ISVC 2012, Rethymnon, Crete, Greece, July 16-18, 2012, Revised Selected Papers, Part II, volume 7432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33190-4 (print), 3-642-33191-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33191-6.

Stefanovic:2012:DCM

[939] Darko Stefanovic and Andrew Turberfield, editors. DNA Computing and Molecular Programming: 18th International Conference, DNA 18, Aarhus, Denmark, August 14-17, 2012. Proceedings, volume 7433 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32207-7 (print), 3-642-32208-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32208-2.

Gudmundsson:2012:CCA

[940] Joachim Gudmundsson, Julián Mestre, and Taso Viglas, editors. Computing and Combinatorics: 18th Annual International Conference, COCOON 2012, Sydney, Australia, August 20-22, 2012. Proceedings, volume 7434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32240-9 (print), 3-642-32241-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32241-9.

Yin:2012:IDE

[941] Hujun Yin, José A. F. Costa, and Guilherme Barreto, editors. Intelligent Data Engineering and Automated Learning — IDEAL 2012: 13th International Conference, Natal, Brazil, August 29-31, 2012. Proceedings, volume 7435 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32638-2 (print), 3-642-32639-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32639-4.

Giannakopoulou:2012:FFM

[942] Dimitra Giannakopoulou and Dominique Méry, editors. FM 2012: Formal Methods: 18th International Symposium, Paris, France, August 27-31, 2012. Proceedings, volume 7436 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32758-3 (print), 3-642-32759-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32759-9.

Stoelinga:2012:FMI

[943] Mariëlle Stoelinga and Ralf Pinger, editors. Formal Methods for Industrial Critical Systems: 17th International Workshop, FMICS 2012, Paris, France, August 27-28, 2012. Proceedings, volume 7437 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32468-1 (print), 3-642-32469-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32469-7.

Bikakis:2012:RWR

[944] Antonis Bikakis and Adrian Giurca, editors. Rules on the Web: Research and Applications: 6th International Symposium, RuleML 2012, Montpellier, France, August 27-29, 2012. Proceedings, volume 7438 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32688-9 (print), 3-642-32689-7 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32689-9.

Xiang:2012:AAPa

[945] Yang Xiang, Ivan Stojmenovic, Bernady O. Apduhan, Guojun Wang, Koji Nakano, and Albert Zomaya, editors. Algorithms and Architectures for Parallel Processing: 12th International Conference, ICA3PP 2012, Fukuoka, Japan, September 4-7, 2012, Proceedings, Part I, volume 7439 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33077-0 (print), 3-642-33078-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33078-0.

Xiang:2012:AAPb

[946] Yang Xiang, Ivan Stojmenovic, Bernady O. Apduhan, Guojun Wang, Koji Nakano, and Albert Zomaya, editors. Algorithms and Architectures for Parallel Processing: 12th International Conference, ICA3PP 2012, Fukuoka, Japan, September 4-7, 2012, Proceedings, Part II, volume 7440 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33064-9 (print), 3-642-33065-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33065-0.

Alvarez:2012:PPR

[947] Luis Alvarez, Marta Mejail, Luis Gomez, and Julio Jacobo, editors. Progress in Pattern Recognition, Image Analysis, Computer Vision, and Applications: 17th Iberoamerican Congress, CIARP 2012, Buenos Aires, Argentina, September 3-6, 2012. Proceedings, volume 7441 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33274-9 (print), 3-642-33275-7 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-33275-3.

Gerdt:2012:CAS

[948] Vladimir P. Gerdt, Wolfram Koepf, Ernst W. Mayr, and Evgenii V. Vorozhtsov, editors. Computer Algebra in Scientific Computing: 14th International Workshop, CASC 2012, Maribor, Slovenia, September 3-6, 2012. Proceedings, volume 7442 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32972-1 (print), 3-642-32973-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32973-9.

Scholl:2012:EGI

[949] Hans J. Scholl, Marijn Janssen, Maria A. Wimmer, Carl Erik Moe, and Leif Skiftenes Flak, editors. Electronic Government: 11th IFIP WG 8.5 International Conference, EGOV 2012, Kristiansand, Norway, September 3-6, 2012. Proceedings, volume 7443 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33488-1 (print), 3-642-33489-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33489-4.

Tambouris:2012:EPI

[950] Efthimios Tambouris, Ann Macintosh, and Øystein Sæbø, editors. Electronic Participation: 4th IFIP WG 8.5 International Conference, ePart 2012, Kristiansand, Norway, September 3-5, 2012. Proceedings, volume 7444 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33249-8 (print), 3-642-33250-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33250-0.

Durand-Lose:2012:UCN

[951] Jérôme Durand-Lose and Nataša Jonoska, editors. Unconventional Computation and Natural Computation: 11th International Conference, UCNC 2012, Orléan, France, September 3-7, 2012. Proceedings, volume 7445 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32893-8 (print), 3-642-32894-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
32894-7.

Liddle:2012:DESa

[952] Stephen W. Liddle, Klaus-Dieter Schewe, A. Min Tjoa, and Xiaofang Zhou, editors. Database and Expert Systems Applications: 23rd International Conference, DEXA 2012, Vienna, Austria, September 3-6, 2012. Proceedings, Part I, volume 7446 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32599-8 (print), 3-642-32600-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32600-4.

Liddle:2012:DESb

[953] Stephen W. Liddle, Klaus-Dieter Schewe, A. Min Tjoa, and Xiaofang Zhou, editors. Database and Expert Systems Applications: 23rd International Conference, DEXA 2012, Vienna, Austria, September 3-6, 2012. Proceedings, Part II, volume 7447 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32596-3 (print), 3-642-32597-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32597-7.

Cuzzocrea:2012:DWK

[954] Alfredo Cuzzocrea and Umeshwar Dayal, editors. Data Warehousing and Knowledge Discovery: 14th International Conference, DaWaK 2012, Vienna, Austria, September 3-6, 2012. Proceedings, volume 7448 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32583-1 (print), 3-642-32584-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32584-7.

Fischer-Hubner:2012:TPS

[955] Simone Fischer-Hübner, Sokratis Katsikas, and Gerald Quirchmayr, editors. Trust, Privacy and Security in Digital Business: 9th International Conference, TrustBus 2012, Vienna, Austria, September 3-7, 2012. Proceedings, volume 7449 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32286-7 (print), 3-642-32287-5 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-32287-7.